A Complete Bibliography of Publications in SIGCSE Bulletin: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 April 2018 Version 1.08

Title word cross-reference

3 [CH98]. χ^{2} [DSB94]. N [Hig93].

* [Tow99].

-D [CH98].

1 [AR95, Bro92, CPS93, DH93, DH94, HAB^{+}94, JCF^{+}93, ML96, Pat93, Pat94, Rei93]. 1-July [Imp99a]. 11 [Ful91]. 1990 [ACMJ90, VS92]. 1991 [Dal91b]. 1992 [Dal92b]. 1993 [KLY93]. 1994 [BG94]. 1995 [LWMG95]. 1996 [FR96a, IAK96]. 1997 [SH97, WEKM97]. 1998 [Car98a, LPJI98]. 1999 [BHH^{+}99, PN99]. 1st [FR96a].

2 [HH95c, JPPM96, Wal99a]. 2000 [Ger98]. 21st [ACMJ90, CEMP97, Gor99b, OQ98]. 22nd [Dal91b]. 23rd [Dal92b].

24-March [JLP99]. 24th [KLY93]. 25th [BG94]. 26th [LWMG95]. 27-June [Imp99a]. 27th [IAK96]. 28th [WEKM97]. 29th [ACM98, LPJI98]. 2nd [SH97].

30th [PN99]. 360 [Det90]. 3rd [Car98a]. 3RolesPlaying [Bob99].

4th [Imp99a].

’68 [Lit90].

80x86 [Det90].

’91 [GL93]. 95 [CP96]. ’97 [CDF^{+}97]. 9X [CB94b].

= [GY97a, MAK^{+}99]. == [Tow99].

A.S. [HJ92]. AAA [AR95]. AACSB [GL94]. AAPT [SG91]. abilities [Chm98, SD93]. ability [Bai91, TJ90]. abound [McC98b]. Abstract [HSS98, AHST93, ABM93, AWB94, BO95, BMB^{+}95, BBC^{+}94, Ber94, BMR^{+}95, BBG^{+}95, BCHB94, Bro94, BSM^{+}94, BBD^{+}95, CDMT94, CFA^{+}94, CHLS95, CB94b, DAUZ94, DFZ^{+}95, Eps93, FCR^{+}94, FF94, GPJ95, GVLZ94, GZ94, GSK^{+}95, GKB^{+}95, GNR^{+}95, HKH^{+}93, Has94, HAB^{+}94, HANWN93, IBF^{+}94, Jes94, JCF^{+}93, KG94, KI93, KACL94, KS94, LGK^{+}93, LAH94, LEDW93, MZS^{+}95, MAG^{+}94c, MAD^{+}95, MC90a, McN94, MSRT95, MBE^{+}93, MCd^{+}94b, MIG^{+}94, MPHL95, MDG^{+}95, Nev94, Noh94, OBH^{+}95, PBEC93, PFR^{+}94, RFH^{+}93, SH95, SVP93, SDJ^{+}94, SVFB94, SCF^{+}93, Sil94, Smi94b, Smi94a, Tro94a, WCMT94, WBBW94, WBCS95, WCDT95, WGD^{+}95, Yer94]. abstraction [Kel93, Lev93, Náv94]. academia [PD95, TW92]. Academic [NRR^{+}90, Bie90, DFH^{+}98, LHA90, MW98b, Rob99a, WMZ^{+}99]. accelerated [SBD97]. access [Blu96, Gro97a]. accessibility [Kno97a]. accessible [Cun96, HF98]. Accommodating [DFH^{+}98]. accompany [Joy90]. account [MRS93]. Accreditation [MAC^{+}90, Roz94, Gor98b, Gor99c, IBF^{+}94, JBCL98, LGBM99, TM92]. accredited [GL94, MM97, MM98]. accrediting [Tay97a]. achievement [CH91]. Achieving [Gur95]. ACM [Car98a, FR96a, SH97, AMN98, BO95, BSM^{+}94, CDF^{+}97, Got98b, KBCP99, Kle99a, Mar98b, MBE^{+}93, NPP^{+}93, PBEC93, PCRR90, RS97, RSLD99, WGD^{+}95]. ACM/AIS/DPMA [CDF^{+}97]. ACM/DPMA/AIS [WGD^{+}95]. ACM/IEEE [AMN98, RSLD99]. ACM/IEEE-CS [RSLD99]. ACOM [DR99]. acquisition [RL95]. across [EFGR94, FK92, FPC96, HMBO98, HANWN93, Wol98]. ACSE [Car98a, FR96a, SH97]. Act [VS92, GPP99]. action [Cle97b, Mar99b, Mar99c, Win93]. Active [McC96, Moo95, CHL96, Fal99]. Activities [RW96, Kle99a, Kur91]. actual [MPTW92]. actually [Tja98].

Ada [Fel90, bY94, Mur91, CB94b, CP96, FACW90, FLP90, FS90, KP90, LAH94, MCD^{+}94a, Rad90, Sil91]. adaptation [MR99]. adaptive [Gro97a, Pre94, SD93, VS92]. Adding [HS97, JPKPM99, SW95]. addition [Bre92]. Addressing [HRBK99]. adequacy [Lar97]. administration [DH95a, DH96a, Ric98, Rug99]. administrative [Bry99]. admissions [NC99b]. ADT [Ber90, Dal90, Eva95]. adults [Mor92]. Advanced [SAL^{+}90, SFR^{+}98, Ack99, AG96, ACRW97, BAH97, Bro93b, CFA^{+}94, Cun96, EL94, HMS93, KS94, Lea90, Mat95, MK99, NKN^{+}95, Owe90, SRL^{+}99, Tic95, TA96]. adventure [CB96, Mos97]. Advisor [Har94a]. advocating [Nul98]. afraid [KG94, MHB^{+}98]. After [BC96, GY97b]. again [bY91]. agenda [RL95]. agent [Rug99]. agents [RM98]. AI [AAW92, IHA^{+}94, Noy93]. aid [Dad90, Moo95, PS91, RG98b, TGGP97]. aided [Gol96a, Gol97a, GP91, GCDP94]. Aids [Bar97a, Sta97]. AIS [CDF^{+}97, WGD^{+}95]. album [RR96]. ALEX [PS91]. Alfonse [BC96, Har98a, Har99a]. algebra [FP95, LHA90, MSSZ97, WN96]. algebraic [Kol98]. ALGOL [Lam92]. Algorithm [JM96, Nap90, Nap96, Nap97b, BBF^{+}94, Chu99, CCMD99, GdLC94, Har91, KLMS95, Kro96a, Kro96b, KMO95, Lev99c, MSPF95, MdL92, MV95, MN90, NS94, Nap98, PH96, Ras99, RST97, SG91, Shi93b, Sta97, SSN99, TSG98, VIRF96, WAK96, YJ95]. Algorithma [CCMD99]. algorithmics [dM98]. Algorithms [Gei91, Bal90, BA97, Ber97b, Bie93, Bra96, BV97, CGP90, Cha96, DS98, Den96, Ele96, Epp92, Fal99, Fek93, FA98, Gha91, Gin96, GEC96, GT98, GT99b, Hag91, Har94b, JF97, Jan97, KS98, KH99, KPT98, KPT99, KCA99, Kor94, KKL99, MBJS90, MB95, McN94, McN96, Mit96, NB98, NC99a, NW99a, Noh94, PR92, Pre99, PWST94, PS91]. Algorthmic [DBKL90]. Alice [Dan92]. alive [RBL^{+}97]. allowing [Joh98]. ALTERNATING [GA91]. Alternative [DFZ^{+}95, Cla93, Cup98, EWC94, FJM98, FH96, HHK95, MCH98, MMB^{+}91, Pra90, WSK97]. always [KG94]. amelioration [FHL^{+}97]. Americans [VS92]. among [Cle97a]. analog [Lev99b]. analogy [Col91a, Eck90, MCH98]. Analysis [BCHB94, FG92, BM96, BBF^{+}94, BC92, Cha96, EN90, Epp92, Fek93, FA98, GT98, JD97b, KMH91, Kro96a, Kro96b, LS93, Lar97, MO96, MB95, MY99, Noy95, Osb92b, PH96, RPF94, SV96, SW95, Tro91]. analysts [Oml99]. analytic [Hay98]. analyze [Noh94, Shi93b]. analyzer [LC90]. Analyzing [Bra96]. anecdote [FP98]. animate [Ber94]. animated [BGG^{+}99, HDC99, Jon96b]. Animating [Har94b, KS98]. Animation [AR98, Ber94, BFG^{+}94, BBG^{+}95, CCMD99, DB98, DST98, JM96, JF97, MdL92, McC98b, PR98a, RPF94, Ras99, RST97, SA95, SSN99, WAK96]. animations [KPT98, Rod96, Sta97]. animator [Ros91a, SG91]. AnLex [PMCGJP97]. Annual [Imp99a, MM97, MM98]. ANSI [FR96b]. AnSin [PMCGJP97]. Antonio [Dal91b]. APPGEN [Osb92b]. Apple [Ful91, Ful92]. applets [YLQ98]. Application [ASW97, CKM^{+}97, Dra98, RL95, Sap91, Sha98]. Application-based [ASW97]. applications [AAHM99, Cas96, DS91, Fri91, HMS93, MPTW92, MC90b, Szu96, Tic95, YLQ98]. applied [AR95, Ciz97, Cup98, ES99, GZ94, HG99, SPSS96, WR90]. Applying [Dal91a, HT98, LC96a, Nul96, MS90]. apprentice [ASW97]. apprenticeship [AR95]. approach [AR95, Ast98b, BHH^{+}90, Ben92, Ben95, Bie90, CB96, Cla98b, Cli91, CH93b, Dal98c, DR99, DFF99, DDG97, DL95, Gha91, Got98a, Had99, Has94, Hay98, Hei93, HP94, Hil93, HW91, HHK95, HL94, Jen98, Joh98, KK98a, Kay92b, Kir96, Kou98, Leo91, LZ98, LAH94, LSS92, Mag94b, McC99c, MM94b, Moo93, Mot91, Mur91, ML90, NS94, NW92b, New93, Noy93, PM95, Pat93, Pax95, Ree95, Rod95, SH96, SR91, Sul93, TP93, WR97, Wah99, WSK97, Zac97a]. Approaches [CPS93, GKB^{+}95, SCF^{+}93, AAK^{+}94, BWW96, CP96, Got99b, LaR93, WBC^{+}98, YSKW95]. Approaching [Gor99a]. Appropriate [PL96, Law97b]. arcade [CGP90]. architecture [Ber93, Can91, CWI96, CM96, EWC94, Fol92, GMAB96, Kno97b, Nev94, Pil93, Sty94, Tan90, dlFCC99]. architecture/programming [Can91]. architectures [Imp99b, Rei95a]. area [Alm98, PSZ99a, RF96]. arena [Har98b]. Argentina [SLM98]. arithmetic [Chu99, MS90]. Arizona [BG94]. array [Hil90, Ras99]. arrays [WW91, Zai96]. art [Cig92, HL94]. ART199 [Mos98]. ART199/CS199 [Mos98]. articles [CC92]. Articulation [KACL94]. Artificial [WV91, AAW92, CH93a, KW96, KM98, LaR93, LR94, McC99b, Pat97, Pax95, Stu94, TW92, TP94, Tos93]. artist [KB97]. arts [APNA90, Arn91, Arn94, AAK^{+}94, Gol99, Has94, KH94, Pra92, SSH98, TLR94, WBCS95]. ASA [MdL92, GdLC94]. ask [KG94, MHB^{+}98]. Asking [Bar97a]. aspects [Sta93]. asperations [Fel99]. assault [GPP99]. assembler [Don90, Wil90]. assembly [Bar95, Coe93, Det90]. assertions [Wal98b]. Assesses [Aik99]. Assessing [CKK94, MWP94, PCLM97, Pet98a, Pet98b, SBD99, SD93]. Assessment [CDMT94, Cul98, HL98, JL98, KCD^{+}96, Kum99, MM99, PS99, Ros97, Sch95, Zac97b]. assessment-possibilities [Cul98]. assets [Tow98]. assignment [MS90, RSB98]. assignments [AR98, Bri93, CPS93, Dad90, DH95a, DH96a, Ele96, FZ96, GPJ95, Mit96, PAC^{+}99, PP97, TS98, Wal98a, Wal99a, WW91]. assistance [KSIR94, Siv96]. assistant [BP94, FRK96, KMO95]. assistants [Kay95, RLR95]. Associate [KLL^{+}98, GHK93, KBCP99]. Associate-level [KLL^{+}98]. Association [Bro94]. ASSYST [JU97]. asynchronous [BS99]. Atlanta [ACM98, LPJI98]. ATM [SR99]. attachment [Got99a]. Attitudes [PP95, MWP94, Mor92, Tow96]. attract [RW96]. Attracting [Bru94, TC93]. attributes [BCHB94]. audio [Lev99b, RRR93]. augmented [RSB98]. Austin [Wal97a]. Australasian [Car98a, FR96a, SH97]. Australia [Car98a, FR96a, SH97]. authentic [HP95]. author [WH96]. authoring [Reb98]. Automata [GR99, Han92, RHND99, RBL^{+}97, RG98b]. Automated [Har94a, KSIR94, RK98, HL98, LC90, Sch95]. Automatic [DH95a, DH96a, Ras99, Hit99, MGH98]. automation [Kir96, Sap91]. Avenue [DH96c]. Average [FA98, Tro91, Cha91a]. Average-case [FA98]. avoid [Mil96]. Avoiding [HGW97]. awareness [Mar99b, Mar99c, PH97].

AWK [Ros91b].

baccalaureate [LN97]. BACCII [CB94a]. Backtracking [Dor92, Zie98]. bad [Car99b]. balance [Wal98a]. balanced [Kay92b]. Bandwagons [Kay96]. barriers [SS98]. base [Uck91]. based [AGT98, Abu92, AMN98, Asl98, ASW97, Ast98b, BK95, BN98, BGG^{+}99, Cas98a, Cas96, CWS98, CC92, Cur97, Dal98c, DB98, DEP97, DL95, EW97, Ell98a, GL93, Gol97b, HP94, Hol98a, JLH^{+}98, JRS96, KK98b, Kir96, Kot98, LG96, Liv97, LH99, Mag94a, Mak98, MO96, Mit96, ML90, Nak98, Nap97a, New93, PD96, PH99, PRS93, PR98a, Rob95a, RHND99, Ros99, Rub92, SH95, SP99, Sch96a, SBK94, SBD99, SR91, TBF97, WP98, Wal96, WN96, Zac94]. bases [Uck91]. basic [LLP95, Mak98, Cas96]. basics [Mak97]. be [DH94, FB97, GSW^{+}98, KCA99, LGK^{+}93, Mar99a, Mos97, Rad90, Wis92]. Begin [Mot91, Tem91]. Begin-BIG [Mot91]. Beginners [Kos99, BT98, Gin96, PMCGJP97, Tam92]. beginning [Llo94]. behavior [Ras97b, Tao95]. behaviour [CCR99]. believe [Goe96]. benefits [DeC96, LR94, WHN93]. best [Bru94]. better [CB94a, LN98, Mak97, Sil90]. Between [JPT^{+}99, Alm91, Cle99a, Eck90, Hvo90, LM98b, MV95, PD95, Wal98a]. Beyond [FP98, Náv94]. biased [LWBS99]. Bibliography [Dal99a, For91a]. BIG [Mot91]. Binary [Mal97, Ada96b, BD96, Cha91a, Gur95, Hig93, Kho90, KMH91, Ras97c, Zie98]. Biologically [Abu92]. BlackBox [War99]. blacks [CH91]. Bloom [HCL96]. Blue [KR96a]. board [Fal99]. book [MCD^{+}94a, Ros97, TGGP97]. Borland [Szu96]. both [DH96c]. bottom [KW94]. bottom-up [KW94]. boundaries [Ast92, Wol98]. box [Cun98]. branch [Ben95]. breadth [BMC95, BMB^{+}95, BBC^{+}94, JPKPM99, Moo93, Noy93, PMB98, PRS93, SW95, SL94, Wil95b]. breadth-first [BMC95, BBC^{+}94, Moo93, Noy93, PMB98, PRS93, Wil95b]. breaking [Pli98]. bridge [PD95, Sch99, Son96]. bridging [Cle98b]. brightest [Bru94]. Bringing [GYT^{+}93, Nac98, CEK^{+}99, EL97]. Brisbane [Car98a]. broad [For91a]. broad-coverage [For91a]. Broadening [RC97]. browser [GMF99]. browsers [BGGR98]. browsing [Kno96]. budgets [SH99a]. build [BLOS98, Kno98a]. builders [Kos99]. Building [Dun91, Eng99a, FK92, Gol98, KPT98, GG99, MPW99, Reb98, Rei92]. built [BMP96, Scr91, Sta97]. built-in [BMP96]. Bulletin [ABDM90, IMAM98]. BURKS [Eng99b, Van98]. business [GL94, Tre98, WUGV99]. bytes [Har99b]. Byzantine [Llo94].

’C [HR93, ACRW97, Ast98b, BAH97, BG93, BDN^{+}94, Chu99, Don90, Dun91, Eps93, FP97, FHS^{+}94, FR96b, GF96, GSH96, Hea95, Hit99, HH95c, HK97, Hug92, LC96b, Mod91, MS98, NKN^{+}95, Osb92a, Osb95, Pat97, Ras97b, Ras97c, Ras99, Rei91, Rei95a, Rob93, Rob95a, Ros96, RH93, Sca96, Sch97, SFR^{+}98, Wic95, Zai96]. C-based [Rob95a]. C-language [GF96]. C.S [Cul97, SWWB97, Tic95]. Cache [Leb99, CCR99, NG94]. CacheSim [CCR99]. CAI [GM91]. calculator [Lee98b]. California [WEKM97]. calls [Hol97, Rob99b]. CALOS [Gol96a, Gol97a]. Campus [PH99, Oli94, Tho98b, TLNS98]. Campus-based [PH99]. campus-wide [Tho98b, TLNS98]. campuses [FHL^{+}97]. Can [Rad90, TJ90, Coh99, Kay98a, KV94, McC99b, Mos97, MIG^{+}94, FB97]. CANLOGS [Mey96]. can’t [DH94, Goe96]. CAP [Sch95]. capabilities [Cle98a]. capstone [Ada98, Ewi93]. Carnegie [GBB^{+}90]. Case [CFA^{+}94, CL92, Ast92, Ber92a, Din97, FA98, GA91, Har91, HW91, Hol99, HAG92, LN98, MY99, MN90, NKN^{+}95, PH96, Pol99, Pro98, SS97, Tro91, Wal97a, GVLZ94, GL96, McK95, ML90]. CASE-based [ML90]. cases [TS95]. CBE [MMK98]. CCRs [OM95]. CD [Eng99a, Van98, VW96b, WGHL97]. CD-ROM [Eng99a, Van98, WGHL97]. CD-ROMs [VW96b]. center [Kno99a]. centered [Ada96c, HHK95]. central [Mye90]. Centralized [Guz95]. centre [GK98]. centric [LM98c]. century [CEMP97, Gor99b, OQ98, Tay98]. certificate [KBCP99]. certification [TM92]. chair [Kle99b]. challenge [CFK^{+}92, MHN98, TW92]. Challenges [MBC^{+}96, WUGV99]. Chance [Ada98]. Chance-It [Ada98]. change [Kay96, RM98, ZBD99]. Changes [NKN^{+}95, HCD^{+}98]. Changing [KGZ99, OBPB97, Cun98, Kum99, OQ98, SY98]. characteristics [OBH^{+}95]. cheating [Dad90]. check [Sch95]. checking [PSZ99a]. chess [Mäk96]. chessboard [Lee98b, Lee99b]. child [STM95]. children [WDM97]. China [HWS92]. choice [FJM98, JCF^{+}93]. Church [Tay98]. CIS [GL96]. City [Dal92b, BR99]. CLAP [MST97]. Clarifying [HR93]. Class [BBC^{+}94, Ber92a, Dra98, DN97, DV98, Dow99, Hea95, Hit99, Nap98, Ras97c, RR96, Sha98, Sie99, SH91, Tic95, TP94, WDM97]. classes [AC96, Ast98b, HS98, Kay98b, PS98, Szu96, Wal99b, Zac94]. classification [EN90]. Classroom [O’N95, RR97, Ada93b, Ber92b, CEK^{+}99, CL92, Gei94, GYT^{+}93, HML^{+}99, HMS93, HRBK99, KK98a, MCW99, Moo95, Nac98, Nul96, OQ98, Pax96, SKL98, Sch92]. classroom-ICTs [OQ98]. classrooms [Mos93]. classwide [KP90]. Clever [dOG93]. click [CLV99]. Client [LWBS99, DMK97, PW99, YB99]. client-server [DMK97]. client/server [YB99]. closed [BW97, FG96, LWC96, MAK^{+}99, PD96, Pre96, PFR^{+}94, Thw94]. closer [Ste91]. club [GY98b]. clustering [EKR98]. Co [Har93, HKH^{+}93]. co-ed [HKH^{+}93]. Co-oping [Har93]. Coatroom [Ber90]. COCS [Pra90]. Code [AR99, BT96, Got98b, Lam92, Mar98b, Pli98, PM96]. codes [Ben95]. codesign [EO95]. cognition [Cle97b]. cognitive [DL95, Hol99, WDB98]. cognitive-based [DL95]. coherence [NG94]. cohesion [Cle97a]. Collaboration [Car99c, Hed99, JLH^{+}98, PS97, WHN93]. collaborations [BK92b]. Collaborative [FBR^{+}99, FCB^{+}99, MPHL95, SS94, Wal97a, CHCL99, Cle98a, Cle99a, DDP96, EHS^{+}99, LZ98, Mak97, MPW99, Siv96, TH98, Wil97, WP97, YR95]. collection [Mit96, RK98, RBL^{+}97]. college [Alm91, AJ93, Cam94, CD95, Kle99a, MAC^{+}90, Mag94b, Osb92a, Sch96b, Tja96, Tur94, TLR94, WBC^{+}98, Wal97a]. colleges [ICD^{+}90]. collider [Har93]. combined [Oli94]. Combining [MPW99]. come [Kno98a, RBL^{+}97]. comes [AHST93]. Comfort [Ber91]. coming [HCD^{+}98]. Command [Gre96a]. Command-line [Gre96a]. commercial [MPTW92, Seb97]. committee [Kle99a]. Communication [Eck90, BCHB94, Car97, Cle98a, Cun91, Der93, FCR^{+}94, KP90, MM93, NW99b, Pol99, RFPA97, Sha91, Sul93, Tay97b, WP97]. communications [BS99, CHLS95, DMW^{+}97, GMAB96, Hug90, Lev97, Lev99a, Lev99b, SG97b, Smi91, Smi94b, Smi98, Tol95b, WCC^{+}92]. community [Ber97a, Bon99, Got98a, Mag94b, SVFB94, Tur94]. comparative [HWS92, LS93, Leo91]. comparatively [Max93]. Comparison [GF96, Det90, EN90, GBC^{+}99, LM98b]. compendium [SH99a]. competency [WV91]. compiler [Ber93, Coo96b, Don95, Els92, Hai97, Lam90, LM98b, PMCGJP97, Rei90, Tem92, Wil90]. compiler-writing [Rei90]. compilers [dOG94, RD98]. Compiling [Hay97, Lam90, Ber93, Nef99]. complementary [Mak98]. complexity [MSPF95, SSN99]. compliance [VS92]. component [AMBC98, Cha91b, Joh94, Mat95, SVP93, Wen90]. components [HTW98, NF97]. comprehensive [Zac97a]. compressed [Mil96]. compression [LN91]. Compuquake [SVFB94]. Computability [Gre90, Hei93]. computation [BFG^{+}94, GEC96, HA95, Joh92, Joh94, Kro92, Kro96a, Kro96b, RHND99]. Computational [MZS^{+}95, Sch99, DBKL90, SSH98, ZJEP95]. Compute [Tro92]. Computer [ACM98, Ama90, ACMJ90, BG94, Car98a, Cle99a, Cup98, Dal91b, Dal92b, DP90, DMW^{+}97, Ele99, FR96a, FW96, Gro97a, HCGW99, HANWN93, IAK96, Imp99a, JBCL98, JLP99, Kay98a, KB97, KLY93, LK98, LWMG95, Lev95, LM98a, LPJI98, LEB^{+}90, McC97, MM98, MCMHS90, MM99, NW99b, PH95, PH96, Pil93, Pra92, PN99, Ros97, Rub95, Sch91, Sch96b, SBK94, SHY96, Siv96, SH97, SPV99, WCMT94, WEKM97, Wis96, WN96, WP97, Ågr99, AGT98, AMN98, AFVH97, Add97, APNA90, Alm91, AP91, AW93, ABM93, ABW93, AWB94, Arn91, AAK^{+}94, AM90, Ast92, Bal90, BK92a, BSK94, Bal96a, Bal97, Bam98, BM96, BBJR96, BMP96, BHH^{+}90, BA98, Ben92, BNB^{+}98, Ber93, Ber92a, Bie90, BBG^{+}95, BKS^{+}99, BGGR98, BGG^{+}99, Bou96, BJB98, Bro97, BCHB94, BSM^{+}94, BABT97, BKLV98]. computer [BD93, BBD^{+}95, CH98, Cam94, Cam96, CK98, CDMT94, Car90, CGK^{+}96, CK99, CFA^{+}94, Che94a, CH91, Chm98, Cla98b, CHLS95, CWI96, CD95, Coh99, CM96, Coo96a, Coo97, CKM^{+}97, Cun91, Cun94, Dal98c, Dal99d, DGEST96, DAUZ94, DH90, DM92, DL95, Duc94, DL90, EGIU99, EWC94, Epp91, Eps95, Ewi93, FK92, Fek95, FG96, FACW90, FLP90, FPC96, FPR98, FRK96, FC94, FMM97, FP95, FH96, Fol92, FCR^{+}94, Fra98, FHL^{+}97, GS93, Gei91, GY98b, GT99a, Gol96a, Gol97a, Gol97b, GEC96, GBL94, Gol96b, GJMS90, Got98a, Got98b, Got99a, Got99b, GP91, GZ94, Gre96b, Gre97, Gre99, Gri99, GKB^{+}95, GNR^{+}95, GKC^{+}98, GK98, GMAB96, GCDP94, Gur98, Haf94, HTW98, HKH^{+}93, HH95a, HJ92, HCD^{+}98, HH95b, HWB92, HW90, Har98b, HP94, Hil93, HT97]. computer [Hil90, Hol98a, HL94, HKSD97, How93, Hus95, IBF^{+}94, JPS90, JMM97, Jan97, Joh92, Joh94, Joh97, Joh99b, JM91, Jon96b, Joy98, Kat94, Kat95, Kay92b, Kay92a, Kay93, Kay95, Kay98b, KCC^{+}98, KW93, KCD^{+}96, KI93, KLL^{+}98, KBCP99, KS94, KPW94, KWJ^{+}96, Kno97b, Kno99a, KA95, Kol98, KD90, KH94, Kot98, Kou98, Kub99, Kum98, KDE^{+}93, Lam95, Lan98, Leb99, Lee96, Lee97, LW97, Lee98a, Lee99a, Lel94, LZ98, LB96, Llo94, Loc91, LR94, MMKW99, Mag94a, Mag94b, MPH96, Mar97, Mar98c, MW98a, MBJS90, MAD^{+}95, MJM96, MM97, McC98a, McC99a, McC96, McC92, MF98, MM93, MS97, MMB^{+}91, MH96c, Mer92, MBE^{+}93, MSSZ97, Moo93, MWP94, Mor92, MD99, Mos93, Muk99, MHN98, Mye90, Mye92, MW98b, Nac98, Nad96, Nap90, Nef93, NKN^{+}95, NP92, NC99b, Nul96, OK95, Ola91, Owe90]. computer [Owe92, OBPB97, PM97, PP95, PD96, PCLM97, PRS93, PRS94, PA90, Pes91, PL96, Pol99, Pre96, PRF96, Rab94, RPF94, Ras97a, Ree96a, Ree96b, Ree97, Rei95a, RF96, RFH^{+}93, RS91, RS94, RC97, Rod95, RW96, Rog90, RSB98, Sab90, SS94, SP96, SBB^{+}97, SP99, SG96, SA95, Sco91, SVP93, Scr91, SS98, SW95, She95, Shi92, Sie99, SBD97, Sty94, SD93, Tan90, TM92, TP93, Tay97a, Tay97b, Tho90, Tho98b, TLNS98, Thu90, TH98, Tow96, Tra98, Tri99, TJ90, TW94, Tym91, VS92, VB98, WV91, Wah99, WBCS95, WCDT95, Wal98c, WMZ^{+}99, Wal99d, WRRT97, Wen90, Wer97, WFGW94, WC94, Yah92, YD99, YR95, YJ99, ZMS95, ZBK^{+}90, dlFCC99, Kum99]. computer-aided [Gol96a, Gol97a, GP91]. Computer-based [WN96, Kot98]. computer-equipped [Bal97]. computer-mediated [Tay97b]. Computer-related [PH95]. Computerized [SD93, Mit98]. Computerland [Dan92]. computers [Eck90, FG91, KV94, MPW99, Mor92, PM97, Rei92, Smi94b]. Computing [Ber97a, Bro94, Cha91a, Den99, DSB94, Jon96a, LGB^{+}99, PRS93, Tur94, Amo93, AJ93, BMB^{+}95, Ber98, Ber91, BCPV98, Bon99, BEG^{+}96, Bro93b, CS98, Car97, CT94, Cun98, Cup98, CH93b, Cyn90, DP90, DI98, DG99, DR99, DDG97, GS99, GHK93, HS93, HMBO98, HHK95, HMS93, HRBK99, ICD^{+}90, IG97, Jes94, JY95, Kin98, KST92, KI93, Kno96, KWJ^{+}96, KJ97, KKA98, LSS92, MZS^{+}95, MSRT95, Mer92, MV98, Mot91, NI98, Nev94, OBH^{+}95, PMB98, PH97, Rog90, Rub92, SK95, SRY98, SL94, SK94, Sig90, SS97, Sta93, STM95, TC93, Tho98b, TLNS98, VBB^{+}98, WH96]. computing-five [SK95]. concept [CHCL99]. Concepts [Gei94, MM94a, Ågr99, BT95, Cam96, Cle99a, DV98, Duc94, JMM97, Jef91, KST92, KA95, LLP95, LWC96, Náv94, Nef93, OM95, Rab94, Rif94b, Sab90, SR99, SR91, WR97, WGHL97]. Conceptual [WDB98]. concern [Kat95, Lee99a]. Concrete [Ast98a]. Concurrency [Hil92, BA96b, BAK99, Fel92, Jac91, KCPC98, Lea90, McD92, PBA98, Tro94b, Yea91]. Concurrent [FB97, JTW90, Ber96, Dav90, EN90, FLP90, HM91, Nev99, OM95, VI94, bY94]. condition [Ben95]. conditions [NSS98]. conduct [Mar98b]. conducting [AIW^{+}96]. Conference [BHH^{+}99, Car98a, FR96a, Imp99a, SH97, BJ98, NW99b]. conferences [Utt99]. conferencing [BM96, Cun94, PM97]. confidence [Tea92]. configuration [Cur97]. congruence [HTW98]. Connect [Stu94]. connection [Blu96, Hvo90]. connectivity [RR97]. conscious [Leb99]. Conserving [Rob99a]. considerations [Yah92]. considered [Kay96]. consistent [SAL^{+}90]. console [WD99]. constants [O’N95]. construct [Fle91]. construction [Coo96b, Lam90, Pil93, Tem92]. Constructivism [BA98]. constructivist [GBS98, Had98a, Had99]. contained [BGG^{+}99, Eng99a]. Content [GY97a, Kat95, SSN99]. contents [NB99]. Contest [NPP^{+}93, PCRR90, BO95, Bro97, FH96, Str91, KAK93]. Contest-1977 [PCRR90]. context [Cle98c, Cle98b, Wha91, ZBD99]. context-bridging [Cle98b]. contexts [NB99]. contextual [Bre95]. continents [DGEST96]. continuous [Nul98]. Contributions [GY98a]. control [Cur97]. controlled [CHL96]. controls [EN90]. controversial [YSKW95]. conventional [Mak98]. conversation [AC96, MF98]. Converses [JM90]. cooperating [Koo95]. cooperation [IHA^{+}94, MMKW99, VL90]. Cooperative [Pre96, CLL^{+}96, DU96]. Copyright [GPP99]. core [Gur98, HH95b, Kub99]. Corn [Cle99b, Rob99a]. coroutines [Rev98]. correctional [Ama90]. Correctness [Wad95]. correlation [Dad90]. Cost [PLG97, Hug90]. counting [WCMT94]. Coupling [Cle97a]. Course [MR99, ABFM90, Abu92, Ada93a, AAW92, AAHF^{+}99, APNA90, Amo93, And90b, Asl98, BMC95, Bar92, Bel92, Bie90, BFG^{+}94, BF94, BS99, Bru94, Bru99, BD93, BBD^{+}95, CO95, Can91, Car95, CK99, Cha93, Cha91b, Cli91, CS99, CM96, Con98a, Coo96a, Coo97, Coo96b, CL98, Cun91, Dal91a, Dal99d, DP90, DH90, DA97, EGIU99, EWC94, Eri96, Etl90, Eva95, Eva96, Ewi93, Fek93, Fek95, FG96, FA98, Fie96, FC94, FW96, FC99b, Fol92, Gam95, Ger94, Gha91, Gre99, GMAB96, Gur98, HSM97, Hai97, HWB92, Har92, HS91, Hei99, Hil93, HCGW99, Hol97, HL94, HCL96, Hus95, IHA^{+}94, Jac91, Joy98, Kay92a, Kay98a, Kay98b, KK98b, Kin92, KSW98, Kno97b, KR96b, KR99, Kor94, Kro92, Lam90, Lam95, LBJ94, Lee90, LN91, Lel94, Lev93, LH92, Loc91, LAH94]. course [LR94, MMKW99, ME92, Mat95, McC99b, Mer90, Mes91, MV98, MM94b, MH91, MP96, Mit96, Moo92, MS98, Mor92, Mot91, MRS93, Nak98, NS96, NKN^{+}95, Noy95, OJ93, Owe90, OCH^{+}94, PM97, Pan90, PCLM97, Pil93, PR92, PW99, Pra90, Pre99, RSBAS98, Reb98, Ree90, Ree97, Rei95b, Rub92, SS94, SF91, Sca96, Sch97, Sch90, SW95, Seb97, Sha91, She98, Sil91, Smi94b, Sob96, SBD97, Sul93, Taf91, Tol98, Tos93, Tow98, Tre98, Tre99, Tro94a, TW97, TZ99, Uck91, ULW99, Van98, VW96b, Wal99d, Wal96, WG93, Wen90, Wil90, WFGW94, WR90, Yea91, bY90, bY94, ZJEP95, BKP^{+}99]. course-long [TZ99]. Courses [BKP^{+}99, Wit96, Ack99, AFVH97, Ast92, Ast98b, BBC^{+}94, BS95, CS98, CCR^{+}97, CKK94, Cor99, Cun96, Det90, DL95, EL94, FPR98, HSC^{+}90, Har94a, HP95, Hug90, HK96, JPKPM99, KG94, KCC^{+}98, KJ97, Kou97, Kou98, LZ98, MPH96, MBJS90, MAG^{+}94c, McC99a, MHP98, MSSZ97, NI98, NP92, Pes91, PL96, Pou90, Ree96a, RLR95, Rod96, SP96, SA95, Scr91, SS97, TC93, TA96, WB93, Wol98, Zie99]. courseware [Bou96, Fag94, Fal99, FLP90, Gol96b, GBC^{+}99, Hol98b, Kum98, Law97a, LWC96, LP99, MH96a, PLG97, TBF97, WP98]. coursework [RK98]. coverage [For91a]. CPU [KH99, Tan92]. Cracow [Imp99a]. create [Tra98]. created [Mes91]. creates [VL90]. Creating [BS95, Eze90, GVLZ94, HKH^{+}93, Hed99, HP95, Rob92, Szu96, BGG^{+}99, Rob96]. Creation [GA91, KS94, Law97a]. Creative [FBR^{+}99, MST97]. credit [Osb92a]. crisis [Rob99a, WMZ^{+}99]. criteria [How94, Sze99]. Critical [LWC96, BF94, Had98b, Kou97, MWP94]. critically [Cun97]. critique [BAH97]. crop [Cle99b]. crossing [Ast92]. CS [BKP^{+}99, Ada98, Alm98, AC96, AR95, ASW97, AR98, AMBC98, BKP^{+}99, BS95, Bro92, BD93, Ciz97, CPS93, CDD^{+}97, Cun95, DH93, DH94, Der93, DFZ^{+}95, Dyc93, Fie96, FF94, Ger94, Got99a, GL96, GL97, HF98, HA95, HAB^{+}94, HANWN93, JCF^{+}93, Kno97c, Kum96, LBK96, LM98a, Lim98, LWC96, MN93, ML96, MAK^{+}99, Pat90, Pat93, Pat94, PM96, Rei93, Rei95b, RP98, RSLD99, Sha99, Tja98, WBBW94, WG93, WDM97]. CS-1 [Ada98, Dyc93, Ger94, Pat90]. CS-2 [Fie96]. CS/CIS [GL96]. CS/IS [Cun95, GL97, Lim98]. CS0 [Coo97]. CS1 [Ada96c, Bar96a, Bar96b, BW97, BDN^{+}94, CH93a, Car99b, Cur92, DFH^{+}98, Eva96, FZ96, FP97, FS90, GPJ95, Gol99, HS97, HT98, JPKPM99, Joy90, KP91, KW99, LS93, Leo91, Lev95, LWB^{+}98, Mar99a, MSPF95, MCD^{+}94a, McN98, Med98, OD94, OCH^{+}94, PLU97, Par95, Pra90, Ree98, Rob93, Rob95a, SG97a, SBB^{+}99, TBF98, Wal97a, Wal98b, Wal99a, Wal94, War99, Wic95, Wil95b, WD99]. CS1-2 [Ada96c]. CS1/2 [Wal99a]. CS1/CS2 [BDN^{+}94, FZ96, LWB^{+}98, War99, Wil95b]. CS199 [Mos98]. CS2 [BDN^{+}94, BT95, Col91b, CS99, Den96, FZ96, FS90, GG99, GPJ95, HCL96, JPKPM99, Joy90, Kel93, LWB^{+}98, MBW^{+}98, McN98, Med98, MS90, OD94, Pro97, Pro99, Sab90, SG97a, TBF98, TZ99, Wal94, War99, Wil95b]. CSAB-accredited [MM97]. CSAC [MM97]. CSAC/CSAB [MM97]. CSAC/CSAB-accredited [MM97]. CSCW [CLL^{+}96]. CSedRes [AW93, AWB94]. CSI [Thw94]. CSP [Nev99, OM95, Ols93]. CSTC [KG99, Kno99b]. Cuarto [SLM98]. culminating [MMB^{+}91]. culture [Ber97a, BCPV98, FMM97, Rou92]. curbing [Dad90]. cured [GSW^{+}98]. Current [DM92, NC99b, SG99a, SG99b, SRL^{+}99, JBCL98]. Curricula [GL93, PRS93, AMN98, Alm91, AMBC98, AJ93, BMB^{+}95, Ber97b, GL94, Gor98b, HBC97, HMBO98, HS97, HWS92, KI93, Lim98, MVS^{+}98, NW92b, OBPB97, PH97, Rub92, SL94, Yah92]. curricular [BN98]. Curriculum [CEK^{+}99, LGB^{+}99, MR99, RSLD99, ZBD99, AMN98, Ack99, BBC^{+}94, BSM^{+}94, BKLV98, Car90, CFA^{+}94, CHLS95, Cle97a, Cle98a, CKM^{+}97, CDF^{+}97, Cul97, DDP96, DI98, DM92, EL97, FK92, FB97, FPC96, For91b, GY97a, GCD^{+}94, GDC^{+}97, Gor99b, GJMS90, Got99a, GL96, GL97, Had98a, HH95b, Has94, HT97, Hil90, HHK95, HANWN93, IG97, JPS90, JMM97, Joh94, Joh97, Joh99b, JRS96, Kay96, KPW94, Kub99, KGZ99, KKA98, KAL^{+}96, Lee96, Lee97, Lee98a, LGK^{+}93, LM98c, MHP93, Mar99c, Maz93, MAD^{+}95, MJM96, McC92, MM99, Mer92, MBE^{+}93, MV98, MDG^{+}95, PMB98, PP95, PZ96, PRS93, PRS94, Pre96, RF96, RFH^{+}93, RS91, RS94, RC97, SH95, SW93, SBK94, SWWB97, Ste94, Tay99b, TG94, Tew95, Thu90, Tol95a, Tol98, TW94, UD97, WCDT95, Wal98c, WG93, WGD^{+}95]. curriculum [Wil95b, WUGV99, ZBK^{+}90, Lit90]. curriculum-cycle [MHP93]. curriculums [GF96]. curve [ZLS98]. cycle [Con98a, HCL96, MHP93]. CYK [HW91].

D [CH98]. Dancing [BEG^{+}96]. DAPPLE [Kot95]. Data [CHLS95, HS93, Jar94, SHY96, BBG^{+}99, Bie93, BS99, CK99, DMW^{+}97, Eps93, Fel90, GGT98, GPJ95, GT99b, Gre90, GZ96, GMAB96, HSS98, Hag91, Hai96, Hug90, Kea93, Koo95, KCA99, Kot95, Kro96a, Kro96b, LN91, Lev97, Lev99a, Lev93, LAH94, MC90a, McN96, MST97, MM94b, Ngu98, NW99a, PR98a, Pre99, SH95, SRY98, Sha91, Sil91, Sil94, Smi91, Smi94b, TGGP97, Tol95b, Uck91, WCC^{+}92, Wei97, Yer94]. data-parallel [Kot95]. Database [DU96, TW92, Ape90, Asl98, Cle98a, DEP97, Din91, Goe96, Lee90, LH92, LHA90, Pig92, RR97, SF91, UD97, YLQ98]. database-part [Cle98a]. databases [DUM98]. dataflow [Can91, Lan93]. DATATUTOR [Hai96]. day [Hec94]. DB [Cun96]. DB/IR [Cun96]. DBMS [Gui99]. DBTool [LH92]. DC [ACMJ90]. DD [JPPMMF96]. DD-Mod [JPPMMF96]. dead [LM98a]. deaf [SH99c, Wan98]. Deakin [DG99]. Dealing [Lev93]. debate [Rob95b]. Debating [Sta93]. Decision [Tol95a]. declarations [Hea95]. declarative [Hei93]. Deconstructing [Mar98b]. decoupling [NW99a]. dedicated [Bou96]. deductive [Uck91]. default [Ant97]. defined [GK97]. Defining [AFVH97, Mar99b, Tur94]. definition [HT97, Mil93]. degree [Bal99, GHK93, GSW^{+}98, HJ92, KI93, LN97, MM97, MM98, NI98, WGD^{+}95]. delivered [Kum98]. delivering [Dyc93]. delivery [Nap97a, RM98, SY98, WP98, ZBD99]. demo [JD97b]. demonstrate [Tay97b]. demonstrating [Joh99a, MC90b]. demonstration [Mye94, Rob96, WCC^{+}92, WCMT94]. demonstrations [GNR^{+}95]. denotational [KOC91, Slo91]. Departmental [Coh99, BK92b]. departments [KCC^{+}98, MM97, TM92]. depth [HL94, SL94]. depth-first [HL94]. Derivation [DNPR94, Den96]. Describing [WBBW94]. description [Tre98]. descriptive [MNC99]. desert [Dal90]. Design [AMBC98, BKCG98, CAH90, CDVF97, Ell96, KMO95, MdL92, Ngu98, Pil93, Pre99, Ras97a, AMN98, Ada93b, Ada93a, Ada96c, BM96, BYC98, Bar93b, Bar96b, BNS^{+}97, BBG^{+}96, Ber93, CH98, CGJ93, DN97, Dow99, FPR98, FF94, GGT98, GSvD99, Had99, HWB92, Hol98b, Kat94, KK98b, KH92, Kno97b, KOC91, Lar97, LP98, Lev99c, Li98, LH92, LM94, Mat95, McC99c, Nef99, NW92a, NW92b, Osb92b, Par95, PR92, Pow99, Pre94, Pro98, Pro99, Sch97, Sil90, Smi91, Sty94, Tan92, TSG98, Tol98, WP98, Wal99d, YJ95, ZLS98, ZZLS99]. designed [FR96b, Ful91, Kro92, KCPC98]. Designer [Pow99]. Designing [AAW92, Coo98, Fal99, FG96, KJ97, Mit96, Reb98, RP98, Sch92, Arm99, Cla98b, FZ96, MNC99]. desk [Mot96]. Desktop [Mot96]. details [Hea95, STM95]. detecting [GT99a]. Detection [Wis92, Wis96]. deterministic [Höf91]. develop [Cle98a, CWS98]. Developing [CM96, GKC^{+}98, Haf94, KG99, Kou97, MPTW92, MM93, Sai97, TLR94, YLQ98, Ciz97, EHS^{+}99, LK95, Mar99c]. Development [RSBAS98, BN98, Cun94, DH96c, Ell98a, FC99a, HSS98, HBC97, HML^{+}99, Kat94, KR96b, Kub99, KMO95, KAL^{+}96, LG96, Lea97, Lim98, LWC96, LSS^{+}96, MPTW92, McC91, MRA^{+}98, MR98, MMK98, Nac98, OD94, PP95, PM95, PCLM97, PLG97, Seb97, TP93, Tom91, VI94, VIRF96, WB99, WUGV99, Yah92, ZC99]. developments [Eps95, LGBM99]. devices [Eck90, VS92]. diagram [HL98, Pow99]. diagrams [Che94b, Cla93, Kum96]. did [Kol98]. Didactic [Tri99]. Didactics [DS99]. DIDAGRAPH [DS98]. didn’t [SG96]. difference [Nap97b]. differences [CT94, Coh99, HTW98]. different [Kin98, Lev93, MS98, NB99, Tro91, YSKW95]. difficult [Mos97]. Digital [Fag94, SRJ99, Cas98b, GKC^{+}98, Hed99, IMAM98, ME92, NW92b, PR98b, GPP99]. Dilemma [Res96, BW96]. dimensional [GEC96]. Dining [Gin90, bY91, Fel92]. diploma [SK94]. directed [Bon99, Tur94]. direction [SRL^{+}99]. directions [JBCL98, TW94, WRRT97, WUGV99]. Disabilities [VS92, Gro97a]. disciplinary [NW92b, RSB98]. discipline [MCMHS90, Mil93]. disciplines [AJ93, Cle97a, EFGR94]. Discovery [Bal96a, Zie99]. Discrete [Bro93a, Hen90, Ber97b, Bri93, Fle93, Hei93, SW93, Wai92, War95]. discussion [Ber92a, CDF^{+}97, HS98]. discussions [Cla98a]. disk [FH91, WT91]. disks [Mye94]. display [Sig90]. disseminating [Utt99]. Distance [BGEL^{+}97, Bou97, Dal99c, Dav98, HG99, TCP^{+}96, AMN98, BS99, CB96, Car97, DH97, DP90, Jon96a, Law97b, LP99, MH96b, Oli94, PS98, RFPA97, VW96b, WPI98, dC99]. distance-learning [BS99]. distance-students [VW96b]. Distributed [BA97, KPT99, Ste94, And90b, Arn95, BAS99, Bro93b, CBPW96, CL98, Ele96, Gha91, JPPMMF96, JTW90, KPT98, McD96, PR92, Tol98]. distribution [KMH91]. diversified [KK98a]. diversity [AAK^{+}94, Ber97a, Cle99b, DFZ^{+}95, DFH^{+}98]. Division [Cas98c, Har99c, Kay98a]. DL [Dav99]. Do [Fre94, Lev99c, Tja98, AIM^{+}95, CR98, KACL94, MM98, MIG^{+}94, PRF96, Sie99, Tra98, ULW99]. Documentation [MJM96, Jac96, Pig92]. Docware [DFF99]. Does [LN98, Mak97, GSW^{+}98, How93, Rob99b, Roz94, SRY98]. dogs [Wil95a]. Doing [HT97]. domain [Bur97, HHK95]. domestication [RHP90]. dominated [BCPV98]. Domino [Cle99a]. done [Mos97]. DORITOS [OM99]. Down [Bre92, DH93, Hil93, LZ98, Ree95, Scr91]. DPLab [BAS99]. DPMA [WGD^{+}95, CDF^{+}97]. drain [BG97]. drawing [KLMS95]. driven [WD99]. drivers [YLQ98]. DSTutor [KOC91]. DUE [Cas98c]. dummy [FC99a]. Dutch [Yer94]. DYNALAB [BBG^{+}95, BEG^{+}96]. Dynamic [Hec94, Sig90, BBG^{+}95, JPPFPMVI97, Lia95, NS94, Zai96]. dynamics [BD99, STBC94]. DYNAMOD [Ros91a]. e-mail [NSTT96]. early [Ast98b, BK92a, Etl90, Hai97, MJ99, Pat93, RSB98]. Easing [LC96b]. easy [Car95, HP93, NS94, ZZLS99, KD90]. easy-to-understand [HP93]. EASY/VI [KD90]. Eating [Cle99b]. EBNF [Pat94]. ed [HKH^{+}93]. editing [Gam95, GBB^{+}90]. editor [Osb91]. Educating [Oml99, Wil97]. Education [ACM98, ACMJ90, BHH^{+}99, BG94, Car98a, Cas98c, Dal91b, Dal92b, FR96a, HTW97, IAK96, Imp99a, JLP99, KLY93, LWMG95, LPJI98, Mod91, PC98, PN99, Sai93, SHY96, SR99, SH97, WEKM97, YD99, AFVH97, AP91, AW93, ABW93, AWB94, Alm98, Ama90, BGEL^{+}97, Bei96, BA98, BD97, BGGR98, BGG^{+}99, CH98, Cam94, CGK^{+}96, CB96, Cas98a, Cas98b, CDD^{+}97, Cle99c, CD95, Con98b, CDVF97, Dal99a, Dav98, FC99a, For94, Fra98, GC98, Gro97b, GCDP94, Har97, Har98b, Har99c, HHK95, HKSD97, Jon96a, Jon96b, Kle99a, KPW94, KWJ^{+}96, Kot95, KDE^{+}93, LK98, LP99, LSS^{+}96, MZS^{+}95, McC97, MF98, MS97, MH96b, MHB^{+}98, MPHL95, MLS97, Nac98, NB99, OQ98, OK95, Owe96, PBEC93, Ras97a, RM98, Ros99, Rou92, SBB^{+}97, Ske93, SRJ99, SPV99, Taf91, Tay96, TCP^{+}96, Tho98b, TLNS98]. education [Tow98, VW96a, WRRT97, WC94, YJ99, MW99a]. education-quo [GC98]. Educational [Har98b, NSS98, O’D98, RH93, BMP96, BNS^{+}97, DEP97, GL93, Gsc94, KPT98, KPT99, Ola91, RS97, Siv96, Tay99a, WP98, WP97, ZMS95]. educators [FACW90, McC98a, MHN98]. Effective [JL98, CBH97, JMM97, Juh99, Kay98b, MAD^{+}95, Mer90, PS97, PLG97, dC99]. effectiveness [KDE^{+}93, TG94]. effects [BJB98, Mor92, TS98]. Efficiency [Gin96]. efficient [DAUZ94, FJM98]. efforts [Cas98b]. Eiffel [Sch97]. eight [Fel90, Gra93, Mye98, Gra93]. eight-minute [Mye98]. eight-year [Fel90]. Electronic [Cla98a, FBR^{+}99, FCB^{+}99, JL98, Mos98, RR96]. electronically [Tho98a]. electronics [GS96]. elementary [FA98, FP95]. Elements [PR92]. elsewhere [FHL^{+}97]. em88110 [VdlFCC97]. email [AC96]. Embedding [RS91]. emergency [Pro99]. emerging [HJ92]. emphasis [Eva96, TA96]. emphasising [Oml99]. emphasize [GS99]. Emphasizing [Bar96b, Dyc93, RF96, Sul93]. Empirical [FG92, RR99, Cla98b, Dal98b, HH95c, PP97, Tja98]. employers [Pfe99]. emulating [VdlFCC97]. emulator [dlFCC99]. enabled [MLS97]. encouraging [WBC^{+}98]. Encryption [KKL99]. endearing [BEG^{+}96]. endorsed [Bal99]. engagement [Mak97]. Engaging [MM94a]. engine [Col91a, WNKK98]. Engineering [BHH^{+}99, Bar97a, Mod91, Sai93, BMP96, BG97, Bri91, Bry99, Cli91, CKK94, Con98a, DFF99, DDG97, DUM98, DA97, Dru98, Ewi93, Fel92, For91a, For91b, For94, GYT^{+}93, Ger94, GG99, GP91, Gro97b, HMBO98, HK96, JMM97, Joh97, Joh99b, KLE92, KI93, KP90, LK98, Leo91, LN97, MAG^{+}94c, MAD^{+}95, MJ99, MHB^{+}98, MC90b, MM94b, Mur91, ML90, Nak98, NB99, Oli94, Osb95, PBA98, PW99, Poo91, Ree97, Rif94b, RS94, RB92, Rou92, Sap91, Seb97, Ske93, Sob96, Son96, SS97, Tew95, TD90, TA96, TLR94, VW96a, Wer90, Wol98, ZJEP95, Zac97a]. engineers [Lam95]. English [Che94a]. enhance [CD95, JRS96, Roz94, Tay99b]. enhanced [Don95, MLS97]. enhancement [AIW^{+}96, HK96]. Enhances [Bar97a]. Enhancing [Fek95, HGPJ^{+}96, Kno97a, SH99c, RG98a]. enough [Gra93]. enrollments [LEB^{+}90]. Enterprise [MLS97]. entities [Ful94]. entity [Ape99, HG98]. entity-relationship [Ape99]. entry [Bru94, ZJEP95]. entry-level [Bru94, ZJEP95]. environment [Bal97, BK95, BHPG99, BAS99, Ber96, BCHB94, CGP90, CB94a, CLL^{+}96, Car90, CWI96, Dal98c, DT98, Der93, Ell98a, Ell98b, Eri96, FC94, FF94, FR96b, GS99, GVLZ94, GBS98, HG98, HKH^{+}93, HBC97, HML^{+}99, HMS93, KR96b, LM98b, LA96, MHP93, MdL92, McD91, MK97, Mos97, NB98, Nef93, OK95, PMB98, RST97, Rif94a, RBG^{+}97, Ros99, SV96, SSH98, SH99b, SH99c, Tom91, Tur94, VL90, VI96, WBCS95]. environmental [PH96]. environments [CGK^{+}96, Coo98, Eze90, GBB^{+}90, GCDP94, Hou97, MPW99, MLS97, WP98]. EOS [Eri96]. epistemic [She95]. Epsilon [Rog90]. equality [Joh99a]. equipped [Bal97]. ERDs [Tre99]. error [Got97]. errors [BHPG99, Sch95]. essay [FJM98]. essay-type [FJM98]. Essential [GSH96, Ast94, AMBC98, Gre96b]. esteem [Tea92]. Estimating [Col91b]. eText [Rif94a]. Ethercom [RRR93]. Ethernet [AAHM99, Bar93a]. Ethical [MSRT95, MRA^{+}98, Got98a, GL97, Lee99a, MR98, SG96, SG97a, TW97]. ethicaldebates [Tow99]. Ethics [Hou97, AR99, Eps95, GJMS90, Got98b, Got99a, Got99b, KG94, Lel94, Mar97, Mar98b, Mar99c, PH96, Rob98, SVP93, Wah99, Wer97]. etiology [FHL^{+}97]. European [O’D98]. evaluate [CK98, Cun97, Lea95, Sil91]. Evaluating [Fle93, Hei99, Kum98, KDE^{+}93, WP98, BD99, Fek96, HSM97, Rob93, Sze99]. Evaluation [ADB^{+}96, Hos96b, Joh99b, GY99a, Had98b, Hos96a, JPPFPMVI96, Kin98, KSW98, ML90, Pig92, PP97, PR98b, RSBAS98, TBF97]. evaluations [AP91]. event [WCMT94, WD99]. event-driven [WD99]. ever [MHB^{+}98]. Every [BMR^{+}95]. everyone [Nar91]. Everything [MHB^{+}98, KG94]. Evolution [Mur91, Kin92, Tay97a]. exam [Gre96a, KS94, SRL^{+}99]. examination [FJM98, Nul98, WB92, dM98]. examinations [AB99, MW98a, WM98]. examined [Gra93]. Examining [Gre97]. example [BG93, Blu96, CH93a, Dix93, Gre96b, Joh98, Pat90, Pet98a, Pet98b, Pre94]. examples [DBKL90, Tam92, Wal99a]. Excel [RST97, DST98]. excellence [FZ96]. exchanges [SDJ^{+}94]. excitement [Joh98]. exclusion [BC96]. ExCon [Nul98]. Executing [Slo95]. execution [Col91b, McD91]. exemplar [SPSS96]. exercise [Col91b, Gib92, Pio92, Slo91, Sze99, TP94, Tro94b]. Exercises [Tre99, Coo96b, FP97, Fle93, dOG93, Mit96, PFR^{+}94]. existing [PS99]. exits [BA96a, Rob95b]. exodus [LWBS99]. expectations [Bro92]. Experience [GdLC94, Gui99, Har90, Har92, Hay98, Ros91a, Sob96, Str91, Bal96b, Ber91, BABT97, Ciz97, DGEST96, Dor97, Ell96, FP98, FMM97, HP95, How93, Kub99, LM91, LEDW93, MMB^{+}91, Poo91, Rob93, RLR95, SH96, SR99, VB98, War95]. experienced [Bru94, LM98b]. Experiences [Mor95, Owe90, SK95, Wei97, AGT98, Der93, GM91, GY97a, GL93, GBB^{+}90, KS94, Lea97, LC96b, MWO^{+}96, MJ99, NRR^{+}90, OBH^{+}95, SLM98, Tay96, Wal98b]. Experiencing [EO95]. experiential [Cla98b]. experiment [BMC95, Bob99, Gol96a, Gol97a, HT98, LLP95, SW99, Thw94, WCMT94]. experimental [BBF^{+}94, Dow99, Epp91, Gha91]. Experimenting [Rev98]. experiments [BK92a, KKC90]. Expert [WR90, Har94a, MPTW92, Pil95, RSBAS98, Tao95, bY90]. expertise [MPTW92, RB92]. explain [JF97]. Explaining [BT96, Hay95]. explanation [JD97a]. Explication [DUM98]. Exploiting [MSSZ97]. exploration [BHPG99, NS96, RR99]. explore [Shu90]. Exploring [BCPV98, FP97, Llo94]. Extended [MCH98, SR91]. Extending [AC96, Zie98]. extensibility [Dix93]. extensible [NS94]. extension [MLS97]. external [DG99]. eyes [LM98a].

fable [Sch91]. face [TH98]. Facilitating [Bar97a, VL90]. facilities [FF94]. factor [Oml99]. faculty [AIW^{+}96, GS96, WBC^{+}98]. failing [SH98]. fails [MW99c]. failure [DH96c]. fair [FH96, GPP99]. fairer [SAL^{+}90]. fall [Sie99]. fantasy [Mos97]. fashioned [Gor99a]. FASRS [JD97b]. FASRS-demo [JD97b]. Faster [SAL^{+}90]. favorite [And90a]. FC [Dav90]. FCLUST [EKR98]. fears [Bry99]. feast [Fel92]. features [Car93, DST98, KW99]. featuring [BBG^{+}95]. February [ACM98, IAK96, KLY93, LPJI98, WEKM97]. feedback [LH99, MW99b, PP97]. Felder [HCL96]. female [Coh99, LB96, Mos93, Tea92, Tow96]. few [MIG^{+}94]. Fibonacci [MSPF95]. field [SP96]. fields [ABM93]. file [McD96, Ros91b]. files [FH91]. Final [AJ93, KI93, DH96b, Tea92]. Finals [NPP^{+}93, Med98]. Finding [Bal97, Wal99a, Ast92]. findings [MV98]. fingertips [Ada96a]. finished [GY97b]. Finite [MD94, Han92, RHND99]. finite-state [Han92]. Finland [JPT^{+}99]. First [Hos96b, Ric98, ABFM90, And90a, BMC95, BBC^{+}94, BFG^{+}94, BW96, BD93, CBH97, Dor97, EHS^{+}99, Gol97a, Gol97b, GMAB96, Had98b, HSM97, HH95a, HH95c, Hos96a, HL94, Kay92b, KKR95, KR96b, KR99, LP98, Lee99a, LWBS99, MH96c, MV98, MM94b, Moo93, MRS93, Noy93, PMB98, Pat94, PRS93, PH93, Ree96b, Rei95b, Rub92, Sca96, SL94, SW91, Smi94b, Tay98, Tro94a, Wal96, WS98, Wil95b, Wol98]. first-year [Dor97, Kay92b]. five [Ada96c, And90a, BMC95, SK95]. five-phase [Ada96c]. five-year [BMC95]. fixed [Hai97]. fixed-point [Hai97]. Flag [Yer94]. flexible [CBPW96, Zag98]. flies [Kea93]. floating [Chu99]. floating-point [Chu99]. focus [DMW^{+}97, Fie96, LWB^{+}98, PZ96]. fold [BSK94]. fools [Sie99]. force [AJ93, PBEC93, RSLD99]. forest [WBBW94]. Fork [Kum96]. form [Wu92]. Formal [Bru99, DH95b, Wad95, WCDT95, Gam95, Hil96, JRS96, Kur91, LBJ94, MH96c, PP95, RBL^{+}97, Sai97, Sob96, Tre98, Tre99, Tro94a, War95]. Formalism [Sai93]. formalizing [Tre99]. formula [Joh99a]. fostering [Ber97a]. Foundation [Cas98b, Cas98c, Har97, MR99, Har99c].

Foundations [PRF96, RB92]. four [BD93, CJ91, CDF^{+}97, DGEST96, DM92]. four-peg [CJ91]. four-year [DM92]. Fourier [JD97b]. fourth [Wu92]. FP [SR91]. FP-like [SR91]. fraction [Per99]. framework [BKCG98, Had98a, Hol99, MO96, Pat97, Wal94, War99, Wha91]. free [Car99a]. frequency [Lev99b]. fresh [Gor98a]. freshman [Coo96a]. freshmen [BLOS98, DNPR94]. friendly [Mos93]. frosting [Ger94]. fruits [Poo91]. full [BD97]. Fun [SCVI91, CL90]. Function [Ras97b, BV97, GK97, Rob99b]. Functional [CL90, JPPFPMVI96, SCVI91, SR91, SLM98, VI96, Wai92]. Functions [Tro92, Joh99a, LC96a]. fundamental [HCD^{+}98]. fundamentals [Str99]. funded [LM91]. Funding [Cas98c]. funny [Joh99a]. future [BGGR98, BSM^{+}94, CEK^{+}99, HCD^{+}98, IMAM98, JBCL98, Kno98b, LW97, MCd^{+}94b, OBPB97, SRL^{+}99]. futures [MBW^{+}98]. fuzzy [EKR98, Kam92].

GA [ACM98]. Game [KST92, CB96, Mos97, She95, SG97b, TD90]. games [Fal99, Gin95]. gap [Cle98b]. gate [Mey96]. Gateway [CSB93, Bal96b, JKB96]. Gathering [McM96, Utt99, Bar97b]. GATutor [PWST94]. Gdañsk [GS96]. Gender [CJ99, FHL^{+}97, GS93, Gre99, HTW98, CT94, FCR^{+}94]. General [McC98a, LR94, MS97, SPV99, Taf91, Tow98, Whi98]. generality [Náv94]. generalized [KMO95]. generals [Llo94]. generating [Osb91, TS95]. Generation [BKP^{+}99, AHST93, JK91, Wil97]. Generator [KJ92, HDC99, PMCGJP97]. generators [War92]. generic [Cle98a, CS99, Fek95]. Genetic [ES99, BV97, JF97, PWST94]. GENIE [GBB^{+}90]. gentle [MSPF95]. gently [KW99]. Geographic [Cam94]. geometric [Ben92, Ben95, Khu94]. geometry [DBKL90]. Georgia [LPJI98]. gestalt [Zac97b]. Get [CFK^{+}92, Kol98, McK95]. Getting [CDD^{+}97, Pan90, SH90, Wer97, WB93]. girls [RW96]. Giving [VB98]. Global [Dav99, MMKW99]. goal [PM95]. goal-oriented [PM95]. goals [Bro92, Cas98b, Mag94a]. going [AIM^{+}95, CJ99, LGK^{+}93]. gone [AIM^{+}95, EBD^{+}90]. Good [War92, Car99b, Gro98, KWJ^{+}96, SC94, Utt99]. Grading [JU97, Hit99, How94, KSIR94, SAL^{+}90]. Graduate [FRK96, Gor99b, Wol98, Amo93, Car90, Gor98a, Gor99a, NC99b, SK94, SG99a, YJ99, bY90]. grammars [LK95, Pro97]. Grant [Har99b]. Granularity [GSvD99]. graph [Cor99, DS98, JK91, Shi93b]. Graphical [JPPFPMVI96, Wer90, LH92, NP92, PWST94, PFR^{+}94]. Graphics [Bri93, Car99a, Car90, Det93, Dix93, Gei91, GEC96, GKB^{+}95, HCGW99, HL94, HANWN93, Jan97, MCMHS90, Muk99, Owe90, Owe92, Rob95a, RP98, Sch91, Sch92, SA95, Sch90, SW95, SPSS96, SPV99, Tic95, WK99a, WK99b]. graphs [Lia95, Mit96]. grasp [Joh98]. Grinnell [Wal97a]. Group [ACK^{+}96, BBG^{+}96, BN98, BKP^{+}99, Bri91, Ell98a, FBR^{+}99, GK98, KFD^{+}99, LGB^{+}99, MR98, Tho98b, ADB^{+}96, BD99, Cla98b, DU96, Dru98, HGPJ^{+}96, KWJ^{+}96, SH98]. Groups [FBR^{+}99, FCB^{+}99, HWB92, FBR^{+}99].

Groupware [Dru98, MBC^{+}96, McM96, Smi98, TH98]. groupware-computer [TH98]. grows [HCGW99]. growth [Kno98b, Kno99b]. GUI [ML96]. Guided [Wal99b, LSS^{+}96, MCW99]. Guidelines [YR95, GDC^{+}97, KBCP99, KWJ^{+}96, PM96]. Guinea [Bam98].

halting [Mye98]. Handling [SBTC94]. hands [Can91, Hil90, Hus95, JTW90, SR99]. hands-on [Can91, Hus95, JTW90, SR99]. Hanoi [CJ91]. happened [Joh99a]. happens [Car99c]. hard [WT91]. hard-disk [WT91]. hardware [Gsc94, Sea93]. harmful [Fra98, Kay96]. Harnessing [PS97]. Hashing [Tro92]. HCI [HWB92]. help [Bre95, CGP90, McC99b, NF97, Shi93b, Tja98]. helpdesk [DG99]. Here [AHST93, Har99a]. heresy [Pat93]. heterogeneous [Eps93]. heuristic [Shi93b]. hierarchical [BFG^{+}94, Li98]. Hierarchies [Náv94]. hierarchy [CCR99]. High [Bro94, AG96, Bro97, BSM^{+}94, CFK^{+}92, FHS^{+}94, FCR^{+}94, Hay97, MZS^{+}95, MBE^{+}93, RSBAS98, Rif94b, RW96, Sap91, SH95, SW93, Sch96b, Str91, WBC^{+}98, WN96, ZMS95]. high-level [Hay97, Sap91]. high-school [RSBAS98, Str91]. HIPE [VI96]. hiring [MW98b, Rob99a, WMZ^{+}99]. Historical [Kno96, Kat95, Lit90]. History [DI98, Lee96, Lee97, Lee98a]. HND [RSB98]. Holland [HTW98]. Home [CBPW96, DP90]. Home-study [CBPW96]. honor [Rog90]. honors [Kay93, LN91]. hooked [CFK^{+}92]. hooks [Ast98a]. Hopfield [Kea93]. Hospital [Pro99]. hosted [Cul98]. hosting [Str91]. hours [JLH^{+}98]. HTML [MHP98, NF97]. human [Cla98b, Cun91, HWB92, MAK^{+}99, Oml99, Cle99a]. human-computer [Cla98b, Cun91, HWB92]. humans [PM97]. humble [Per99]. Hypercard [Han92, Coe93, DH90, Fri91]. hypergraphs [KLMS95]. Hypermedia [HFKSM98, CGJ93, DA97, Har96, Kot98, MH96a]. HyperTalk [Gol99, KP91]. hypertext [Bar93b, FF94, Mak97, Moo95]. hypothesis [HTW98]. hypothetical [Ber93].

I-CASE [GVLZ94]. I-IV [BBC^{+}94]. I.S [GDC^{+}97]. I.S. [CDF^{+}97]. IBM [Det90]. Iconic [CBH97, CB94a]. ICTs [OQ98]. Identifying [RB92]. IEEE [AMN98]. IEEE-CS [RSLD99]. If [Dal90, Kno98a]. II [Bro93a, DBKL90, EO95, Lee98a, Pli98]. Iliad [Arn91]. illustrations [GSvD99]. illustrative [Ast94]. Image [KKC90, SPSS96, ME92, McN98, SV96]. images [FP97, Gei91]. imbalance [FHL^{+}97]. imbalances [GS93]. immersion [Rab94]. impact [Car97, GP91, Mar99c, Rob92, TM92, Tho98b, TLNS98]. impacts [Tay97b]. imperative [Bre95, Ter95]. imperatives [Cul99]. implement [BA96b]. Implementation [Höf91, Ada93b, BKCG98, Ber90, CAH90, CGJ93, Ell96, Hea95, KOC91, KMO95, Lev99a, LWBS99, MR99, MD99, PM95, RSBAS98, Zac97b]. implementation-biased [LWBS99]. implementations [Ful94, Gur95, Maz93]. implemented [GBS98]. Implementing [KP90, MC90a, PM96, Tay99a]. implications [MV98]. importance [TG94]. important [Wic95]. improve [FR98, LH99, TJ90]. Improved [Wis96, MRS93, SS95]. Improvement [MR99]. improves [Tow96]. Improving [ABM93, Bai91, BD97, Ber97b, BK92b, Der93, MK97, PS99, Coh99, PS98]. in-service [HG99]. inadequate [Kot98]. including [MC90b]. inclusion [Ree96b]. Incoming [BLOS98]. Incorporating [PW99, Ree98, Nap98]. independent [CBPW96]. Indiana [KLY93]. Indianapolis [KLY93]. individualized [Rei90, TS98]. induction [RSB98]. inductive [McC99c]. industrial [Ell96, PH99]. industry [Cli91, Con98b, GYT^{+}93, IBF^{+}94, MCW99, Son96, Wil97]. industry-guided [MCW99]. Inexpensive [Bec94, Tic95, Smi91, VIRF96]. infinite [Sie99]. influence [BCPV98, Kol98]. influences [Tja96]. informally [Wal98b]. Informatics [AG96, Bei96, DA97, GS96, Hol99, Mak97, Mak98]. Information [GL94, GCD^{+}94, Gor98b, Gor99c, LGK^{+}93, LM98c, LGB^{+}99, Zag98, BKS^{+}99, Cam94, CLL^{+}96, Cla98a, CDF^{+}97, Cun96, DR99, Fri91, Gor99b, HSC^{+}90, HH95a, HAG92, HWS92, KBCP99, Kou98, LLP95, MPHL95, MDG^{+}95, MLS97, RF96, Ric98, WGD^{+}95, Wil97]. infrastructure [BBG^{+}95, FK92, Ree96a]. INGRES [Goe96]. inheritance [BA96b, Ber98, BT96, Din97, Ree96b]. initial [AGT98, EW97, FC99a, MWO^{+}96, MVS^{+}98, WH96]. initiative [JBV97]. inner [BR99]. Innovation [Imp99a, YJ99, DJ99, Pet98a, Pet98b]. innovations [IG97]. Inquiry [AGT98, Jul92]. Inquiry-based [AGT98]. insertion [TS95]. insight [SBK94]. insight-based [SBK94]. Inspections [Hil96]. Inspiring [Fel99]. instances [Kor94]. institute [MPW99, SW93]. institution [FP98]. institutional [PS97]. institutions [APNA90, LEB^{+}90]. instructed [NG94]. instruction [AGT98, BNB^{+}98, DT98, EW97, Fek96, Imp99b, Kot98, KGZ99, KCPC98, MWO^{+}96, MCd^{+}94b, Moo95, Mur91, NW92a, SP99]. Instructional [BW97, Mye94, Ast98a, Ciz97, Coo98, Ful91, Joh98, KH92, KD90, McC98b, Nap97a, Tay99b, Tri99]. instructions [Ben95]. instructor [PP97]. instuction [LEB^{+}90]. integer [Ben95, Hig93]. integers [Ben92]. integratable [AMN98]. integrate [GL97]. integrated [CSB93, CL98, DR99, Gro97b, HP94, Joh94, LK98, PRS93, PRS94, RSB98, Sea93, WDM97, YLQ98, ZC99]. Integrating [Bei96, BKS^{+}99, CLL^{+}96, DDP96, Gam95, GL96, Hil90, Joh97, Kno97b, LGB^{+}99, MVS^{+}98, MW98a, NI98, Owe96, Pes91, Rod96, SA95, SL94, Ske93, Tay96, Uck91, UD97, Wil95b, WC94, WD99, YJ95, AC96, BMP96, CC92, CP96, Fek93, FF94, GJMS90, Had98a, Has94, Sob96]. Integration [FC99a, Got99a, Joh92, SW93, TBF98, JMM97, Li98, Mar99c, MAD^{+}95, TW92]. Integrative [GPJ95]. Intel [Det90]. intellectual [LWB^{+}98, YD99]. Intelligence [WV91, AAW92, CH93a, KM98, LaR93, LR94, McC99b, Pax95, Stu94, TW92, TP94, Tos93]. Intelligent [Rug99, JD97a, LM91, RM98, WNKK98]. intensity [BG97]. intensive [Ada93a, SF91]. Inter [DN97, BK92b, PS97, RSB98]. Inter-class [DN97]. inter-departmental [BK92b]. inter-disciplinary [RSB98]. inter/intra [PS97]. inter/intra-institutional [PS97]. Interact [GR99]. Interaction [CHCL99, Ash97, AR98, Cla98b, HWB92, LM98a, MAK^{+}99, WBC^{+}98, Cle99a]. Interactive [ACK^{+}96, JKB96, MWO^{+}96, MH96a, RM96, BKCG98, BHH^{+}90, BJR94, BGG^{+}99, CGK^{+}96, CWI96, Coe93, CDVF97, CBPW96, CSB93, GSvD99, Gri99, HGPJ^{+}96, Jan97, JPKPM99, Nap97a, PMB98, Rif94a, Rod95, RBG^{+}97, Sch92, TGGP97, Wan98, WDM97, WK99a, WK99b]. Interdisciplinary [Sch96a, Ber98, GZ94, ME92, Sch91, Sch99, TP93]. interest [McC98a]. interesting [Sha98, Wal99a]. interface [Bar93b, Cun91, Gur95, KK98b, YB99]. interfaces [Wer90]. interim [MW98b, RSLD99]. internal [CD95]. internals [Shu90]. International [JBV97, GPP99, SDJ^{+}94, Wol98, ZBD99]. internationalise [Cle98a]. internationalizing [Kou98]. Internet [Cun96, MS97, Sch96a, Ack99, Bei96, BN98, Blu96, BGGR98, Car97, Car98b, Coo98, Cun97, Ell98a, Gur98, HGPJ^{+}96, KAK93, Lea97, Mar98a, MRA^{+}98, MR98, MP96, Mul91, PL96, PSZ99b, Ros99, SBD99, TCP^{+}96]. Internet-accessible [Cun96]. Internet-based [SBD99]. internship [CD95]. interpersonal [SBTC94]. interpreter [BK95, HP93, JPPM96, Lam92]. interpreter-based [BK95]. interpreter/visualizer [JPPM96]. interpreters [Bru99]. interpreting [Fek96]. interrupt [Hol97]. intervention [MW99c]. interviews [Tea92]. intra-institutional [PS97]. intracorporate [VL90]. Intractability [Khu94]. IntraED [MW99a]. intranet [MW99a, Nak98]. intranet-based [Nak98]. introduce [Jef91, Sab90, Wal98b, WGHL97]. introduced [FB97]. Introducing [Cam96, DH95b, Duc94, Hai97, Hol94, Kam92, MBJS90, Mer92, Rab94, Sap91, VS92, Wai92, Bro97, ES99, Pil95, Tem91]. Introduction [BHH^{+}90, Ada96c, BSK94, CB94b, Cun95, Dix91, Gei91, Gui99, Hay97, Hol98a, JTW90, MSPF95, Mil96, Moe95, She98, STM95, Tol95a]. Introductory [BKP^{+}99, CH93b, MS97, Wit96, Abu92, Ack99, APNA90, Ast98b, BMC95, Bal90, BMB^{+}95, Bar92, BABT97, Bru94, CCR^{+}97, Car97, CK99, Cha91b, CH91, Coo96b, CKM^{+}97, Dal91a, Dal99d, DL95, Fek93, FPR98, FW96, FR96b, Gri99, HS98, HP94, HP95, Hil93, HCGW99, IHA^{+}94, Joh98, Joh92, JRS96, Kay98b, KCC^{+}98, KSW98, Kno96, Kor94, Lam95, LBJ94, LZ98, LH92, Loc91, McC99a, MD94, MH96b, MHP98, Moo93, Mor92, Mot91, OK95, Pio92, RPF94, Ree96a, RFH^{+}93, RS94, RLR95, Ros99, RT98, SS94, SP96, SA95, Sch90, SH91, SS97, SBD97, TW94, WR97, WT91, WSK97, WFGW94, WWDM96, WH96, Yer94, Zac94, Zie99]. invariants [Arn94, Ast91, Gin95, Tam92, Wal98b]. invented [Pli98]. Inventing [MN90]. investigation [MV95, Moo93, PS99]. IR [Cun96]. Irish [Con98b]. ISC’98 [LM98c]. ISETL [BHH^{+}90]. island [Dal90]. Isomorphism [CP94]. Issues [ICD^{+}90, JCF^{+}93, LGB^{+}99, FC99a, GL97, Maz93, MAG^{+}94c, MSRT95, MRA^{+}98, MR98, PH96, Pro99, Ras97a, RF96, SG96, SG97a, SBTC94, SG99a, SG99b, TW97, YSKW95]. iteration [Hai97]. Iterative [Seb97, HK96]. iterator [Ras97c, Zie98]. ITICSE [Imp99a]. ITICSE-99 [Imp99a]. ITiCSE’99 [BKP^{+}99, FBR^{+}99, KFD^{+}99, LGB^{+}99]. IV [BBC^{+}94].

JAVA [YLQ98, Arm99, BA97, BN98, BNB^{+}98, BT98, BGGR98, CH98, CWS98, CS99, Cul97, DV98, EHS^{+}99, GG99, GBS98, Had98a, Had98b, Har98a, Hos96b, Hos96a, JPT^{+}99, KSW98, KW99, KR99, Muk99, Nap97b, Nap98, Nev99, RP98, RHND99, SBB^{+}97, Sch97, SPV99, Wei97, WSK97, WK99a, WK99b, YB99]. Java-based [RHND99]. JavaScript [MHP98, WS98]. JAWAA [PR98a]. JDBC [YLQ98]. JDuck [GG99]. JFLAP [GR99, RG98b]. job [Ast92, Tea92]. joint [CDF^{+}97, HH95a, WGD^{+}95]. Jose [WEKM97]. journal [RG98a]. journals [RG98a]. joy [WDH96]. Joyce [Ele96, McD92]. Joyce/Linda [Ele96]. joysticks [SBB^{+}99]. judgements [PH95]. July [Car98a, FR96a, Imp99a]. June [Imp99a].

K-12 [PBEC93]. Kansas [Dal92b]. keep [AIM^{+}95]. keeping [Bru94]. kernel [MK98]. killer [Eps95]. kindergarten [BNS^{+}97]. KLYDE [BBF^{+}94]. know [KG94, MHB^{+}98]. Knowing [Ada96b]. knowledge [Hol99, Kol98, MC90b, PSZ99a, Rug99, WNKK98]. Kolb [HCL96]. KOMBIZ97 [SW99]. KPPCDL [Ros99].

lab [Bal97, Coo96b, Gei94, GS96, Har90, Har96, Joy90, Kno96, Kno97a, MD94, Ree90, SW99, SBB^{+}99, Thw94, Tja98]. lab-rich [Bal97]. Laboratories [GL93, BK92a, BW97, CSB93, FG96, HH95b, JRS96, KLE92, KWJ^{+}96, Kno97c, LS93, LWC96, MS97, MAK^{+}99, Nap90, Pre96, RS91, RS94, Tur94, Wal94]. Laboratory [Bob99, Hus95, Kur91, MR99, PA90, AAW92, BBG^{+}95, CGP90, Cha91b, Col91b, CM96, DMW^{+}97, DH96c, Ele96, Epp91, Epp92, Eri96, EO95, Gro97a, Hei93, JY95, Kat94, KJ97, KM98, Lev97, Lev99a, MK98, MK99, MB95, McC91, McN98, MH91, NW92a, OD94, Ols93, PM95, Pil93, PFR^{+}94, Rei92, RBG^{+}97, Sap91, SR99, Smi91, Smi94b, Tan90, Tay96, Tol95b, Tur94, WT91, Wen90, WB92]. Laboratory-style [PA90]. labs [Bal96b, Bro92, CS99, EL94, HA95, JKB96, PD96]. Language [Hos96b, Bar95, BK95, Bre95, BW96, Car93, Coe93, Det90, FF94, FS90, GF96, Had98b, Har92, HH95c, Hol94, Hos96a, JCF^{+}93, KW99, KKR95, KR96a, KJ92, LC96a, Lev95, McD91, Nef99, PH93, Rad90, RL95, SW91, SR91, Tem92, WS98, Wil90]. languages [BF94, Bru99, BP95, EN90, dOG93, Hay98, Kin92, Kur91, Max93, McC92, Mer90, Mes91, MN93, NS96, Pan90, RL95, RBL^{+}97, Van98, Yea91]. languages-comparatively [Max93]. Large [Kay98b, KCC^{+}98, CS98, CCR^{+}97, Har94a, Imp99b, LEB^{+}90, MC90b, PS98, ULW99]. large-scale [PS98]. largest [DP90]. later [SK95]. Latest [Eps95]. law [Kay92a]. layer [GZ96]. leading [OCH^{+}94]. learn [Bre95, KV94, Tho98a, YDM99]. learned [AIW^{+}96, BNS^{+}97, DL95]. Learning [Cas98a, FBR^{+}99, FCB^{+}99, HWB92, HKSD97, Mit98, MD99, Ros96, TH98, Abu92, AMN98, ASW97, Bal96a, BAK99, BS99, Bou97, BWW96, CB94a, CLL^{+}96, CGK^{+}96, Car97, CHL96, CK99, Cle99a, CKM^{+}97, CSB93, Dal98c, DH97, DT98, DP90, DJ99, DeC96, DAUZ94, DR99, Der93, DU96, Ell98a, ES99, Fal99, FW96, GY98b, Gol96a, Gol97a, GBS98, Gre96b, Gre97, HG99, HG98, HP95, HCL96, Hus95, Jan97, JPT^{+}99, JKB96, Kir96, LLP95, Law97b, LC96a, LN98, LM98a, LM98b, LH99, Mag94b, Mak97, Mar99b, McC96, MPW99, Moo95, MS98, Mos97, MPHL95, MLS97, PMB98, Pre96, Ric98, Rif94a, RFPA97, RBG^{+}97, SS94, SP99, SH99b, SH99c, Sta97, Tay99b, Tja96, ULW99, Utt99, Wal97a, Wan98, WPI98, Wil97, WFGW94, WDB98, YR95, Zag98, Zie99, ZC99]. LearningWorks [Gol98]. Lecture [NSTT96, Ell98b, MCH98, Rod95, WCMT94]. lecture/tutorial [Ell98b]. lecturer [BP94]. Lectureroom [WCC^{+}92]. lecturers [Tra98]. lectures [Bal97, Mak98]. led [DJ99]. legacy [Mor95]. lemmas [JM90]. length [Tro91]. lengths [Cha91a]. less [LN97]. lesson [DL95]. Lessons [AIW^{+}96, WM98]. Let [Tem91]. level [Bru94, FB97, Hay97, KLL^{+}98, KBCP99, KCPC98, McD96, Reb98, SK98, Sap91, SCF^{+}93, SH91, Str91, ZJEP95]. Levels [Kel93, Kin98, Lev93]. liabilities [Tow98]. liberal [APNA90, Arn91, Arn94, AAK^{+}94, Gol99, Has94, KH94, Pra92, SSH98, TLR94, WBCS95]. libraries [GPJ95, Hed99, Par97, TG94]. Library [CS99, Arn95, Cas98b, GKC^{+}98, IMAM98, JPPMMF96, Kot95, LN98, Rob95a, RP98]. Licensing [Wer99, Got98a]. life [Con98a, Pat97, Son96]. life-cycle [Con98a]. lightweight [Ber96]. like [Dal90, MM98, SR91, VI96]. limited [HS91, WB99]. Linda [Ele96, McD92]. Line [PC98, AB99, Car99b, Gol96b, Gre96a, GK98, JL98, Kno97c, Kum99, Med98, PS99, RG98a, RT98, WM98]. lingual [HBC97]. link [GZ96, Tay99b, Ber97b]. linked [Kho90, KMH91, MS90, Wha91, WLC96, Zai96]. linking [NF97, RL93, Tay99b]. links [McC97]. LISP [SV96, Mer90]. list [Hit99, Kho90, KMH91]. lists [MS90, Tro91, WLC96, Zai96]. literacy [Arn91, EL94, Fag94, GBL94, LW97, TW97]. Literate [MM94a, Wit96, SC94]. literature [Che94a, Cun95]. LLparse [BJR94]. Local [Yah92, HG99, VW96a, VW96b]. locker [JPPFPMVI97]. Logic [SH95, Wad95, BKLV98, GSK^{+}95, Hei93, Mey96, Mye90, Nef93, Pow99, Sch95, Slo95, Tre99, Uck91]. logical [Ful94]. long [Sha99, TZ99]. LongNbrs [Sha98]. look [DeC96, Gor98a, MM98, Ste91]. Loop [Gin95, Rob95b, Arn91, Arn94, Höf91, Tam92, Wal98b]. loops [BA96a, Col91a, Sie99]. lost [Dal90]. lot [McC99b]. Lotus [BJB98, Cle99a]. Louisiana [PN99, JLP99]. low [Hug90, Moe95]. low-cost [Hug90]. low-tech [Moe95]. lower [FB97]. lower-level [FB97]. LRparse [BJR94]. LUPSort [GA91].

MA [McM96]. machine [Abu92, Bar95, Bro97, Don95, ES99, Liv97, MD94, SH91, Slo95, YJ95]. Machinery [Bro94]. Machines [PH97, Pli98, Ågr99, Han92, Kro92, RHND99, Win93]. Macintosh [Ful91, Ful92]. made [dOG94, KKL99, ZZLS99]. Mail [PH93, BCHB94, NSTT96]. maintenance [McK95]. major [HW90, Rad90, ULW99, WBCS95]. majors [BS95, BBD^{+}95, HL94, Joy98, Kay93, Kay98a, LR94, PM96, Rei95b, SVP93, SS97, Tic95, Wal99d, Zac97a]. make [Rob99b]. makes [Nap97b]. Making [FH91, Ara91, RBL^{+}97, RK97]. male [BCPV98]. managed [Ros97]. Management [HWS92, BG97, Bou97, Ele99, HMOS90, Kay98b, LHA90, NRR^{+}90, Pat96, Ske93, Sul93]. Managing [CCR^{+}97, HSC^{+}90, LG96, WT91, Pat96, Siv96, SSN99]. manifesto [Ola91]. Manpower [HAG92]. manual [Har96]. manuals [Mes91]. many [Rob99b]. mapping [CHCL99]. March [ACM98, BG94, Dal91b, Dal92b, JLP99, LWMG95, LPJI98, PN99, WEKM97]. mark [MW99b]. mark-up [MW99b]. marker [MGH98]. marking [JL98, MW99b, PS99]. Martian [FP97]. MASPAR [Duc94]. master [Ros91b, GSW^{+}98, MMB^{+}91, SBD97]. mastery [ULW99]. matching [Nar91]. Materials [KFD^{+}99, KJ97, Kno97a, MAK^{+}99, SK98, VW96b, KFD^{+}99]. Math [Tay99b, Gur98]. math/computer [Gur98]. Mathematica [SW93]. Mathematical [Cor99, AG96, Gin95, JLH^{+}98, Mye90, Nef93, TBF98]. Mathematics [Sco91, Alm91, Ber97b, Bri93, Bro93a, Fle93, Hen90, Hvo90, SW93, Wai92, War95, WC94]. Mathematizing [Sil94]. maturity [CKK94]. maximum [MB95]. may [BSM^{+}94, STM95, Wis92]. maze [SBB^{+}99]. MC68000 [Coe93]. means [Rad90, Sch90]. mechanics [HFKSM98]. mechanics-MechANIma [HFKSM98]. MechANIma [HFKSM98]. mechanism [Nap97a]. media [SH96, WWDM96, WDM97]. mediated [BCHB94, FCR^{+}94, Tay97b, WP97]. medium [Car97, Lev99a, Mos98, Wic95]. meet [HJ92, SY98]. Megaprogramming [HTW97]. Melbourne [SH97]. Mellon [GBB^{+}90]. membership [Dal92a]. memory [CCR99, JPPFPMVI97, O’N95]. Men [Mye92, Ber91]. mentoring [ABM93, HWB92]. mentors [GY99b]. merge [PH93, TR93]. merging [JM91, Tro91]. Message [Kle99b, Wal95, Ash97, Nev94, Wic95]. message-passing [Nev94]. metaphor [JPPFPMVI97]. metaphors [MPW99]. method [Bob99, Hit99, LSS^{+}96, MCH98, Pol99, Sob96, Zac97b]. methodologies [PS99]. methodology [Ant97, CHL96, Eva95, PRS94, VI94]. Methods [DH95b, GJMS90, ABW93, CK98, JRS96, KGZ99, PP95, Shi93a, Tre98, Tre99, Tro94a, WCDT95, War95]. metrics [Lea95]. MIC [Don95]. MIC-1 [Don95]. Microcode [Fol92, Ful92, SS95]. Microcomputer [EO95, Ewi93]. Microkernels [CO95]. Micron [Tho90]. microprocessor [Mat95, Ste91]. microprogram [Don95]. microprogrammed [Ber93, Tho90]. middle [Pil95]. Millennium [GPP99]. millionth [DL90]. mind [BP95]. mindset [Guz95]. mini [Jac91]. mini-course [Jac91]. Minimal [Tro92]. minimising [Ric98]. Minimum [dC99]. MINIX [CAH90, Har90]. minorities [MIG^{+}94]. minority [LEB^{+}90, Ola91]. minute [Mye98]. MIPPETs [SBB^{+}99]. MIS [Mil93, WUGV99]. Misconceptions [MNC99, HGW97]. Missouri [Dal92b]. Mistakes [Mil96]. Mod [JPPMMF96]. Model [GDC^{+}97, Ape99, BFG^{+}94, BSM^{+}94, Con98a, Din97, DMK97, GS99, Gor99a, HG99, JD97a, Kot98, Law97a, MBE^{+}93, Ras97c, She95, Tos93, ULW99, WGD^{+}95, WPI98, WC94, Zac97b, Zie98]. modeling [Ber94, Blu96, Cyn90, Mar95]. modelling [BJB98, HG98, MO96, Taf91]. models [Ber94, Cas98a, Cla93, HF98, HKSD97, OCH^{+}94, RHND99, SL94, Tow96, WD99, WDB98, YJ95]. modem [Lev99b]. Modern [MM94a, Noy95, SR91]. modes [Oli94]. Modula [Det93, HH95c, JPPM96]. Modula-2 [Det93, HH95c, JPPM96]. modular [OK95]. modularization [Bai91]. module [And90b, Arn95, Bar93b, JD97a]. Modules [Wal98b, AAW92, ASW97, CSB93, CR98, Nap97a, Tja98, Tra98]. monitor [VI94]. Monitoring [HSM97, BYC98, Jac96]. monitors [Dun91, OM95]. monkeys [LM98a]. Monopoly [TD90]. moral [Got98a]. Mosaic [Mor95]. Most [Scr91]. mother [Car93]. motivate [Gro98, Sch90]. Motivating [Tay98, CHL96, Gur98, Pre94]. motivation [FMM97]. movable [Fel92]. Moving [MCW99]. MP [Duc94]. MP-1 [Duc94]. MPS [MD99]. Mr [Hex94]. MSIM [SS95]. MSIS [Gor99a]. MSL [Els92]. much [GS99, Kol98]. Multi [Jes94, Kub99, WDM97, Bra96, HBC97, Kat94, Lan93, LM94, NB98, NW92b, SK98, SH96, WWDM96]. multi-disciplinary [NW92b]. multi-level [SK98]. multi-lingual [HBC97]. Multi-media [WDM97, SH96, WWDM96]. multi-paradigm [LM94]. Multi-Pascal [Jes94]. multi-phase [Bra96]. multi-processor [Kat94]. multi-purpose [Lan93]. Multi-team [Kub99]. multi-windowed [NB98]. multilevel [SH91]. Multimedia [EFGR94, GC98, AHST93, ACK^{+}96, AFVH97, CB96, CGJ93, CDVF97, Ell96, Ell98a, Ell98b, Gri99, HDC99, Hol98b, MO96, MCW99, MMK98, MS98, ML96, MAK^{+}99, PLG97, Ree97, RFPA97, SH99a, Tay96, Wan98, WS98, WGHL97, YSKW95]. multimedia/hypermedia [CGJ93]. multimodality [Hed99]. multiparadigm [BP95]. Multiple [LBK96, MS90, FJM98, RS97]. multiple-choice [FJM98]. Multiplying [Lee99b]. multiprogramming [HMOS90]. Multithreaded [She98]. MuPMoTT [MO96]. music [Rub95]. mutual [BC96]. My [And90a, Ada96a, Har99a].

N [CP94]. N-Queens [CP94]. Napier [Lee98b]. Nashville [LWMG95]. NATAL [GM91]. National [Cas98b, Har97, Yer94, HJ92, Har99c, Tay99a, Cas98c, MR99]. natural [KKL99, LC96a, RL95]. nature [Cle97b]. near [Kno98b]. necessary [Chm98, Mar99b]. Necessity [Car93]. need [BNS^{+}97, GSW^{+}98, Wu92]. needed [Gor98a]. needs [KACL94, OBH^{+}95, SBK94, SY98]. negotiated [Mag94b]. neither [DH96c]. Net [Aik99]. NetCp [FC94]. nets [Jef91]. Network [BYC98, Pat96, CHCL99, DDG97, Din91, DMK97, DS91, Ele99, GMAB96, Har96, Jac96, Kea93, KW96, MPH96, McD91, Str99, WW91]. networked [FST98, MK98, MPW99]. Networking [ACE^{+}90, DS91, BBJR96, Bar93a, SR99, Tym91]. Networks [JM96, ABFM90, AAHF^{+}99, Ash97, BS99, DMW^{+}97, FC94, GMAB96, Hol98a, Kea93, Lan98, MC90b, PSZ99a]. Neural [JM96, ABFM90, Kea93, KW96, MC90b, WW91]. Neuralis [KW96]. Never [BP95]. new-fashioned [Gor99a]. newsletter [ABDM90]. Next [BKP^{+}99, Cle98c, Wil97]. Nifty [PAC^{+}99]. Nightmare [DH96c]. nine [Cas99, DKT95, Kea93, MSPF95]. nine-tails [Cas99, DKT95]. nineties [ACE^{+}90]. non [BS95, BBD^{+}95, Cam94, DS91, Höf91, HL94, Joy98, Kay93, OBH^{+}95, SS97, SBD97, ULW99]. non-computer [Cam94]. non-deterministic [Höf91]. non-major [ULW99]. non-majors [BS95, BBD^{+}95, HL94, Joy98, Kay93, SS97]. non-network [DS91]. non-traditional [OBH^{+}95, SBD97]. nonmajors [Bie90, Gur98]. normal [Wu92]. Notation [Wit96]. note [CD93, Ful94]. Notes [Cle99a, BJB98]. novel [BWW96, Pax95]. novice [Bai91, BHPG99, BC92, LM98b, LA96, MY99, NF97, WK99a, WK99b]. novices [AAW92, Rab94]. NSF [Cas98c, AIW^{+}96, Joh94, LM91]. NSW [FR96a]. NTU [DH97]. number [War92]. numbers [Sco91]. numerical [BV97, Noy95, Shi93a].

O.S [PD95]. Oberon [Rad93]. Object [Ada96c, CF98, Cul99, DeC96, Dix93, HAB^{+}94, MM94a, Rei91, Rei95a, Sch97, Shi93a, ZBK^{+}90, Ada96b, Ada98, Ast98b, BHPG99, Bel92, Ber90, BD96, Con98a, CP96, DH93, DH94, DV98, Din97, DUM98, Dix91, FR98, Fie96, Güv95, Guz95, Had99, Has94, HGW97, Hol99, Hol94, KKR95, KR96a, KR96b, Lam92, Maz93, McC99c, NS94, OJ93, Par97, Pro98, Pro99, Rei93, RK97, Sca96, SBD99, Tan92, Tao95, Tem91, Tem92, TG94, Tew95, Wal96, War99, Wic95, WH96, WGHL97]. object-based [Ast98b]. Object-centered [Ada96c]. Object-orientation [DeC96, Hol99, Wic95]. Object-Oriented [MM94a, HAB^{+}94, Rei91, Rei95a, Sch97, Shi93a, Ada96b, Ada96c, Ada98, BHPG99, BD96, Con98a, CP96, DH93, DH94, DV98, DUM98, Dix91, FR98, Fie96, Güv95, Guz95, Had99, KKR95, KR96a, KR96b, Maz93, McC99c, NS94, OJ93, RK97, SBD99, Tan92, Tao95, Tem91, Tem92, TG94, Tew95, Wal96, War99, WH96]. object-ownership [Din97]. objective [FJM98]. objectives [Mar99b]. objects [KR99, Ras97b, WDH96]. OBOA [JD97a]. observations [Wal98b]. Observing [Tho98a]. Occam2 [MH91]. odds [Sco91]. off [Got97, Nap96]. offering [MM97]. office [JLH^{+}98]. Oklahoma [FRK96]. Old [Wil95a, AAHF^{+}99]. older [Mor92]. OLE [Kir96]. OLEIS [DT98]. On-Line [PC98, AB99, Kno97c, Med98, RG98a, Car99b, Gol96b, GK98, JL98, Kum99, PS99, RT98, WM98]. one [AR99, Dal90, Got97, TP94, Wal99d]. one-semester [Wal99d]. Online [Cle99c, TBF97, Law97a, MW99b, McC99a, McC99b, Tay97b, YDM99]. only [DS99, OJ93]. OO [Nef99]. OOP [Luk94, Pat97]. OOPL [dVdVI97]. OOPS [Eps93]. OOPSLA [Pfe99]. Open [DT98, Ros97, Thw94, Dav98]. Operating [Don90, RL93, And90b, Bar92, CO95, Car95, Car94, CAH90, Dow99, Eri96, Goh92, Gol96a, Gol97a, GBC^{+}99, Har90, Har92, Har94b, Har96, HMOS90, Hol97, Hug92, Jef91, Jon96b, Lea90, LC96a, MH91, Moe95, MD99, O’G98, Ree90, RR99, She98, Shu90, Tos93, WR97, WB92, Zie99]. operator [Din97]. oping [Har93]. Opportunities [SDJ^{+}94, BN98, BMR^{+}95, Cas98c, MAC^{+}90]. optimisation [BV97]. optimising [Ric98]. option [GS96]. oral [MM93, Sul93]. ordered [GK97, Tro91]. organiser [HSS98]. Organization [DH96b, Fek95, Mag94a, Nad96, PD96, Scr91, SH91, Wen90]. Organizational [MAG^{+}94c]. organizer [SP99]. Organizing [BO95]. orientation [Coo96a, Coo97, DeC96, Hol99, Hol94, Wic95]. Oriented [MM94a, Ada96b, Ada96c, Ada98, BHPG99, Bel92, Ber90, BD96, Cam94, CF98, Con98a, CP96, DH93, DH94, DV98, DUM98, Dix91, FR98, Fie96, Güv95, Guz95, Had99, Has94, HAB^{+}94, KKR95, KR96a, KR96b, Maz93, MAG^{+}94c, McC99c, NS94, OJ93, PM95, Par97, Pig92, Pro98, Pro99, Rei91, Rei93, Rei95a, RK97, SF91, Sca96, Sch97, SBD99, Shi93a, Tan92, Tao95, Tem91, Tem92, TG94, Tew95, Wal96, War99, WH96, ZBK^{+}90]. Orleans [JLP99, PN99]. OSs [OM99]. Othello [ES99]. Other [Wis96, AJ93, FP98, GL93, Wal98a, Wal99d]. otherwise [MW99c]. Our [Cle99b, Mos93, SH98, SLM98, CHL96, Fel99, Tre98]. overcoming [Bry99]. overloading [Din97]. overture [KP91]. overview [BBG^{+}96, Bie90, HH95b, KPW94, Win96]. OWL [Szu96]. own [BLOS98]. ownership [Din97]. OZ [Osb91].

paced [BS95]. package [DMK97, Hai96, JK91, JD97b, KW96, MH96b, WK99a, WK99b]. packages [Det93]. PACKET [MPH96]. page [KH99, Kos99]. pages [VW96b]. pairs [GK97]. palindromes [Sha98]. Panel [BKLV98, AIM^{+}95, AFVH97, ABDM90, ACE^{+}90, BGEL^{+}97, BNS^{+}97, CCR^{+}97, CDD^{+}97, CKM^{+}97, CDF^{+}97, DMW^{+}97, DFH^{+}98, EBD^{+}90, FACW90, FHL^{+}97, GSW^{+}98, GBB^{+}90, GJMS90, GL97, HSC^{+}90, HCD^{+}98, HKSD97, ICD^{+}90, IG97, IMAM98, JBCL98, KCC^{+}98, KLL^{+}98, KJ97, Law97b, LM98c, Lit90, LEB^{+}90, MAC^{+}90, MVS^{+}98, MCMHS90, MBW^{+}98, MHB^{+}98, MSSZ97, Nap97a, NI98, NKN^{+}95, NSS98, OBPB97, PAC^{+}99, PS97, PCRR90, SBB^{+}97, SAL^{+}90, SWWB97, SFR^{+}98, TBF98, VBB^{+}98, WBC^{+}98, WRRT97, WP97, YSKW95, ZBK^{+}90]. paperless [PP97]. papers [IMAM98, Ric98]. Papua [Bam98]. paradigm [BGGR98, BW96, BP95, Car98b, Had98a, Kam92, LM94, MPH96, Maz93, Rei93, Rei95b, Tem91, Tol98]. Paradigms [HH95b, Abu92, Cle97a, DBKL90, Dor97, JM91, LBK96]. Parallel [Bur97, Ele96, FG92, HA95, KKA98, LSS^{+}96, BAK99, CL98, Duc94, FG91, Gha91, Har91, HS91, HS93, HW91, Jes94, JY95, Joh92, Joh94, Juh99, KST92, Kot95, KAL^{+}96, LSS92, MZS^{+}95, MBJS90, Mer92, NC99a, Nev94, Rif94a, Rif94b, SH90, SK95, SCF^{+}93, Tol95a, TR93, YJ95]. parallelism [JTW90, Sch91, Smi94a, SPSS96]. Parameter [Fle91, BG93, NS96]. Parser [LK95, KW94]. Parsing [HP93, BJR94, HW91, KS98]. part [Cle98a, LK98, Lee98a, Mak98, Mar99b, Mar99c, Pet98b]. partially [TS98]. partially-individualized [TS98]. participants [BG97]. participation [Bro94, CT94]. participative [Jen98]. particular [LR94]. partnership [RHP90]. Pascal [Jes94, And90a, Dav90, Dix91, Dix93, GF96, JTW90, MC90a, Mul91, O’N95, Sch95, Smi94b, Web96]. Pascal-FC [Dav90]. passage [Ras97b]. passing [Ash97, BG93, Fle91, NS96, Nev94]. past [Dal99b, Kay96]. path [Cha91a, Har91]. Pattern [EW97, Nar91, EL97, FPR98]. Pattern-based [EW97]. Patterns [CL99, NW99a, AMBC98, CH98, GGT98, Ngu98, PCLM97, Pre99, Ree98, Wal96]. PC [Det93]. PCs [BLOS98, WT91]. PDP [Ful91]. PDP-11 [Ful91]. pedagogical [CC92, EL97, Eze90, War92]. pedagogies [ACK^{+}96]. pedagogy [CL99, Eva96, Kum98, MCD^{+}94a, Mul91, RL95, Win96, WH96, ACK^{+}96]. Peer [KFD^{+}99, WFGW94, CKM^{+}97, FW96, Sul94]. peg [CJ91]. pennies [Hec94]. Pennsylvania [IAK96]. people [PH97]. Perfect [Tro92]. Performance [FG92, Bar93a, Blu96, GP91, Hei99, Juh99, Leo91, MZS^{+}95, TS98]. peripheral [Eck90]. permutation [Lia95]. Permutational [FJM98]. personal [BCHB94, FLP90, Gro98, HT98, Kat94, Smi94b]. personality [HTW98]. personnel [KLL^{+}98]. perspective [BT96, CT94, DUM98, Fel90, GBL94, IBF^{+}94, Lit90, O’D98]. Perspectives [IG97, KCD^{+}96, Kno96]. Petri [Jef91]. phase [Ada96c, Bra96]. Philadelphia [IAK96]. philosophers [Fel92, Gin90, bY91]. philosophy [KAK93, KPW94, LR94, Pat90]. Phoenix [BG94]. photo [RR96]. physical [Ber98, Ful94]. Pi [Rog90]. Pictures [Ast91]. pilot [Wil97]. pipeline [BCPV98, KK98a]. pitfalls [BT98, GVLZ94]. place [Bal97, OCH^{+}94]. Placement [SAL^{+}90, ACRW97, BAH97, CFA^{+}94, KS94, NKN^{+}95, SFR^{+}98, SRL^{+}99]. plagiarism [Car99c]. plague’ing [Wis92]. plan [New93]. plan-based [New93]. planetary [FP97]. Planned [Kno99b]. planning [OBPB97]. platform [BMP96]. play [PLU97, Sul93]. Playing [CLV99, DH95b, KST92, SP96]. plight [Ola91]. plurality [Hed99]. pluribus [dVdVI97]. point [Chu99, Coh99, Hai97]. pointer [Wha91]. points [Tol95a]. Poland [Imp99a]. policy [Cul98]. polynomial [Bre92]. portability [Moo92]. portable [Fag94, Fel92, Rob96, Tho98b, TLNS98]. portfolios [KCD^{+}96, Van95]. possibilities [Cul98]. Possible [Cla93, MBW^{+}98]. post [AP91, YJ99]. post-symposium [AP91]. Poster [Ack99, CF98, Cas98b, Cla98a, Cle98a, Coo98, Dal98c, DI98, Dru98, Ell98b, FR98, FST98, Hol98a, Hol98b, Joh98, Kin98, KPT98, Kol98, Kum98, LP98, LN98, LZ98, McN98, MV98, MMK98, MS98, Mos98, Nak98, OQ98, PR98b, RSBAS98, Ric98, RG98b, RM98, RT98, SY98, Smi98, Tho98a, Tra98, Wan98, WNKK98, Wol98, Zag98]. postgraduate [SH96]. postlingually [Wan98]. power [Eps93, Joh98]. practical [Kou98, LAH94, Rab94, UD97, WR97, Wal99d, Wu92]. Practice [Dor97, Ape90, DU96, Ele99, KWJ^{+}96, RL93, Utt99]. practices [Gro98, SC94]. practitioner [Got98a]. practitioner-licensing [Got98a]. pragmatic [Dal98c]. pragmatism [SCVI91]. PRAM [MGH98]. pre [Alm91]. pre-college [Alm91]. precision [MS90]. precursor [Hen90]. predicting [Juh99]. predominantly [CH91]. preferable [Wis92]. preliminary [Hol98b]. Preparation [Fek96, DFH^{+}98, KLL^{+}98, Tay97a]. Preparing [Kus94]. present [Kat95]. presentation [BJ98, O’N95, PFR^{+}94]. presentations [Mak98]. presenting [Ben92]. primary [Rad90, ZMS95]. primitives [Mul91]. principles [Bru99, Dal91a, DeC96, KH94, KJ92, MAD^{+}95, RS94, Van98]. prior [Tea92]. Prisoner [Res96]. private [MAC^{+}90]. proactive [Got98a]. Problem [Win93, Cas99, CP94, DDP96, Ell98a, Gib92, Guz95, Hex94, HRBK99, Joy98, KK98b, Kea93, Kir96, KA95, Koo95, Kor94, KMO95, Llo94, MB95, Mye98, Pre94, Ree98, Ros91b, Stu94, TJ90, Zac97b, dM98, Gra93, Yer94]. problem-based [KK98b, Kir96]. problem-solving [Ree98, TJ90]. problems [AAHF^{+}99, CS98, Che94b, Che94a, Cor99, DeC96, Jon96a, MAC^{+}90, NPP^{+}93, Pou90, Rad90, Shi93b, Zie98]. procedural [GS99]. procedures [Pat93]. proceedings [JLP99, ACM98, ACMJ90, BG94, Car98a, Dal91b, Dal92b, FR96a, IAK96, Imp99a, KLY93, LWMG95, LPJI98, PN99, SH97, WEKM97]. Process [KFD^{+}99, BAK99, Ber96, BJB98, CKK94, Dyc93, EO95, Gro98, HMOS90, HT97, HT98, MO96, OD94, PSZ99b, Shi92]. processes [CHCL99, RRR93, Sig90, Siv96]. Processing [KKC90, Asl98, Che94b, CL98, Duc94, HS91, KAL^{+}96, ME92, McN98, SV96, SCF^{+}93, SPSS96, Tol95a]. processor [Hil90, Kat94, Liv97, VdlFCC97, dlFCC99]. procurement [Ber92b]. product [DH96c, UD97]. production [PM96]. profession [Den99, Mar98c]. Professional [EL94, Cun95, Mar98b, MSRT95, Osb92a]. Professionalism [LGB^{+}99, NI98, WG93, LGB^{+}99]. professionals [PH95, Wer99]. profiling [HAG92, Noh94]. Program [Har99c, Wad95, Wal90, Wis96, Arn91, BBG^{+}95, Chu99, DL90, FF94, HH95a, Juh99, KR96b, Kos99, LC90, LM94, Mac97, MR99, MV95, McM96, Mos97, PM96, Ros91a, SKL98, STM95, TLR94, Wha91, ZC99]. programmer [BC92, LA96]. programmers [Bai91, Dal98b, NF97, Osb92a, Sch97, WHN93]. programmes [BJB98]. Programming [Cle98c, Joh98, KAK93, LP98, Mäk96, Max93, McC92, MM94a, NPP^{+}93, PCRR90, Van95, Win96, Wit96, Ada96c, Arn94, Arn95, AC96, AB99, BO95, BK95, BHPG99, Bel92, BAS99, Ber90, Ber96, BF94, BW96, Bru99, Bur97, CB94a, CBH97, Can91, CJ99, Cha93, Cha91b, CPS93, Cun91, Cur92, CR98, Dad90, Dal91a, Dal99d, Dav90, DJ99, DH95a, DH96a, DH93, DH94, Dix91, Dor97, Dra98, EW97, EHS^{+}99, ES99, EL94, Eze90, Fek93, FLP90, FB97, FG91, FH96, FR96b, Gei94, Gib92, Gin96, GBB^{+}90, Gre96a, Gro98, dOG93, Guz95, Had98b, Had99, HSM97, HS98, HCD^{+}98, HS93, Has94, Hay98, Hec94, Hen90, HP94, HP95, HM91, HAB^{+}94, HH95c, Hou97, Hvo90, Jen98, JPPMMF96, JPKPM99, JCF^{+}93, KK98b, Kea93, Kin92, KLMS95, KSW98, KR96a, KR96b]. programming [Kot95, Kur91, KJ92, Kus94, LBJ94, Leb99, Li98, Lia95, MMKW99, MHP93, MK98, MH96b, MH96c, MHP98, Mer90, Mes91, MCd^{+}94b, Mit96, MS98, MN93, MS90, ML96, NS96, Náv94, Nef93, Nev99, NP92, OM95, Pan90, Pat90, PH93, PP97, Rad90, Ree95, Rei91, Rif94a, Rif94b, Rob92, Rob95b, RLR95, RL95, Ros99, RT98, SCVI91, SH90, SG91, Sca96, SH95, Sch97, Sch96b, STBC94, SY98, She98, Shi93a, SC94, SW91, Slo95, SWWB97, SR91, Str91, SD93, SLM98, Tan92, Ter95, Tol95b, TJ90, TZ99, Uck91, Van98, VI96, VIRF96, Wai92, Wal98a, WW91, WS98, WSK97, Wil95b, WM98, Wol92, WC94, WWDM96, WK99a, WK99b, YB99, Yea91, bY94, Zac94, Zac97a, Zac97b, ZBK^{+}90]. programming-are [HCD^{+}98]. Programs [FG92, And90a, Ara91, Arm99, BSM^{+}94, BD93, CH98, CDF^{+}97, DNPR94, DM92, Fek93, FRK96, FHL^{+}97, GT99a, GHK93, GDC^{+}97, Gor98a, HKH^{+}93, HCD^{+}98, How94, JU97, JPPFPMVI96, Joh99a, KSIR94, KI93, KLL^{+}98, KBCP99, Lea95, MM97, MM98, MY99, MN93, Nad96, Nap98, NC99b, Rob96, Rog90, RK97, Sap91, Sch95, Tay97a, VI94, Wis92]. progress [For94, KLE92, Mer92]. progressive [Mar99c]. Project [GY97b, Tro92, Ada93a, Ada98, Asl98, Bar96a, Ber98, BG97, Cas96, Cla98b, CCMD99, Els92, Eng99b, FP98, FC94, GG99, Goh92, GZ94, HSS98, Hag91, Har94a, HWB92, KLMS95, Kub99, KP90, LG96, Lea90, Lea97, Lee90, MMKW99, MAG^{+}94c, Mer90, MC90b, Nul98, NRR^{+}90, Oli94, Pil93, SF91, SP99, Sha99, SH96, Sha91, Sul93, Tem92, Wil97, WDM97, ZMS95, PC98]. Project-after [GY97b]. project-based [SP99]. project-intensive [Ada93a, SF91]. project-oriented [MAG^{+}94c]. Projects [RL93, BK95, Ber93, Bou97, Bri91, Bro93b, BD99, Cha93, Cur97, DH96b, DU96, Dru98, Ell96, Etl90, Gam95, Hus95, JPKPM99, Joh94, Kea93, Li98, Mäk96, McD96, Mor95, PH99, Pat90, Pig92, Pou90, RR97, Rei90, Rob92, Rub95, Sab90, SBTC94, STBC94, Sil91, Son96, TZ99, Wer90, Wil95b]. Prolog [Kam92, MGH98, New93]. prologue [Kay96]. promise [BGEL^{+}97]. promises [GVLZ94]. promote [Alm98]. Promoting [YD99, Hus95, PH97]. Proofs [Wad95, Cor99, GT98]. properties [Whi98]. property [YD99]. Proposal [Gol96b, EGIU99, Gro97b, Har98c]. Proposed [WGD^{+}95, Cup98, Dal98c, MDG^{+}95, Nac98, Pra90]. props [Ast98a]. Protocol [BC92]. Protocols [GMAB96, GZ96, NG94, SG97b]. prototype [Car95, Lev97]. prototypes [MPTW92]. prototyping [RHP90]. proves [CBH97]. provide [LN98, VW96b]. Providing [LWB^{+}98, MW99b, NW92a, SH96]. psychological [CT94, Tay97b, Win96]. psychology [GZ94]. public [Bur97, CH91, Nac98]. publication [Kno97c, Shi92]. Publishing [Dal98a, KG99]. pumping [JM90]. purpose [Lan93]. purposes [Gsc94, KPT98, KPT99]. pursuit [Smi94a]. pushdown [Han92]. puzzle [CJ91, Pru93]. PVM [MK97]. Pythagorean [Gib92].

Quality [AAHM99, KFD^{+}99, Bal99, BD97, BG97, Cun97, GS96, HT97, MV95, MY99, Rou92, Roz94, Ske93, Sze99]. quantitative [LSS92]. quantities [Ber98]. Queens [CP94, Gra93]. Queensland [Car98a]. query [Asl98]. quest [FZ96]. question [Ros97]. Questioning [Bar97a]. Questions [Bar97a, FJM98, dM98]. queue [Sil90]. queueing [FP95]. quickly [Joh98]. Quicksort [CD93]. quiz [Dad90]. quo [GC98].

Rôle [DH95b]. Rôle-Playing [DH95b]. racquetball [Wal97b]. railroad [Tos93]. Raising [Tea92]. random [JD97b, Mit96, War92]. randomization [GT99b]. rapid [RHP90]. rat [Dor92]. rats [GZ94]. Ray [Sch90]. RCOS.java [Jon96b]. Re [PBA98, Gra93]. Re-engineering [PBA98]. re-examined [Gra93]. reaction [Gol97b]. readable [Ara91]. reading [SAL^{+}90, Wal99b]. Ready [AHST93, Har98a]. Real [HW90, Bro93b, FP98, FH91, FST98, McC91, Nac98, OM99, Sco91, Son96, VB98]. real-life [Son96]. real-time [Bro93b, McC91]. real-world [Nac98, VB98]. reality [BGEL^{+}97, Bel96, Gol98, PD95]. realization [CDVF97]. Reasoning [Fek93, Ant97, ACRW97, RHND99]. reasons [DH94]. Reciprocal [Sul94]. Reconstructing [Got98b]. Recursion [Pro97, GS99, Hay95, New93, Rev98, WLC96, WDB98]. Recursive [Chu99, Den96, Rob99b]. Redesign [HJ92]. redesigned [HSM97]. Reexamining [APNA90]. Refereed [IMAM98]. reference [Ast94]. Reflections [Dal99b, Rob99a]. register [SH91]. reinforcing [BT95, WR97]. related [MRA^{+}98, MR98, NB98, PH95]. relational [Ape99, DEP97, LHA90, Uck91]. relationship [Alm91, Ape99, HG98, MV95]. relevance [Lar97]. relevant [Dal99a, RS97, Wil95b]. remote [MWO^{+}96]. Removing [Hea95, STM95]. rendering [GEC96, KK98a]. RenderMan [Owe92]. reopening [Rob95b]. repeat [SH98, SH99b]. replacement [KH99]. Report [BKP^{+}99, FBR^{+}99, KFD^{+}99, LGB^{+}99, MM97, Wal99c, ADB^{+}96, AJ93, BHH^{+}99, CDF^{+}97, Gol99, GCD^{+}94, HGPJ^{+}96, KAK93, KLE92, KI93, KWJ^{+}96, MM98, Mer92, PBEC93, Pfe99, RSLD99, ACK^{+}96, BBG^{+}96]. Repository [Kno98b, IHA^{+}94]. representation [Hai96, Khu94]. representations [Ben92]. Reprogrammable [Gsc94]. Republic [HWS92]. Requirements [KKR95, RF96, Bar97b, WUGV99, dC99]. Research [ABW93, DJ99, LEDW93, AW93, AWB94, Alm98, BJ98, Ciz97, CDD^{+}97, Cun94, Cun95, Cun97, CEMP97, Dal98a, Dal99a, Dal99b, ES99, Fek96, Har98b, How93, JPS90, KCC^{+}98, LM91, LM94, LEB^{+}90, MVS^{+}98, MRA^{+}98, MR98, RL95, SG99a, SG99b, WB93]. Research-led [DJ99]. research-welcome [CEMP97]. Resource [Aik99, AW93, AWB94, Gol97b, HGPJ^{+}96, HMOS90]. Resources [BKP^{+}99, Lan98, BNB^{+}98, BGG^{+}99, Cun97, Ell98a, Ell98b, EHS^{+}99, GKC^{+}98, HS91, KG99, LN98, McC98a, McC99a, McC99b, Siv96, WB99]. respect [WUGV99]. responsible [Mar99b, Mar99c]. restricting [Cle99b]. Restructuring [Lam95]. Results [Dal92a, NPP^{+}93, BBC^{+}94, Dal98a, FST98, Gol97a, HH95c, Hol98b, Lev95, MVS^{+}98, MJ99]. Retaining [HF98]. retention [Coh99, LB96, VBB^{+}98]. Rethinking [Fie96]. retrieval [Cun96]. retrospective [Etl90]. reusability [BT96]. reuse [KG99, LAH94, NF97, Tew95]. Review [KFD^{+}99, WAK96, Gol96b, MW98b]. reviewer [Har98c]. reviews [Sul94]. revisited [Ada96b, Gin90, Gor99c, Lit90, Res96, Tro94c, Yer94, bY91]. Revitalizing [BBD^{+}95]. Reviving [Wol92]. rewards [Bec94, PH99]. rich [Bal97, Bec94]. right [Lev99c, YD99]. Rigorous [Wad95]. rigour [SCVI91]. Río [SLM98]. risks [PH99]. road [LN97, McC99a]. RoboProf [Dal99d]. robot [Eps95, KM98]. Robotics [Ste91, Stu94]. RoBOTL [LC96b]. role [FCR^{+}94, HF98, HT97, Mye90, Nac98, PW99, Rog90, Sul93, Tan90, Tow96]. role-play-team-project [Sul93]. ROM [Eng99a, Van98, WGHL97]. ROMs [VW96b]. room [Pro99, WCMT94]. roommate [Ast92]. root [Hig93]. roots [Ada96b]. Roundtable [GBB^{+}90]. RSI [PH97]. RTN [HP93]. rubrics [KS94]. rules [Fle91, Uck91]. running [Bre95].

S [Det90]. S/360 [Det90]. sabbatical [BMR^{+}95]. safe [Nev99]. same [Ber91, OCH^{+}94]. San [Dal91b, WEKM97]. say [How93]. scalable [Has94]. scale [BD97, HAB^{+}94, PS98]. Scaling [DAUZ94, FPR98]. scenario [Eps95]. Scheduled [LS93]. scheduler [KH99]. scheduling [EN90, Tan92]. Scheme [Hay97, KH94, Lam92, Mar95, BT95, RFH^{+}93]. Scholastic [NPP^{+}93, PCRR90]. school [AG96, Bro97, Bro94, BSM^{+}94, CFK^{+}92, FHS^{+}94, FCR^{+}94, HW90, LW97, MBE^{+}93, Pil95, RSBAS98, Rif94b, RW96, SH95, SW93, Sch96b, Son96, Str91, WBC^{+}98, WN96, ZMS95]. school/college [WBC^{+}98]. schools [GL94]. Science [ACM98, ACMJ90, BG94, Car98a, Cas98b, Cas98c, Dal91b, Dal92b, FR96a, Har97, IAK96, Imp99a, JLP99, KLY93, Kum99, LWMG95, LPJI98, MR99, PN99, SHY96, SH97, SS97, WEKM97, YD99, AGT98, AMN98, AFVH97, Add97, APNA90, Alm91, AP91, AW93, ABM93, ABW93, AWB94, Ama90, AAK^{+}94, AM90, Ast92, Bal90, BK92a, BSK94, Bal96a, Bam98, BBJR96, BHH^{+}90, BA98, Ber94, BNB^{+}98, Bie90, BBG^{+}95, BKS^{+}99, BEG^{+}96, BGGR98, BGG^{+}99, Bou96, Bro97, BSM^{+}94, BABT97, BKLV98, BD93, BBD^{+}95, CH98, CK98, CDMT94, CGK^{+}96, CK99, CFA^{+}94, Che94a, CH91, Chm98, CHLS95, CFK^{+}92, CD95, Coh99, Coo96a, Coo97, CKM^{+}97, Cun91, Cup98, Cyn90, DGEST96, DP90, DAUZ94, DH90, DM92, DL95, Epp91, Ewi93, FK92, FG96, FACW90, FPC96, FPR98, FRK96, FW96, FMM97, FH96, FHL^{+}97]. science [GS93, Gei91, GY98b, Gol97b, Gol96b, GJMS90, GZ94, Gre96b, Gre97, Gre99, Gri99, GNR^{+}95, GKC^{+}98, GK98, Haf94, HTW98, HKH^{+}93, HH95a, HJ92, HCD^{+}98, HH95b, Har98b, Har99c, Hil93, HT97, Hil90, HL94, HKSD97, How93, IBF^{+}94, JPS90, JMM97, Joh92, Joh94, Joh97, Joh99b, JM91, Jon96b, JBCL98, Kat95, Kay92b, Kay93, Kay95, Kay98a, Kay98b, KCC^{+}98, KV94, KW93, KB97, KCD^{+}96, KS94, KPW94, KWJ^{+}96, Kno99a, KA95, Kol98, KH94, Kou98, Kub99, Kum98, KDE^{+}93, Lam95, LK98, Leb99, Lee96, Lee97, Lee98a, LM98a, LB96, Llo94, Loc91, LR94, LEB^{+}90, Mag94b, MZS^{+}95, Mar98c, MW98a, MBJS90, MAD^{+}95, MJM96, McC97, MM97, MM98, McC98a, McC99a, McC96, McC92, MF98, MM93, MM99, MS97, MMB^{+}91, MH96c, Mer92, MBE^{+}93, MSSZ97, Moo93, MWP94, Mos93]. science [MPHL95, MDG^{+}95, MHN98, Mye90, Mye92, MW98b, Nac98, Nap90, Nef93, NKN^{+}95, NC99b, Nul96, OK95, Ola91, OBPB97, PMB98, PP95, PRS93, PRS94, PA90, Pes91, PL96, Pra92, Pre96, PRF96, RPF94, Ras97a, Ree96a, Ree96b, Ree97, RF96, RFH^{+}93, RS91, RS94, RC97, Rod95, RW96, Rog90, Rub95, Sab90, SS94, SP96, SBB^{+}97, SP99, SSH98, Sch99, SG96, Sco91, SVP93, SBK94, SS98, Shi92, Sie99, SBD97, SPV99, SD93, TM92, TP93, Tay97a, Tho98b, TLNS98, Thu90, Tow96, Tra98, Tri99, TW94, VB98, WCMT94, WBCS95, WCDT95, Wal98c, WMZ^{+}99, Wal99d, WRRT97, WFGW94, WC94, Yah92, YR95, YJ99, ZJEP95, Zac97a, ZMS95, ZBK^{+}90]. sciences [GHK93, ICD^{+}90, KI93, LZ98, Pra92, VBB^{+}98]. Scientific [BWW96, Moo93, BK92a, Zac97b]. scientists [Kay98a]. scope [NS96]. scoring [KS94]. scotty [Jac96]. search [Ada96b, Cha91a, Cor99, HP93, Hig93, Kho90, KMH91, Ras97c]. Searching [Har91, Bra96]. seasons [Mey96]. SECD [Slo95]. second [Cha93, RL95]. secondary [LW97, SW93, TM92]. secretary [Wal99c, Wal95]. secretary/treasurer [Wal99c, Wal95]. secure [MK98, MK99]. security [Amo93, EGIU99, Hus95]. Seed [Cle99b, Rob99a]. segment [MB95]. SEI [For91b]. selection [Kum96, dVdVI97]. Self [Ast94, BF94, PR98b, BS95, BGG^{+}99, Eng99a, Sch95, Tea92]. self-assessment [Sch95]. self-contained [BGG^{+}99, Eng99a]. Self-evaluation [PR98b]. self-paced [BS95]. Self-reference [Ast94]. Self-sufficiency [BF94]. semantics [Bru99, KOC91, Kur91, KJ92, Slo91]. semaphores [Hil92]. semester [Bar96a, Sha99, Wal99d]. semester-long [Sha99]. Seminar [Nev99, AM90, Ast98b, Bel92, Cas98c, Cul97, EW97, Kay93, Kay95, Kou97, Kou98, Lel94, Reb98, Wal99b, Zac97a]. seminars [FST98, Tay97b]. seminars-real [FST98]. senior [AM90, Kat94]. seniors [Bry99]. seniors-overcoming [Bry99]. SEPA [Joh99b]. sequence [Ape90, CBH97, Coo96b, PD96, SBD97]. serial [HS93, WCC^{+}92]. served [Nap96]. server [CDVF97, DMK97, YB99]. Service [GY98b, AAHM99, EL97, HG99]. session [Ack99, ABDM90, ACE^{+}90, EBD^{+}90, FACW90, GBB^{+}90, GJMS90, HSC^{+}90, Hug90, ICD^{+}90, Lit90, LEB^{+}90, MAC^{+}90, MBJS90, MCMHS90, PCRR90, SAL^{+}90, Wal90, ZBK^{+}90]. sessions [HG99]. set [Cha93, GK97, NW92a, PM96]. sets [Güv95]. setting [Gol99, Pra92]. settings [WP97]. seventh [BMR^{+}95]. shape [BSM^{+}94, Cun98]. Shaping [PZ96]. shared [EHS^{+}99, Ros99]. Sharpening [GY99a]. shell [JD97a]. shift [BGGR98]. shop [WGHL97]. shop-using [WGHL97]. shortest [Har91]. Should [Shu90, GS99, KCA99, LGK^{+}93, Mar99a]. show [SRY98]. side [Bei96]. SIGCSE [ACM98, ACMJ90, BG94, Car98a, Dal91b, Dal92b, FR96a, IAK96, Imp99a, JLP99, KLY93, LWMG95, LPJI98, PN99, SH97, WEKM97, ABDM90, AP91, Dal92a, IMAM98, Kle99b, Wal95, WBC^{+}98, Wal99c]. SIGCSE-98 [ACM98]. SIGCSE/SIGCUE [Imp99a]. SIGCUE [Imp99a]. signal [Har99a]. signals [JD97b]. significance [Sze99]. SIGs [RS97]. Sim [GT99a]. similar [NB99]. Similarities [Wis96, Wis92]. similarity [GT99a]. Simple [Pil93, Arn95, Bar95, Ber96, Cam96, Juh99, RK97]. simpleIO [WK99a, WK99b]. simplified [STM95]. simulated [Pat96, Str99].

Simulating [Bar97b, Ber93, Ber92b, Cam96, Rev98, RHND99]. Simulation [HMOS90, Mag94a, Nad96, SG97b, Taf91, Thu90, TR93, Blu96, Cyn90, DMK97, Fol92, Han92, HDC99, Kno97b, MdL92, MD94, Pro98, Pro99, Rei95a, She95, Tan92, Wal97b]. Simulator [GZ96, Bar93a, Bar95, CCR99, Don95, Ful91, Ful92, KCPC98, Lan93, Liv97, Mey96, MD99, NP92, PBA98, Sea93, SS95, SH91, Sty94, Tan90]. simultaneous [NB98]. singing [Sie99]. single [Col91a, Har91, KP90]. single-source [Har91]. site [Alm98, Reb98, VW96a]. site-level [Reb98]. sites [Hei99]. Situated [MM94a]. situations [MS98]. sized [Lev99a]. Skeptic [Aik99]. skill [Cun94]. Skills [Bar97a, BK92b, BJ98, Fek95, GY99a, Gre96b, Haf94, HW90, JPS90, Kou97, Mar99b, MM93, MWP94, NW99b, Pol99, SBD99, Sul93, TP93]. Slow [Jul92]. slowing [CD93]. SLR [KW94]. small [BD93, Lev99a, MAC^{+}90, Tur94, FLP90]. SMALL-Ada [FLP90]. Smalltalk [SW91]. social [Mar99c, PH96, SG97a, Sta93, Tay97b, TW97]. societal [GL97]. society [Rog90]. Socket [Tol95b, YB99]. soft [Bei96]. Software [BHH^{+}99, Bar97a, Bry99, DFF99, GS96, GCDP94, HMBO98, Mod91, NB99, Oli94, Osb95, Sai93, Tew95, TD90, TA96, Ada93a, BW97, BNS^{+}97, BN98, BFG^{+}94, BG97, Bou97, Bri91, Bro93b, Bur97, CS98, Cli91, CKK94, Con98a, Con98b, CBPW96, DS98, DUM98, Dru98, Etl90, FC99a, Fel92, For91a, For91b, For94, Fre94, Gam95, GYT^{+}93, Ger94, GG99, GP91, Gro97b, Gro98, HML^{+}99, HP94, HT97, HT98, HK96, JMM97, Joh97, Joh99b, JK91, KH92, KLE92, KP90, KAL^{+}96, Lea97, Leo91, LM94, LN97, MAG^{+}94c, MAD^{+}95, MJ99, McC91, MHB^{+}98, MM94b, Moo92, Mur91, ML90, Nac98, Nak98, NRR^{+}90, O’D98, OK95, OD94, PP95, PH99, PCLM97, PW99, Poo91, Ree96a, Rif94b, RS94, RB92, Rou92, SV96, Seb97, SH96, Sha91, Ske93, Sob96, Son96, Sul93]. software [Sze99, Tja96, Tom91, TLR94, VW96a, Wer90, Wer99, Wol98, WDH96, WH96, YD99]. software-based [HP94]. software-starting [WDH96]. Solid [Mar95, Ber94]. solution [MW99a, Tow99]. solutions [AAHF^{+}99, Jon96a, Rad90]. solve [CS98, DKT95]. Solving [Koo95, Che94b, DDP96, Joy98, KA95, KMO95, Ree98, TJ90]. Some [Che94a, GM91, Lea97, AGT98, BWW96, Cle97a, HH95c, HS97, SR91]. Songs [Cha96]. sophomore [KCPC98]. sophomore-level [KCPC98]. sort [TS95, TR93]. sorting [DB98, Jul92, McN94, MN90]. soundcard [Lev99b]. source [Har91]. sources [FACW90]. Spanish [Che94a, DA97]. speaking [Max93]. special [ABDM90, PCRR90, SH99b]. specialists [Wil97]. Specialization [Fra98, BBJR96]. specific [Kor94]. Specification [TSG98, Gam95, McD91, Win93]. specifications [Hil96, Sai97]. specify [Sil91]. spice [HS97]. SPIN [Pru93]. SPIN-OUT [Pru93]. spouse [Ast92]. spreadsheet [KA95]. spreadsheets [Juh99]. spring [KLMS95]. SQL [Mit98]. square [Hig93]. SR [Har92]. stable [Ast92]. stack [Wil90]. stage [TP94]. STAMPS [Liv97]. Standard [CS99]. standards [HJ92, KS94, MJM96, RS97, TM92, Tay97a, Tay99a]. Stanford [Rob93, RLR95]. started [CDD^{+}97, SH90, Wer97, WB93]. starting [WDH96]. state [Cig92, Han92, Kat95, Liv97, MD94, MW98b]. state-machine [Liv97]. States [HWS92, Tay99a]. static [MY99]. stationary [JD97b]. statues [PH97]. status [Cas98b, GCD^{+}94, JBCL98]. Stella [Taf91]. step [Cle98c]. steps [MVS^{+}98]. stereotyping [Gre99]. still [Wic95]. stimulating [Mit96]. STL [Mal97]. Story [Ger98, WDM97]. Strategic [WRRT97]. Strategies [Din91, JMM97, Rob98, Der93, DFZ^{+}95, Kay98b, Kin98, PH97, Tay99b]. strategy [Cle98a, EHS^{+}99, SSN99]. stroke [Col91a]. Structure [BA96a, SHY96, Cle98b, Dal98c, GGT98, GBB^{+}90, MD99]. Structured [Par95, Cha91b, MO96, Mil93, Rob95b, Tan92]. structures [BBG^{+}99, Bie93, Eps93, Fel90, GT99b, Hag91, Hei93, Jar94, JY95, Kea93, Koo95, KCA99, Kro96a, Kro96b, Lev93, LAH94, McN96, MST97, MM94b, Ngu98, NW99a, PR98a, Pre99, Sil91, TGGP97, Wei97, Wha91]. Structuring [Smi98]. Student [GY99b, Har94a, Kro92, Mes91, MMK98, Ber96, BK92b, Bon99, CHL96, Ciz97, Dru98, Fel99, GY98a, GF96, Gol97b, GP91, Gre97, dOG94, Guz95, HSC^{+}90, Hei99, HL98, How94, JU97, KSIR94, Lea95, Leo91, LH99, Mac97, MV95, MY99, MWP94, Nap98, OBH^{+}95, Pou90, PSZ99a, RG98a, SBTC94, STBC94, SY98, SG99a, SG99b, Sta97, SD93, Tay99a, TS98, WPI98, WHN93, Wis92]. student-built [Sta97]. student-controlled [CHL96]. student-directed [Bon99]. student-made [dOG94]. Students [Bar97a, MM94a, Arn94, AAK^{+}94, BEG^{+}96, BJ98, BJB98, Bro97, Bru94, Cam94, Car97, CF98, Car99c, CHL96, CFK^{+}92, Cun97, CR98, Cyn90, DG99, ES99, FHS^{+}94, Fle91, Fre94, Gro97a, Haf94, HTW98, HKH^{+}93, Joh98, Kin98, KACL94, Kol98, KH94, LP98, LB96, Llo94, MW99b, MH96b, MMB^{+}91, MH96c, MV98, Mos93, Mot96, MRS93, Mye92, Pfe99, PH95, Pil95, RSBAS98, Ree96b, Ree97, Rif94b, RSB98, Sap91, SH95, Sch90, SH96, SH98, SH99b, SH99c, SBD97, TP93, Tea92, Ter95, Tho98a, TH98, Tja96, Tja98, VS92, VW96b, VB98, WS98, WB93]. studies [CFA^{+}94, CL92, Dal98b, LR94, NKN^{+}95, RSB98, SS97]. studio [Tom91]. study [Ast92, BD97, Ber92a, CBPW96, Din97, GJMS90, Har91, HW91, HAG92, HWS92, Leo91, LN98, Mag94a, MNC99, MY99, MN90, Pol99, Pro98, RRR93, SS98, SH98, Tja98, Wal97a]. studying [BBF^{+}94, FLP90, Kur91, MH96b, Nad96, RG98b]. Style [LC90, CK99, HKSD97, PA90, Sch95]. styles [HCL96, LM98a, WDB98]. subgroup [AJ93]. subject [PR98b, RF96]. subjective [GY99a]. submission [DH95a, DH96a, Mac97]. submodule [CO95]. subprograms [O’N95]. subsequent [TS98]. subset [ACRW97, BAH97]. subsets [Rad90]. success [Chm98, DH96c, GF96, HKH^{+}93]. Successful [GHK93, BMC95, MW99c]. Successfully [CP96]. sufficiency [BF94]. sum [MB95]. Summary [AP91, BBC^{+}94]. summer [SW93]. SunOs [Ber96]. super [Har93]. superscalar [VdlFCC97, dlFCC99]. supervised [LS93]. supplementary [MH96b]. Support [CGJ93, AW93, Cun94, DH97, EHS^{+}99, FK92, KLL^{+}98, LA96, Mak97, MP96, Ree96a, Smi94b]. supported [CHCL99, FW96, Joh94, TH98]. Supporting [BD99, VW96a, DFZ^{+}95, MO96, Mye92, NSTT96]. supports [IHA^{+}94, MS98]. surface [ZZLS99]. survey [CK98, Dal92a, DH90, DM92, Dra98, Hol98b, Lar97, Lev95, Maz93, MM97, MM98, MV98, MN93, NC99b, PP95, Pra90]. Swan [SHY96]. swapper [CAH90]. Sweden [Cle99a]. swimming [GZ94]. Swinburne [PC98]. Sydney [FR96a]. symbolic [Kro96a, Kro96b]. Symposium [ACM98, ACMJ90, BG94, Dal91b, Dal92b, IAK96, JLP99, KLY93, LWMG95, LPJI98, PN99, WEKM97, AP91]. synchronization [KP90, RRR93]. synergy [DN97]. Syntactic [Che94b]. syntax [Luk94, Osb91, Sch95]. synthesis [Got99a]. Synthesizer [KJ92]. Synthesizer-Generator [KJ92]. System [Hol97, SHY96, BYC98, Ber94, Car95, Car94, CAH90, Coe93, Cun94, Cur97, DT98, FP95, Gol98, GZ94, GdLC94, Güv95, HMOS90, HM91, HL98, Jan97, Jef91, Jon96b, KOC91, LLP95, LHA90, MO96, MPTW92, MD99, NSTT96, PD96, PMCGJP97, PS98, PR98b, PWST94, PSZ99a, SKL98, SBD99, TSG98, Tos93, Wan98, Whi98, Zie99]. system-based [PD96]. Systems [EO95, RL93, SK98, And90b, BM96, Bar92, BKS^{+}99, CO95, Cam94, CLL^{+}96, CGJ93, Cla98a, CDF^{+}97, Din91, Don90, Dow99, Eri96, Ewi93, Fri91, GL94, Goh92, Gol96a, Gol97a, GBC^{+}99, GCD^{+}94, Gor98b, Gor99b, Gor99c, HSC^{+}90, Har94a, HH95a, Har90, Har92, Har94b, Har96, Hol97, Hug92, HWS92, Kou98, LM91, Lar97, Lea90, LC96a, LGK^{+}93, LSS^{+}96, Mac97, MK99, McC91, McD96, MSSZ97, MH91, Moe95, MD99, NW92b, NP92, NW99b, O’G98, Oml99, Osb92b, PD96, Pil95, Pre94, PR98b, RSBAS98, Ree90, Ric98, RR99, She98, Shu90, Ste91, Ste94, Tao95, Tay97b, Tos93, WR97, WGD^{+}95, Whi98, WAK96, WB92, WR90, Wol92, bY90, dC99].

tables [Cla93]. TAC [Pil95]. tail [Rev98]. tail-recursion [Rev98]. tails [Cas99, DKT95]. take [CR98]. taking [MRS93]. tale [BSM^{+}94, WB93]. Tanenbaum [Don95]. Tanzania [JPT^{+}99]. taped [Tow96]. tapestries [AAK^{+}94]. target [GZ94]. targeted [Smi98]. TAs [Kay95]. task [AJ93, PBEC93, RSLD99]. tasking [KP90]. taught [DH94, HW90, KCA99]. Taxman [Tro94c]. taxonomy [HCL96]. teach [Ape99, AB99, Bre95, CGP90, CLL^{+}96, DMK97, GNR^{+}95, Har96, JPPFPMVI97, Kay98a, KJ92, Lev99c, Mar98a, MBC^{+}96, NC99a, PM97, PRF96, RS94, Ros96, SC94]. teacher [GBB^{+}90, HG99, SDJ^{+}94, Tay96, Tay97a]. teachers [Kus94, LW97, SW93, TM92]. Teaching [Ågr99, Ada96a, Add97, Aik99, Arn94, Bal90, Bam98, Bel96, BG93, BFG^{+}94, Bie93, BG97, BJB98, Car99b, Car97, Cla98b, Cun97, Cyn90, DGEST96, Dav90, Det90, DDG97, Dow99, FBR^{+}99, FCB^{+}99, Fel90, FHS^{+}94, FG91, FG92, FP95, GGT98, GS99, GT98, Gre96b, GSK^{+}95, HS91, HK97, Hug92, Imp99b, JPS90, Jes94, KG94, KH92, Kir96, KFD^{+}99, KA95, KH94, KAL^{+}96, LaR93, Lia95, Lim98, McC99b, MJ99, MCD^{+}94a, McD92, MH96c, MM94a, MM94b, Muk99, Nev94, Noy93, Noy95, O’G98, OM99, Osb92a, Owe92, Par97, Pat94, Pat97, Pol99, PP97, Ree96b, Rif94b, SSH98, SG96, SG97a, SVP93, SW91, Son96, SR91, Str99, Tam92, Tom91, TW97, WB99, WWDM96, Yea91, bY90, ZZLS99, dlFCC99, Ant97, Arn95, Ast98a, BBG^{+}99, BM96, Bar93b, Bec94, Bie90]. teaching [BJ98, CCR99, CBH97, CK98, Car94, CGJ93, Cle99a, CWI96, DS98, Dal99c, DJ99, DH93, DR99, DFZ^{+}95, DV98, DFF99, EHS^{+}99, Eri93, Fri91, Goe96, GBL94, GT99b, Got98b, GKB^{+}95, GKC^{+}98, GK98, GdLC94, Gui99, Güv95, HS98, Hai96, HFKSM98, HGPJ^{+}96, Hay98, HP94, HM91, Hil93, IHA^{+}94, Jen98, JPPMMF96, JY95, Joh99b, Juh99, Kay95, Kin98, KCD^{+}96, KST92, Kno99a, KG99, KKR95, KR96a, KR99, Kro96a, Kro96b, Kum96, KM98, Lan98, LW97, LP98, Lev99b, LZ98, LAH94, LWBS99, LSS92, Mag94b, MHP93, Mar99a, MO96, MPH96, MAG^{+}94c, McC91, McC99c, McD91, MK97, McN96, MST97, Mot96, ML90, Nef93, NW99b, OM95, Osb92b, Owe90, PP95, Pat96, PRS94, Pax95, PA90, Pet98a, Pet98b, Pre94, Pro98, PSZ99b, PS91, Ree95, RD98, RLR95, Rob98]. teaching [RL95, Rod95, RBG^{+}97, RT98, SK98, SV96, SK95, SY98, SCF^{+}93, SG97b, Sta93, Sty94, SLM98, Tay99b, TGGP97, TR93, Tre99, Tym91, Utt99, Wah99, War95, Wei97, WDB98, YB99, Zac94, Zac97a, ZLS98, ZC99, ZBD99, dC99]. teaching-managing [Pat96]. Team [STBC94, BO95, BD99, Cur97, Kub99, Pig92, Pou90, SF91, SBTC94, Smi98, Sul93]. team-oriented [Pig92, SF91]. teams [Cas96, HSC^{+}90, HML^{+}99]. teamwork [Cle98a]. tech [Moe95]. Technical [ACM98, ACMJ90, BG94, Dal91b, Dal92b, IAK96, JLP99, KLY93, LWMG95, LPJI98, PN99, WEKM97, AP91, Kin98, Kou97]. technique [KST92, Mac97]. techniques [Bec94, Ell96, JMM97, KR99, Lev99c, Lia95, MC90b, PFR^{+}94, SAL^{+}90, Wal90, WB92]. technological [Fag94]. technologies [BMP96, Lim98, RFPA97, RHP90]. Technology [Imp99a, LGB^{+}99, Ada96a, ADB^{+}96, Amo93, AC96, Ast98a, Ciz97, DS99, DR99, Gro97a, HAG92, KI93, KBCP99, LC96a, MW98a, MCW99, MP96, MLS97, NSS98, OJ93, Owe96, PBEC93, PS97, RS97, Ree97, RF96, Rob98, SBD99, Tay99a, Tay99b, YSKW95, Zag98]. TechTalk [JLH^{+}98]. Telecommunications [JM91, BBJR96, Tay96]. telelearning [AAHM99]. Telematics [DG99]. teleteaching [Sch96a]. television [BJB98]. Template [CS99]. templates [Ras97b]. Tennessee [LWMG95]. terminal [GSW^{+}98]. tertiary [TC93]. test [Fre94, HTW98, TP94]. testbed [Bro93b]. Testers [BBG^{+}99]. Testing [dOG94, RK97, BBC^{+}94, Ell96, Hit99, SD93, WNKK98]. tests [Kum99]. TETE [WNKK98]. Texas [Dal91b]. text [Che94b]. text-processing [Che94b]. textbook [Gri99]. textbooks [For91a, MS97, Shi92]. Texts [Wis96, Kat95]. textual [CB94a]. their [BLOS98, BSM^{+}94, Fre94, MV98, Rad90]. them [AIM^{+}95, PRF96]. theme [Gur98, MSPF95, McN98]. Themes [AAK^{+}94]. Theorems [GR99, RG98b]. theoretical [UD97]. Theory [GR99, Ape90, Cor99, DS98, DeC96, DU96, Ele99, FP95, Goe96, Kro92, RL93, RBL^{+}97, RG98b, WV91]. There [Luk94]. thesis [Tay98]. ThetenthStrand [Tow99]. Thetis [FR96b]. thing [Joh99a]. Thinking [BAK99, BD96, BF94, FR98, MWP94]. third [MRS93]. Thirtieth [JLP99]. threads [Dal98b, HA95]. Three [Bal96b, GEC96, BSK94, Rei95b, WC94]. Three-dimensional [GEC96]. three-fold [BSK94]. throughout [DDP96, MAD^{+}95]. thunks [BG93]. TIC [Pil95]. TIC-TAC-TOE [Pil95]. tiered [WC94]. Time [KMH91, Bro93b, McC91, OM99]. times [Col91b]. Tinyturing [Lam90]. tips [AIW^{+}96]. Tkined [Jac96]. Tkined/scotty [Jac96]. TOE [Pil95]. together [BGG^{+}99, Car99c, Hed99, HCL96, McM96]. together-collaboration [Hed99]. Too [MIG^{+}94, DS99]. Tool [Ara91, ADB^{+}96, Arm99, BJ98, BV97, Car94, Che94b, CC92, DEP97, EKR98, Eri93, Gam95, GG99, Goe96, GSK^{+}95, Hug90, JLH^{+}98, Joy98, Juh99, LH92, MO96, NG94, Osb92b, Pax96, Pow99, RD98, RG98a, RHND99, RT98, Sch95, Tri99, Tym91, ZLS98, ZC99]. toolbox [AW93, AWB94, SG91]. toolkit [BC96, Rei90]. Tools [DB98, KR99, NS96, OM95, Rob96, Bei96, Bie93, BJR94, Cun96, GL96, Jac96, KAL^{+}96, McC98b, Reb98, RBL^{+}97, SH99a, Sch92, Seb97, SWWB97, TR93]. toolset [VIRF96]. Top [DH93, DH94, Hil93, LZ98, NC99b, Ree95]. Top-down [DH93, Hil93, LZ98, Ree95]. TOPAS [Ber94]. topics [Dra98, Kin98, SW93, TBF98, Tri99]. total [Ske93]. Tournament [PLU97, Res96]. Tower [CJ91]. toy [ADB^{+}96]. TQM [Nul96]. tracing [MPH96, Sch90]. track [AAW92, Cup98, MR99]. track-applied [Cup98]. tracking [GZ94]. traditional [CO95, CP96, Ell98b, Ger94, Joh97, Joh99b, MLS97, OBH^{+}95, SBD97]. Traffic [Pro98]. Trainer [Web96]. Training [Kay95, Goe96, HG99, Lar97, Zag98, BHH^{+}99]. transfer [FP98, SH91]. transformation [WUGV99]. transition [HS93, SFR^{+}98]. Translation [Pio92]. translations [Che94a]. Transparent [Bre95]. transputer [MH91]. traveled [LN97]. treasurer [Wal99c, Wal95]. tree [Gur95, Ras97c, Rug99]. trees [Ada96b, BD96, Cha91a, Mal97, Ras97c, Zie98]. trends [DM92, NC99b]. trial [Cle99a]. tricks [Wil95a]. trimming [KLMS95]. trip [McC99a]. troubleshooting [Jac96]. truth [Cla93]. TU [GS96]. Turbo [Dix91, Dix93, Hug92, MC90a, Smi94b, VI96]. Turbo-like [VI96]. Turing [Bro97, Han92, Hol94, RHND99, Tay98, TP94]. Turning [Tow98, ADB^{+}96]. tutor [Mit98]. Tutorial [CB94b, Goe96, Tro94a, Zac94, BKCG98, BV97, Coe93, Ell98b, Hug90, MBJS90, PWST94, RT98, Wal90]. Tutorial-based [Zac94]. Tutorial-formal [Tro94a]. tutorials [CWS98, RM96]. tutoring [Güv95, JD97a, KOC91]. twenty [Tay98]. twenty-first [Tay98]. Two [Dal98b, Fag94, Got99b, WW91, Ack99, AJ93, BSM^{+}94, CD95, ICD^{+}90, JM91, KLE92, KI93, Kle99a, Tro91, WB93]. two-year [AJ93, CD95, ICD^{+}90, KI93, Kle99a]. Tying [BGG^{+}99, HCL96]. type [FJM98, Yer94]. types [GPJ95, Gre90, HSS98, MC90a, SH95, Sil94].

U.S [FHL^{+}97]. ugly [Car99b]. Umbriel [Ter95]. Umbriel-imperative [Ter95]. UML [Arm99]. Undergraduate [Cas98c, CEMP97, FMM97, Har97, KLE92, KPW94, Poo91, Tro92, ABFM90, AAW92, Alm91, BBJR96, Bar93a, BJB98, Cas98b, CHLS95, CDF^{+}97, FB97, Fel99, FC94, For91b, For94, GDC^{+}97, Had98a, HKH^{+}93, Har92, HW90, Har99c, Has94, How93, HK96, JMM97, Joh94, Joh99b, Kay93, KKA98, Leb99, LM94, MZS^{+}95, MVS^{+}98, MBJS90, Maz93, MAD^{+}95, MJM96, McD91, Mer92, MH91, MPHL95, Mur91, ML90, NI98, OJ93, Owe96, PP95, PZ96, Pil93, Pra92, Pre96, Ree90, RR99, Sap91, SS98, Seb97, SCF^{+}93, Smi91, SG99b, Ste94, TG94, Tew95, Thu90, Tol95a, Tre98, TLR94, WCDT95, WG93, WGD^{+}95, bY94]. undergraduates [Cun95, Hol94, Kea93, LM91, LEDW93, MM93, RLR95, Shu90, Tym91, WBBW94]. understand [HP93, Tao95, ULW99]. Understanding [Bar97a, KW94, BCPV98, NSS98, WAK96]. unified [Eva95, Jar94]. uniform [Gur95]. unifying [Joy98, WDM97]. unit [AMN98]. United [HWS92, Tay99a]. universal [RBG^{+}97]. universities [KLE92, Lev99a]. University [BABT97, Car98a, Car98b, CH91, TM92, VL90, Wil97, BR99, Dav98, DG99, GS93, SLM98]. UNIX [Dun91, Gre96b, Mac97, Osb92a, Ada93b]. unlevel [SP96]. unrestricted [MK99]. unsophisticated [Hay95, Ter95]. unum [dVdVI97]. update [RLR95, Ros91b]. updates [Har99c]. UPE [Rog90]. Upgrading [Pra90]. upheaval [SVFB94]. upper [Kay98a]. Uppsala [Dal99c]. upside [Scr91]. Upsilon [Rog90]. USA [ACMJ90, BG94, Dal91b, Dal92b, IAK96, KLY93, LWMG95, LPJI98, PN99, WEKM97, ACM98]. usability [FC99a]. usage [FRK96, Gre96a, Lev95, O’N95, PCLM97]. Use [BM96, Ell98b, KWJ^{+}96, Lev99b, PM97, Bar92, BBG^{+}96, Car95, DH97, Ell98a, FR96b, Ful91, GPP99, Gol97b, Gol99, JF97, KS94, KOC91, McC98a, McC96, Moo95, PL96, PSZ99a, PSZ99b, RL93, SH95, SH91, Smi98, Sta93, Sty94, UD97]. used [CK98, Rad90]. useful [Ber90]. USENET [Eri93]. User [Cun91, McD96, Bar93b, FC99a, Mes91, WH96]. User-level [McD96]. users [LM98b]. uses [Mye94, Wal90]. Using [Ape99, Ash97, Ast98b, BK92a, BA96b, Ber94, Ber98, BDN^{+}94, CS98, CK99, CD95, CWS98, Con98a, Cun94, DMK97, FR98, GT99b, GR99, GL97, GNR^{+}95, Gro98, HML^{+}99, Hug90, HK96, Jef91, Juh99, Kho90, Kro96a, Kro96b, KJ92, Lea95, Leo91, Mar98a, McK95, McN96, McN98, MHP98, MP96, ML96, Nap97a, NC99a, NF97, Noh94, PFR^{+}94, RFH^{+}93, RS94, Rob93, RLR95, RFPA97, RM98, Sab90, SBB^{+}97, SHY96, Shi93b, SC94, Sil91, Sta97, Tao95, Tay97b, Tos93, TZ99, Van98, VW96b, Wil90, Wit96, YB99, YSKW95, Ågr99, Arn94, AB99, ASW97, BHH^{+}90, BG93, BJB98, Bur97, Cas96, Cle99a, Coe93, DH90, Duc94, FP98, GBB^{+}90, GdLC94, HGPJ^{+}96, Har96, Hay97, Hol94, Hug92, JU97, JLH^{+}98, Kam92, KW99, KH94, Koo95, KP90, Lea90, Lea97, Lia95]. using [LAH94, MBC^{+}96, Muk99, NSTT96, Owe92, Pat97, Pax96, PR98a, Pol99, Rei95a, RK98, Rob98, Sap91, SW93, SBD99, SAL^{+}90, SR99, Slo95, Smi94b, SBB^{+}99, Str99, Szu96, TH98, Van95, WW91, Wic95, WGHL97, YLQ98, bY94]. Utilising [WPI98]. utility [GT99a]. Utilizing [MH91, MB95].

vadis [GC98]. Validation [KFD^{+}99]. value [HS98]. values [DSB94]. variables [O’N95, Wha91]. variations [MSPF95]. VCOCO [RD98]. vectors [Hol97]. verification [Fek93, Wal90, Wha91]. verifiers [Kos99]. version [Don95]. versus [HH95c, LP98, Ree96b]. Very [Pil93, HL94, Imp99b]. VI [KD90]. via [Car97, Cla98b, EL97, Gam95, GY98b, Hig93, Sai97]. VIBDaST [TGGP97]. Victoria [SH97, BABT97]. video [Mil96, SRJ99, Tow96]. video-taped [Tow96]. view [GYT^{+}93, Har98c, Jar94, Joy98, LWBS99, Whi98]. Viewing [Tow96]. views [FHL^{+}97, Uck91]. ViLi [SV96]. VINCE [RT98]. VINCE-an [RT98]. Virtual [Car98b, FST98, JLH^{+}98, Ågr99, Bel96, Fal99, FH91, Gol98, HG98, Hou97, Joy90, SVFB94, SW99, TGGP97]. visible [GMF99]. Vision [SV96]. Visions [BMB^{+}95]. Visual [KSW98, SW95, SWWB97, WSK97, Bar95, BJR94, Cla93, GT98, GNR^{+}95, HBC97, JPKPM99, NS96, Rob92, Tja96, Cas96]. visualisation [CH98, FR98, KPT99, RD98]. visualisations [Nap98]. visualiser [Nap98]. Visualization [McN94, SHY96, Ash97, AR98, BBG^{+}96, BWW96, DV98, EKR98, GEC96, JPPFPMVI96, McN96, Nap90, NS94, Nap96, Nap97b, Nap97a, NB98, NC99a, SKL98, Sch92, TR93, BBG^{+}96]. visualization-based [Nap97a]. visualization-easy [NS94]. visualization-some [BWW96]. visualizations [LK95]. visualizer [JPPM96]. visualizers [BBG^{+}99]. Visualizing [KH99, WLC96, GZ96, Jan97, Juh99]. VLSI [CM96, Wal99d]. VNET [Tym91]. volleyball [Wal97b]. voyage [Rad93]. vs [CB94a, HW90, Sch96b, Thw94, Zai96]. VSC [Pil93].

wait [Har99a]. want [Pfe99]. wanted [KG94, MHB^{+}98]. wants [YDM99]. Waring [Hex94]. Warning [McC98b]. Washington [ACMJ90]. Watson [OK95]. way [Coh99, DKT95, Dor92, Joh99a, MST97]. weakest [CHL96]. web [Whi98, Asl98, Alm98, AC96, AB99, Blu96, BGGR98, BGG^{+}99, Cas98a, CWS98, CS99, Cul98, Cur97, DB98, Ell98b, EL97, FC99b, GMF99, Gol97b, HGPJ^{+}96, Har99b, Hei99, Hol98a, JPT^{+}99, JLH^{+}98, JRS96, Kno96, Kos99, Kum98, LG96, Law97b, LN98, Lim98, LH99, MH96a, McC98b, Nap96, Nap97b, NB98, Owe96, Pax96, PR98a, RR97, RK98, RM96, RBG^{+}97, SPV99, VW96a, VW96b, WB99, YLQ98]. WEB-based [Asl98, BGG^{+}99, DB98, Gol97b, Hol98a, JLH^{+}98, JRS96, LG96, LH99, PR98a]. WebCT [Gol97b]. WebLab [RBG^{+}97]. webs [Reb98]. Websites [Eng99a]. Webucation [Pax96]. Webware [FC99b]. welcome [CEMP97]. Wellington [BABT97]. were [Dal90, KG94, MHB^{+}98]. Westwood [DH96c]. Where [AIM^{+}95, EBD^{+}90, LGK^{+}93, Wer99]. which [AR99, Ber90, CB94a]. WHILE [Arn91, Col91a]. whimsical [Jul92]. white [CH91]. Who [YDM99, KACL94]. Wide [Ell98b, HGPJ^{+}96, Kno96, LN98, MH96a, Nap96, Nap97b, NB98, Owe96, RM96, RBG^{+}97, PSZ99a, Tho98b, TLNS98]. will [Kno98a]. windowed [NB98]. Windows [Osb95, Szu96, Cas96, DEP97]. Windows-based [Cas96, DEP97]. WinRDBI [DEP97]. wireless [SR99]. within [AFVH97, FC99a, Gol99, MO96, RS91, WBCS95, Wal98b, Wal98c]. without [Add97, BG97, Wad95]. Witwatersrand [GS93]. Women [BABT97, KV94, KW93, SRY98, AIM^{+}95, ABM93, Ber91, BCPV98, EBD^{+}90, FMM97, HF98, HKH^{+}93, How93, MIG^{+}94, Mye92, SP96, SS98, TC93, VBB^{+}98]. word [Imp99b]. work [Car99c, FP98, HT97, Oli94]. workbench [BBF^{+}94]. worked [SG96]. Working [BKP^{+}99, FBR^{+}99, KFD^{+}99, LGB^{+}99, ADB^{+}96, BN98, Ell98a, GY98a, GK98, HGPJ^{+}96, KWJ^{+}96, MR98, Tho98b, ACK^{+}96, BBG^{+}96]. Workplace [LGB^{+}99]. works [NSS98]. workshop [AIW^{+}96, Dor97, HG99, Sai97]. workstation [Shu90]. World [Fag94, NB98, BD96, Cun98, DP90, HW90, MRS93, Nac98, VB98, WWDM96, Ell98b, HGPJ^{+}96, Kno96, LN98, MH96a, Nap96, Nap97b, Owe96, RM96, RBG^{+}97]. worst [TS95]. worst-cases [TS95]. would [Dal90]. Writing [Cur92, FPC96, Wal98c, BK92b, BJ98, CC92, Cun94, Els92, FK92, Haf94, JPS90, Kay98a, Pes91, Rei90, TP93, Van95]. written [Kum99, Sul93]. WW [Pli98]. WWW [CGK^{+}96, DH97, HK97, Mak98, Nap97a, NSTT96, Sch96a, TBF97, WP98]. WWW-based [Mak98, TBF97].

XDP [Arn95]. XINU [Car95]. XINU-an [Car95]. XTANGO [Har94b].

Y2K [HRBK99]. YAATCE [Wah99]. YAATCE-yet [Wah99]. yacc [LK95]. YAP3 [Wis96]. YAP’ing [Wis92]. year [And90a, AJ93, BMC95, Bal96b, BMR^{+}95, BD93, CBH97, CD95, CDF^{+}97, DH96b, DM92, Dor97, EHS^{+}99, Fel90, Gol97b, HSM97, HH95a, ICD^{+}90, Kay92b, KI93, Kle99a, KKR95, LaR93, LP98, MH96c, MRS93, Ree96b, Ric98, Tea92, Ger98]. years [Fag94, NSS98, SK95]. York [Bri91]. Youngster [STM95].

Z [Sai97]. Zealand [Cle99a]. Zen [Dal91a].

References

Akhtar:1999:NCO

[AAHF^{+}99] Shakil Akhtar, Nizar Al-Holou, Mark Fienup, Gail T. Finley, Robert S. Roos, and Sam Tannouri. The networks course: old problems, new solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):360–361, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arco:1999:QSE

[AAHM99] J. M. Arco, B. Alarcos, A. M. Hellín, and D. Meziat. Quality of service over Ethernet for telelearning applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):68–70, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1994:TTD

[AAK^{+}94] David Arnow, Owen Astrachan, James Kiper, Robert Workman, Paula Whitlock, Brent Auernheimer, and John Rager. Themes and tapestries: a diversity of approaches to computer science for liberal arts students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 374–375, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1992:DLM

[AAW92] Robert M. Aiken, Dean Allemang, and Thomas Wehrle. Designing laboratory modules for novices in an undergraduate AI course track: artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1): 76–80, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1999:LPE

[AB99] David Arnow and Oleg Barshay. On-line programming examinations using Web to teach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):21–24, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1990:SNB

[ABDM90] Robert M. Aiken, Della T. Bonnette, Lionel E. Deimel, and James E. Miller. SIGCSE’s newsletter — the Bulletin from 1969 to 1990 (special panel session): the Bulletin from 1969 to 1990 (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):253, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abunawass:1990:FUC

[ABFM90] Adel M. Abunawass, Omar Bukhres, Theresia G. Fisher, and Kenneth Magel. A first undergraduate course in neural networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):240–244, February 1990. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1993:IMW

[ABM93] Vicki L. Almstrum, Anita Borg, and J. Paul Myers, Jr. Improving mentoring for women in computer science fields (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):294, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abunawass:1992:BBM

[Abu92] Adel M. Abunawass. Biologically based machine learning paradigms: an introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1): 87–91, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1993:RMC

[ABW93] Vicki L. Almstrum, Debra Burton, and Cheng-Chih Wu. Research methods in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):306, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1996:ECI

[AC96] David M. Arnow and Dayton Clark. Extending the conversation: integrating email and Web technology in CS programming classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):93–95, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Auernheimer:1990:NNP

[ACE^{+}90] Brent Auernheimer, Vint Cerf, Susan Estrada, Russ Hobby, Craig Partridge, Eugene Spafford, and Steven S. Wolff. Networking in the nineties (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 252, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1996:IMP

[ACK^{+}96] Elizabeth S. Adams, Linda Carswell, Amruth Kumar, Jeanine Meyer, Ainslie Ellis, Patrick Hall, and John Motil. Interactive multimedia pedagogies. Report of the Working Group on Interactive Multimedia Pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):

182–191, 1996. CODEN SIGSD3. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/Pape.bib.

Ackermann:1999:PSI

[Ack99] Ernest Ackermann. Poster session. Internet curriculum. two courses: introductory and advanced. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 179–180, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ACM:1998:PST

[ACM98] ACM, editor. Proceedings of the 29th SIGCSE Technical Symposium on Computer Science Education (SIGCSE-98), 25 February–1 March 1998, Atlanta, GA, USA, volume 30(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Austing:1990:PST

[ACMJ90] Richard H. Austing, Lillian N. Cassel, James E. Miller, and Daniel T. Joyce, editors. Proceedings of the 21st SIGCSE Technical Symposium on Computer Science Education, 1990, Washington, DC, USA, 1990, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1990. CODEN SIGSD3. ISBN 0-89791-346-9. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1997:RAP

[ACRW97] Owen L. Astrachan, Gail Chapman, Susan H. Rodger, and Mark Allen Weiss. The reasoning for the advanced placement C++ subset. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):62–65, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1993:PIS

[Ada93a] Evans J. Adams. A project-intensive software design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):112–116, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1993:DIU

[Ada93b] Joel C. Adams. The design and implementation of a Unix classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):223–227, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1996:TTM

[Ada96a] Elizabeth S. Adams. Teaching with technology at my fingertips. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):229, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1996:KYR

[Ada96b] Joel C. Adams. Knowing your roots: object-oriented binary search trees revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):36–40, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1996:OCD

[Ada96c] Joel C. Adams. Object-centered design: a five-phase introduction to object-oriented programming in CS1-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):78–82, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1998:CIO

[Ada98] Joel C. Adams. Chance-it: an object-oriented capstone project for CS-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):10–14, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1996:ETT

[ADB^{+}96] Vicki L. Almstrum, Nell Dale, Anders Berglund, Mary Granger, Joyce Currie Little, Diane M. Miller, Marian Petre, Paul Schragger, and Fred Springsteel. Evaluation: turning technology from toy to tool: report of the working group on evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):201–217, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Addison:1997:TCS

[Add97] Paul Addison. Teaching computer science without a computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):30–33, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1997:DMC

[AFVH97] William Joseph Adams, Edward A. Fox, Gloria Melara Vides, and Rachelle Heller. Defining multimedia courses within a computer science education (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):379–380, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Andreeva:1996:IAM

[AG96] Elena Andreeva and Igor Gushchin. Informatics in an advanced mathematical high school. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):231, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aagren:1999:TCC

[Ågr99] Ola Ågren. Teaching computer concepts using virtual machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):84–85, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abernethy:1998:IBC

[AGT98] Ken Abernethy, Paula Gabbert, and Kevin Treu. Inquirybased computer science instruction: some initial experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):14–17, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1993:HCM

[AHST93] Elizabeth S. Adams, Rachelle S. Heller, Elliot Soloway, and Barbee Mynatt Teasley. Here comes the multimedia generation! ready? or not? (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 297, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1999:SAN

[Aik99] Robert M. Aiken. A skeptic assesses the Net as a resource for teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):79–82, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:1995:WWG

[AIM^{+}95] Sandra Honda Adams, Rachel D. Isles, Louise E. Moses, Domenick Pinto, and Frances L. Van Scoy. Where have the women gone and how do we keep them from going? (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):360–361, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:1996:LLT

[AIW^{+}96] R. Aiken, G. Ingargiola, J. Wilson, D. Kumar, and R. Thomas. Lessons learned: tips from conducting an NSF faculty enhancement workshop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):3–7, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Austing:1993:FRC

[AJ93] Richard Austing and Therese Jones. Final report: curricula for two-year college task force subgroup: computing for other disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):310, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1991:RBP

[Alm91] Vicki L. Almstrum. The relationship between pre-college mathematics and the undergraduate computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):124–129, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1998:WSP

[Alm98] Vicki L. Almstrum. A Web site to promote CS education as a research area. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):42–45, December 1998.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arras:1990:SSC

[AM90] Richard J. Arras and Lewis Motter. The senior seminar in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):29–36, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aman:1990:CSC

[Ama90] James R. Aman. Computer science in correctional education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):147–151, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1998:DPE

[AMBC98] Owen Astrachan, Garrett Mitchener, Geoffrey Berry, and Landon Cox. Design patterns: an essential component of CS curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):153–160, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abunawass:1998:IUB

[AMN98] Adel Abunawass, Martin Maskarinec, and Kathleen Neumann. An integratable unit based computer science distance learning curriculum design for the ACM/IEEE curricula 1991. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):18–20, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amoroso:1993:GCC

[Amo93] Edward G. Amoroso. A graduate course in computing security technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):251–255, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:1990:MFF

[And90a] Oliver D. Anderson. My five favorite first year Pascal programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):60–64, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Andrianoff:1990:MDS

[And90b] Steven K. Andrianoff. A module on distributed systems for the operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):176– 180, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Antoniou:1997:NMT

[Ant97] Grigoris Antoniou. A new methodology for teaching default reasoning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2):35–38, June 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1991:SPS

[AP91] Vicki L. Almstrum and David C. Platt. Summary of postsymposium evaluations from 1991 SIGCSE technical symposium on computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (4):27–30, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Apenyo:1990:DST

[Ape90] Kofi Apenyo. A database sequence: theory then practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):5–10, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Apenyo:1999:UER

[Ape99] Kofi Apenyo. Using the entity-relationship model to teach the relational model. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):78–80, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Allen:1990:RIC

[APNA90] J. Thomas Allen, Hayden Porter, T. Ray Nanney, and Ken Abernethy. Reexamining the introductory computer science course in liberal arts institutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 100–104, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1995:ACA

[AR95] Owen Astrachan and David Reed. AAA and CS 1: the applied apprenticeship approach to CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1): 1–5, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1998:AVI

[AR98] Owen Astrachan and Susan H. Rodger. Animation, visualization, and interaction in CS 1 assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):317–321, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abi-Raad:1999:CEW

[AR99] Maurice Abi-Raad. Code of ethics? which one? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):73–77, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arab:1991:TMP

[Ara91] Mouloud Arab. Tool for making programs more readable. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):31–35, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armentrout:1999:TDJ

[Arm99] Anna Armentrout. A tool for designing Java programs with UML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):180, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1991:IWL

[Arn91] David Arnow. The iliad and the WHILE loop: computer literacy in a liberal arts program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):78–81, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1994:TPL

[Arn94] David Arnow. Teaching programming to liberal arts students: using loop invariants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):141–144, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arnow:1995:XSL

[Arn95] David M. Arnow. XDP: a simple library for teaching a distributed programming module. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1): 82–86, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ashton:1997:UIN

[Ash97] Paul Ashton. Using interaction networks for visualization of message passing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):272–276, March 1997. CODEN SIGSD3. ISBN 0-89791-889-4. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1997.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Os/minix.bib.

Aslam:1998:WBQ

[Asl98] Sohail Aslam. WEB-based query processing in a database course project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):297–301, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1991:PI

[Ast91] Owen Astrachan. Pictures as invariants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):112–118, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1992:FSR

[Ast92] Owen Astrachan. On finding a stable roommate, job, or spouse: a case study crossing the boundaries of computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):107–112, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1994:SRI

[Ast94] Owen Astrachan. Self-reference is an illustrative essential. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):238–242, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1998:CTH

[Ast98a] Owen Astrachan. Concrete teaching: hooks and props as instructional technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):21–24, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1998:UCE

[Ast98b] Owen Astrachan. Using classes early, an object-based approach to using C++ in introductory courses (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):383, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:1997:ABM

[ASW97] Owen Astrachan, Robert Smith, and James Wilkes. Applicationbased modules using apprentice learning for CS 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):233–237, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1993:CTR

[AW93] Vicki L. Almstrum and Cheng-Chih Wu. The CSedRes toolbox: a resource to support research in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):21–26, December 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:1994:RRC

[AWB94] Vicki L. Almstrum, Cheng-Chih Wu, and Debra Burton. A resource for research in computer science education: the CSedRes toolbox (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):409, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1996:SEL

[BA96a] Mordechai Ben-Ari. Structure exits, not loops. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):51, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1996:UII

[BA96b] Mordechai Ben-Ari. Using inheritance to implement concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):180–184, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1997:DAJ

[BA97] Mordechai Ben-Ari. Distributed algorithms in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):62–64, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1998:CCS

[BA98] Mordechai Ben-Ari. Constructivism in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):257–261, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1997:WIC

[BABT97] Judy Brown, Peter Andreae, Robert Biddle, and Ewan Tempero. Women in introductory computer science: experience at Victoria University of Wellington. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):111–115, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1997:CAP

[BAH97] Mordechai Ben-Ari and Kevlin Henney. A critique of the advanced placement C++ subset. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 7–10, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bailie:1991:IMA

[Bai91] Frances K. Bailie. Improving the modularization ability of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):277–282, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1999:TPP

[BAK99] Mordechai Ben-Ari and Yifat Ben-David Kolikant. Thinking parallel: the process of learning concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):13–16, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1990:TIC

[Bal90] Doug Baldwin. Teaching introductory computer science as the science of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):58–62, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1996:DLC

[Bal96a] Doug Baldwin. Discovery learning in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):222–226, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1996:TYE

[Bal96b] Doug Baldwin. Three years’ experience with gateway labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):6–7, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1997:FPC

[Bal97] Doug Baldwin. Finding a place for computer-equipped lectures in a lab-rich environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):34–37, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Balbin:1999:YDQ

[Bal99] Isaac Balbin. Is your degree quality endorsed? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):60–63, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bampton:1998:TCS

[Bam98] Alvin Bampton. Teaching computer science in Papua New Guinea. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):25–27, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bareiss:1992:ICU

[Bar92] Catherine C. Bareiss. An introductory course on the use of operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):173–175, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnett:1993:EPS

[Bar93a] B. Lewis Barnett III. An Ethernet performance simulator for undergraduate networking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):145– 150, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barrett:1993:HMT

[Bar93b] Martin L. Barrett. A hypertext module for teaching user interface design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):107–111, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnett:1995:VSS

[Bar95] B. Lewis Barnett III. A visual simulator for a simple machine and assembly language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):233–237, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bareiss:1996:SPC

[Bar96a] Catherine C. Bareiss. A semester project for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):310–314, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barrett:1996:EDC

[Bar96b] Martin L. Barrett. Emphasizing design in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):315–318, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:1997:SAQ

[Bar97a] David Barnes. Students asking questions: Facilitating questioning aids understanding and enhances software engineering skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):38–41, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Techreports/UKent.bib; http:
//www.cs.kent.ac.uk/pubs/1997/524.

Barrett:1997:SRG

[Bar97b] Martin L. Barrett. Simulating requirements gathering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):310–314, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:1999:DED

[BAS99] Mordechai Ben-Ari and Shawn Silverman. DPLab: an environment for distributed programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):91–94, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:1994:CTB

[BBC^{+}94] Keith Barker, Andrew P. Bernat, Robert D. Cupper, Charles F. Kelemen, and Allen B. Tucker. Class testing the breadthfirst curriculum (abstract): summary results for courses I-IV. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):405–406, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burd:1995:RCS

[BBD^{+}95] Barry Burd, J. Glenn Brookshear, Rick Decker, Frances G. Gustavson, Mildred D. Lintner, and Greg W. Scragg. Revitalizing the computer science course for non-majors (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):388–389, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berque:1994:KWS

[BBF^{+}94] Dave Berque, Jeff Bogda, Brian Fisher, Tim Harrison, and Nick Rahn. The KLYDE workbench for studying experimental algorithm analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):83–87, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Birch:1995:DDC

[BBG^{+}95] Michael R. Birch, Christopher M. Boroni, Frances W. Goosey, Samuel D. Patton, David K. Poole, Craig M. Pratt, and Rockford J. Ross. DYNALAB: a dynamic computer science laboratory infrastructure featuring program animation (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):29–33, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1996:OVU

[BBG^{+}96] Joe Bergin, Ken Brodlie, Michael Goldweber, Ricardo Jiménez-Peris, Sami Khuri Marta Patiño-Martínez, Myles McNally, Tom Naps, Susan Rodger, and Judith Wilson. An overview of visualization: its use and design. Report of the Working Group on Visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):192–200, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927

(electronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/Pape.bib.

Baker:1999:TVT

[BBG^{+}99] Ryan S. Baker, Michael Boilen, Michael T. Goodrich, Roberto Tamassia, and B. Aaron Stibel. Testers and visualizers for teaching data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):261–265, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnard:1996:CSU

[BBJR96] Anthony C. L. Barnard, Barrett R. Bryant, Warren T. Jones, and Kevin D. Reilly. A computer science undergraduate specialization in telecommunications and computer networking.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):324–328, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bishop-Clark:1992:PAN

[BC92] Catherine Bishop-Clark. Protocol analysis of a novice programmer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):14–18, September 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bynum:1996:AYA

[BC96] Bill Bynum and Tracy Camp. After you, Alfonse: a mutual exclusion toolkit. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):170–174, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brooks:1994:APM

[BCHB94] Marc W. Brooks, John R. Chang, Ryan C. Horner, and David W. Brown. Analysis of personal mail attributes in a computer mediated communication environment (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):410, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bjorkman:1998:EPT

[BCPV98] Christina Bjorkman, Ivan Christoff, Fredrik Palm, and Anna Vallin. Exploring the pipeline: towards an understanding of the male dominated computing culture and its influence on women. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):64–69, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bryant:1993:FCC

[BD93] Robert Bryant and Paul De Palma. A first course in computer science for small four year CS programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):31–34, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berman:1996:TAB

[BD96] A. Michael Berman and Robert C. Duvall. Thinking about binary trees in an object-oriented world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):185–189, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berglund:1997:IEQ

[BD97] Anders Berglund and Mats Daniels. Improving education quality, a full scale study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):330–334, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1999:SET

[BD99] Judy Brown and Gillian Dobbie. Supporting and evaluating team dynamics in group projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):281–285, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berman:1994:UCC

[BDN^{+}94] A. Michael Berman, Rick Decker, Dung X. Nguyen, Richard J. Reid, and Eugene Wallingford. Using C++ in CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):383–384, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:1994:ITT

[Bec94] Byron Weber Becker. Inexpensive teaching techniques with rich rewards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):41–44, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boroni:1996:DDE

[BEG^{+}96] Christopher M. Boroni, Torlief J. Eneboe, Frances W. Goosey, Jason A. Ross, and Rockford J. Ross. Dancing with DynaLab: endearing the science of computing to students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):135–139, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beise:1996:IIT

[Bei96] Catherine M. Beise. Integrating Internet tools into the soft side of informatics education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):107– 108, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bellin:1992:SCO

[Bel92] David Bellin. A seminar course in object oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):134–137, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bell:1996:TVR

[Bel96] D. H. Bell. Teaching virtual reality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2): 56–61, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benham:1992:GAP

[Ben92] James W. Benham. A geometric approach to presenting computer representations of integers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):27– 28, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benham:1995:GAI

[Ben95] James W. Benham. A geometric approach to integer condition codes and branch instructions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):35–38, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1990:CAW

[Ber90] Joseph Bergin. Coatroom: an ADT which is useful in implementation of object oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):45–ff., September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernstein:1991:CEC

[Ber91] Danielle R. Bernstein. Comfort and experience with computing: are they the same for women and men? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):57–60, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berman:1992:CDC

[Ber92a] A. Michael Berman. Class discussion by computer: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):97–101, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berry:1992:SPC

[Ber92b] Michael W. Berry. Simulating procurement in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):15–19, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergmann:1993:SCH

[Ber93] Seth D. Bergmann. Simulating and compiling a hypothetical microprogrammed architecture with projects for computer architecture and compiler design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2): 38–42, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berard:1994:UTS

[Ber94] Anthony D. Berard, Jr. Using the TOPAS solid modeling and animation system to animate science models (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):409, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berk:1996:SSE

[Ber96] Toby S. Berk. A simple student environment for lightweight process concurrent programming under SunOs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):165–169, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernstein:1997:CDC

[Ber97a] Danielle R. Bernstein. Computing, diversity and community: fostering the computing culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 101–105, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berry:1997:IDM

[Ber97b] Jonathan Berry. Improving discrete mathematics and algorithms curricula with LINK. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):14–20, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergmann:1998:UIC

[Ber98] Seth Bergmann. Using inheritance for computing with physical quantities: an interdisciplinary project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):45–47, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bloom:1994:SSC

[BF94] Bard Bloom and Alan Fekete. Self-sufficiency and critical thinking in the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):9–18, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biermann:1994:THM

[BFG^{+}94] Alan W. Biermann, Amr F. Fahmy, Curry Guinn, David Pennock, Dietolf Ramm, and Peter Wu. Teaching a hierarchical model of computation with animation software in the first course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):295–299, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1993:TPP

[BG93] John Bergin and Stuart Greenfield. Teaching parameter passing by example using thunks in C and C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):10–14, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beck:1994:PST

[BG94] Robert Beck and Don Goelman, editors. Proceedings of the 25th SIGCSE Technical Symposium on Computer Science Education, 1994, Phoenix, Arizona, USA, March 10–12, 1994, volume 26(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1994. CODEN SIGSD3. ISBN 0-89791-646-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biffl:1997:TSE

[BG97] Stefan Biffl and Thomas Grechenig. Teaching software engineering and project management to 300 participants without drain of quality or intensity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):51–53, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:1997:DEP

[BGEL^{+}97] Keith Barker, Judith Gal-Ezer, Pam Lawhead, Kurt Maly, James E. Miller, Pete Thomas, and Elizabeth S. Adams. Distance education (panel): promise and reality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):369–370, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boroni:1999:TIA

[BGG^{+}99] Christopher M. Boroni, Frances W. Goosey, Michael T. Grinder, Jessica L. Lambert, and Rockford J. Ross. Tying it all together: creating self-contained, animated, interactive, Web-based resources for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):7–11, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boroni:1998:PSI

[BGGR98] Christopher M. Boroni, Frances W. Goosey, Michael T. Grinder, and Rockford J. Ross. A paradigm shift! The Internet, the Web, browsers, Java and the future of computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):145–152, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baxter:1990:ICS

[BHH^{+}90] Nancy Baxter, David Hastings, Jane Hill, Peter Martin, and Robert Paul. Introduction to computer science: an interactive approach using ISETL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):31–33, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bagert:1999:RCS

[BHH^{+}99] Donald J. Bagert, Thomas B. Hilburn, Gregory W. Hislop, Nancy R. Mead, Susan A. Mengel, and Hossein Saiedian. A report on the 1999 Conference on Software Engineering Education and Training. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):346–347, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barr:1999:ENP

[BHPG99] Matthew Barr, Sam Holden, Dave Phillips, and Tony Greening. An exploration of novice programming errors in an objectoriented environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):42–46, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biermann:1990:OCA

[Bie90] Alan W. Biermann. An overview course in academic computer science: a new approach for teaching nonmajors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):236–239, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biernat:1993:TTD

[Bie93] Martin J. Biernat. Teaching tools for data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):9–12, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Borstler:1998:SCT

[BJ98] Jürgen Börstler and Olof Johansson. The students conference — a tool for the teaching of research, writing, and presentation skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):28–31, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boulet:1998:TPM

[BJB98] Marie-Michèle Boulet, Faouzi Ben Jebara, and Serge Boudreault. Teaching process modelling to undergraduate computer students: effects of using Lotus Notes and television programmes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):32–36, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blythe:1994:LLV

[BJR94] Stephen A. Blythe, Michael C. James, and Susan H. Rodger. LLparse and LRparse: visual and interactive tools for parsing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):208–212, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1992:USE

[BK92a] Doug Baldwin and Johannes A. G. M. Koomen. Using scientific experiments in early computer science laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):102–106, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bickerstaff:1992:ISW

[BK92b] Douglas D. Bickerstaff and Judith D. Kaufman. Improving student writing skills: inter-departmental collaborations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):42–45, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barr:1995:EIB

[BK95] John Barr and L. A. Smith King. An environment for interpreter-based programming language projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):159–162, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnett:1998:DII

[BKCG98] Lewis Barnett, Joseph F. Kent, Justin Casp, and David Green. Design and implementation of an interactive tutorial framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):87–91, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruce:1998:PLC

[BKLV98] Kim B. Bruce, Phokion G. Kolaitis, Daniel M. Leivant, and Moshe Y. Vardi. Panel: logic in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):376–377, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1999:RNG

[BKP^{+}99] Joseph Bergin, Amruth Kumar, Viera K. Proulx, Myles Mc-Nally, Alyce Faulstich Brady, David Mutchler, Stephen Hartley, Richard Rasala, Charles Kelemen, Rocky Ross, and Frank Klassner. Resources for next generation introductory CS courses: Report of the ITiCSE’99 Working Group on Resources for the Next Generation CS 1 Course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4): 101–105, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bock:1999:ICS

[BKS^{+}99] Douglas B. Bock, Robert W. Klepper, John F. Schrage, G. Gregory Stephen, and Bernard Waxman. Integrating computer science and information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):56–60, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beil:1998:IFB

[BLOS98] Donald H. Beil, Donna Lange, Elissa Olsen, and Anthony Spiecker. Incoming freshmen build their own PCs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):181–184, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blum:1996:ICW

[Blu96] Howard Blum. Internet connection for Web access: an example for performance modeling and simulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):62–64, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baram:1996:UCC

[BM96] Giora Baram and Munir Mandviwalla. Use of computer conferencing in teaching systems analysis and design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):37–39, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1995:VBI

[BMB^{+}95] Doug Baldwin, Jerry Mead, Keith Barker, Allen Tucker, and Lynn R. Ziegler. Visions of breadth in introductory computing curricula (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):386–387, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bagert:1995:SFY

[BMC95] Donald Bagert, William M. Marcy, and Ben A. Calloni. A successful five-year experiment with a breadth-first introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):116–120, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Batlle:1996:BEP

[BMP96] Joan Batlle, Joan Martì, and Lluìs Pacheco. A built-in educational platform for integrating computer engineering technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):229, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernstein:1995:ESY

[BMR^{+}95] Danielle R. Bernstein, Bill Marion, Catherine M. Ricardo, Joy Teague, and Lee Wittenberg. Every seventh year: sabbatical opportunities (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):376–377, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1998:GWG

[BN98] Joseph Bergin and Thomas L. Naps. Group 3 (working group): curricular opportunities of Java based Internet software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):270, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:1998:JRC

[BNB^{+}98] Joseph Bergin, Thomas L. Naps, Constance G. Bland, Stephen J. Hartley, Mark A. Holliday, Pamela B. Lawhead, John Lewis, Myles F. McNally, Christopher H. Nevison, Cheng Ng, George J. Pothering, and Tommi Teräsvirta. Java resources for computer science instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4): 18–38, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berghel:1997:YLA

[BNS^{+}97] Hal Berghel, Cathleen Norris, Elliot Soloway, Harriet Taylor, and Joe Turner. You learned all you need to design educational software design in kindergarten (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):139, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bagert:1995:OTA

[BO95] Donald Bagert and Barbara Boucher Owens. Organizing a team for the ACM programming contest (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):402, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bobkowska:1999:LEM

[Bob99] Anna Bobkowska. Laboratory experiment with the 3RolesPlaying method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):123–126, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bonner:1999:SDC

[Bon99] Adam Bonner. A student-directed computing community. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):155–159, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boufaida:1996:WCD

[Bou96] Mahmoud Boufaida. What courseware dedicated to computer science? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):8–14, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boulet:1997:DLM

[Bou97] Marie-Michèle Boulet. Distance learning of the management of software projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):136–138, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Buckalew:1994:LA

[BP94] Chris Buckalew and Alan Porter. The lecturer’s assistant. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):193–197, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Budd:1995:NMP

[BP95] Timothy A. Budd and Rajeev K. Pandey. Never mind the paradigm, what about multiparadigm languages? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):25–30, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyle:1999:UIC

[BR99] Roger Boyle and Ann Roberts. The University in the inner city. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):119–122, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bradley:1996:AMP

[Bra96] Michael J. Bradley. Analyzing multi-phase searching algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):5–8, September 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brewer:1992:PA

[Bre92] R. K. Brewer. Down with polynomial addition! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):225–229, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brette:1995:TRC

[Bre95] Jean-François Brette. Transparent running and contextual help to learn and to teach an imperative language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):7–12, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Briggs:1991:GPS

[Bri91] Jim Briggs. Group projects in software engineering at york. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):48–50, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bridges:1993:GAD

[Bri93] Susan Bridges. Graphics assignments in discrete mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):83–86, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1992:CLG

[Bro92] David T. Brown. CS 1 labs: goals and expectations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):29–34, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1993:DMI

[Bro93a] David T. Brown. Discrete mathematics II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):13–17, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1993:STA

[Bro93b] Richard A. Brown. A software testbed for advanced projects in real-time and distributed computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (1):247–250, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1994:HSP

[Bro94] David W. Brown. High school participation in the association for computing machinery (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 410, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brogi:1997:TMC

[Bro97] Antonio Brogi. A Turing machine contest for introducing high school students to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 23–27, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruce:1994:AKB

[Bru94] Kim B. Bruce. Attracting (& keeping) the best and the brightest: an entry-level course for experienced introductory students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):243–247, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruce:1999:FSI

[Bru99] Kim B. Bruce. Formal semantics and interpreters in a principles of programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):331–335, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bryant:1999:SES

[Bry99] Robert Bryant. Software engineering for seniors-overcoming the administrative fears. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):83–86, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boerner:1995:CSP

[BS95] Gerald L. Boerner and Carol Backer Stoker. Creating selfpaced courses for CS majors and non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):394–395, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blum:1999:ADL

[BS99] Howard Blum and David Sachs. An asynchronous distancelearning course in data communications and networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):52–55, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:1994:TFI

[BSK94] Doug Baldwin, Greg Scragg, and Hans Koomen. A three-fold introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):290–294, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:1994:TTH

[BSM^{+}94] David W. Brown, Michael A. Sheets, Randy L. Myers, Jeremy A. Freed, Allan Cameron, Patricia Amavisca, Theresa Cuprak, Brian Pollack, and Chris Stephenson. A tale of two high school computer science programs and how the ACM model high school computer science curriculum may shape their future (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):394–395, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berque:1995:NSR

[BT95] Dave Berque and Gloria Townsend. A new scheme for reinforcing concepts in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):327–330, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biddle:1996:EIC

[BT96] Robert Biddle and Ewan Tempero. Explaining inheritance: a code reusability perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):217– 221, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biddle:1998:JPB

[BT98] Robert Biddle and Ewan Tempero. Java pitfalls for beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):48–52, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burkhart:1997:PPU

[Bur97] Helmar Burkhart. Parallel programming using public domain software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):224–228, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burke:1997:GAT

[BV97] E. K. Burke and D. B. Varley. A genetic algorithms tutorial tool for numerical function optimisation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (3):27–30, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brilliant:1996:FPP

[BW96] Susan S. Brilliant and Timothy R. Wiseman. The first programming paradigm and language dilemma. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):338–342, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beheshti:1997:ISC

[BW97] Mohsen Beheshti and Bill Waller. Instructional software for closed laboratories in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):40–41, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brodlie:1996:SVS

[BWW96] Ken Brodlie, Jason Wood, and Helen Wright. Scientific visualization-some novel approaches to learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):28–32, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yue:1990:TGE

[bY90] Kwok bun Yue. Teaching a graduate expert systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):49–52, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yue:1991:DPR

[bY91] Kwok bun Yue. Dining philosophers revisited, again. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):60–64, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yue:1994:UCC

[bY94] Kwok bun Yue. An undergraduate course in concurrent programming using ada. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):59–63, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barr:1998:NMS

[BYC98] Bob Barr, Sung Yoo, and Tom Cheatham. Network monitoring system design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):102–106, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chappelow:1990:DIS

[CAH90] Stephen W. Chappelow, Steven F. Ackerman, and Stephen J. Hartley. Design and implementation of a swapper for the MINIX operating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):55–59, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:1994:GIS

[Cam94] Harold G. Campbell. Geographic information systems education for non-computer oriented college students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):11–14, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:1996:ICC

[Cam96] Robert A. Campbell. Introducing computer concepts by simulating a simple computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):9–11, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Canning:1991:HDA

[Can91] James T. Canning. A hands-on dataflow architecture/ programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):29–32, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carlson:1990:EGC

[Car90] Wayne E. Carlson. An environment of a graduate curriculum in computer graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):15–20, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carrasquel:1993:NML

[Car93] Jacobo Carrasquel. Necessity is the mother of language features. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):59–64, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cartereau:1994:TOS

[Car94] Michel Cartereau. A tool for operating system teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):51–55, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carissimo:1995:XEU

[Car95] Joel W. Carissimo. XINU-an easy to use prototype for an operating system course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):54–56, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carswell:1997:TII

[Car97] Linda Carswell. Teaching via the Internet: the impact of the Internet as a communication medium on distance learning introductory computing students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):1– 5, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carrington:1998:PAS

[Car98a] David A. Carrington, editor. Proceedings of the ACM SIGCSE 3rd Australasian Conference on Computer Science Education, ACSE 1998, The University of Queensland, Brisbane, Queensland, Australia, July 8–10, 1998, volume 3 of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1998. CO-DEN SIGSD3. ISBN 1-58113-018-X. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carswell:1998:VUT

[Car98b] Linda Carswell. The “Virtual University”: toward an Internet paradigm? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):46–50, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carlisle:1999:GF

[Car99a] Martin C. Carlisle. Graphics for free. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2): 65–68, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carrasquel:1999:TCL

[Car99b] Jacobo Carrasquel. Teaching CS1 on-line: the good, the bad, and the ugly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):212–216, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:1999:CPW

[Car99c] Janet Carter. Collaboration or plagiarism: what happens when students work together. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):52–55, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Castaneda:1996:WBA

[Cas96] Sheila E. Castaneda. Windows-based applications in Visual Basic using project teams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):232, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Casey:1998:LTW

[Cas98a] Des Casey. Learning “from” or “through” the Web: models of Web based education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):51–54, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:1998:DLU

[Cas98b] Lillian N. Cassel. A digital library for undergraduate education: goals and status of efforts at the US national science foundation (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):272, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:1998:FON

[Cas98c] Lillian N. Cassel. Funding opportunities at the National Science Foundation Division of Undergraduate Education (NSF DUE) (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):381, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

CastroG:1999:MNT

[Cas99] Fernando Castro G. More on the nine-tails problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):47, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Calloni:1994:IPB

[CB94a] Ben A. Calloni and Donald J. Bagert. ICONIC programming in BACCII vs. textual programming: which is a better learning environment? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):188–192, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cook:1994:TIA

[CB94b] Dave Cook and Eugene Bingue. Tutorial introduction to Ada 9X (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):373, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carswell:1996:AGA

[CB96] Linda Carswell and David Benyon. An adventure game approach to multimedia distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):122–124, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Calloni:1997:IPP

[CBH97] Ben A. Calloni, Donald J. Bagert, and H. Paul Haiduk. Iconic programming proves effective for teaching the first year programming sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):262–266, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Connelly:1996:HSS

[CBPW96] Christopher Connelly, Alan W. Biermann, David Pennock, and Peter Wu. Home-study software: flexible, interactive, and distributed software for independent study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):63–67, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cote:1992:IPT

[CC92] Vianney Côté and Guy Custeau. An integrating pedagogical tool based on writing articles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):38–41, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Concepcion:1999:AAA

[CCMD99] Arturo I. Concepcion, Lawrence E. Cummins, Ernest J. Moran, and Man M. Do. Algorithma 98: an algorithm animation project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):301–305, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carrasquel:1997:MLI

[CCR^{+}97] Jacobo Carrasquel, Michael J. Clancy, Eric Roberts, Joseph L. Zachary, and David G. Kay. Managing large introductory courses (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):386–387, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cabeza:1999:CCS

[CCR99] M. Luisa Córdoba Cabeza, M. Isabel García Clemente, and M. Luz Rubio. CacheSim: a cache simulator for teaching memory hierarchy behaviour. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):181, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chaudhuri:1993:NSQ

[CD93] R. Chaudhuri and A. C. Dempster. A note on slowing Quicksort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):57–58, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:1995:UII

[CD95] Norman Cohen and Wanda Dann. Using an internal internship to enhance computer science education in a two-year college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):44–47, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clancy:1997:GSC

[CDD^{+}97] Michael J. Clancy, Ed Dubinsky, Richard A. Duggan, Marian Petre, Vicki L. Almstrum, and Mickey McDonald. Getting started in CS education research (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):375–376, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Couger:1997:PRP

[CDF^{+}97] J. Daniel Couger, Gordon B. Davis, David L. Feinstein, John T. Gorgone, and Herbert E. Longenecker, Jr. I.S. ’97 (panel): a report and panel discussion of the joint ACM/AIS/DPMA information systems curriculum for four year undergraduate programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):381–382, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caristi:1994:ACS

[CDMT94] James Caristi, Nell Dale, Bill Marion, and Joe Turner. Assessment in computer science (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):401–402, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coninx:1997:DRI

[CDVF97] K. Coninx, B. Daems, F. Van Reeth, and E. Flerackers. Design and realization of an interactive multimedia server in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):65–70, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chang:1999:CBF

[CEK^{+}99] Carl Chang, Gerald Engel, Willis King, Eric Roberts, Russ Shackelford, Robert H. Sloan, and Pradip K. Srimani. Curriculum 2001: bringing the future to the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):70–73, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cupper:1997:URW

[CEMP97] Robert D. Cupper, Rhonda Eller-Meshreki, and Gerald Pitts. Undergraduate research-welcome to the 21st century. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):367–368, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:1998:OOS

[CF98] J. Carter and A. Fowler. Object oriented students? (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):271, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chapman:1994:CSA

[CFA^{+}94] Gail Chapman, Sarah Fix, Owen Astrachan, Joseph Kmoch, and Michael Clancy. Case studies in the advanced placement computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):371, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:1992:GHS

[CFK^{+}92] M. Cohen, M. Foster, D. Kratzer, P. Malone, and A. Solem. Get high school students hooked on science with a challenge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):240–245, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:1993:STD

[CGJ93] Margaret Christensen, Michael Giamo, and Trevor Jones. Support for teaching the design and implementation of multimedia/hypermedia systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):242–245, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carlson:1996:WIL

[CGK^{+}96] David Carlson, Mark Guzdial, Colleen Kehoe, Viren Shah, and John Stasko. WWW interactive learning environments for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):290–294, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cagnat:1990:ALE

[CGP90] J. M. Cagnat, V. Gueraud, and J. P. Peyrin. The arcade laboratory: an environment to help teach algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):37–42, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chitsaz:1991:ABI

[CH91] Mohsen Chitsaz and Karen Holbrook. The achievement of blacks in introductory computer science at a predominantly white public university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):2–8, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carmony:1993:EAI

[CH93a] Lowell A. Carmony and Robert L. Holliday. An example from artificial intelligence for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 1–5, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Curl:1993:ICN

[CH93b] Linda A. Curl and Brent J. Hussin. Introductory computing: a new approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):131–135, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Callaghan:1998:DVD

[CH98] Michael Callaghan and Heiko Hirschmüller. 3-D visualisation of design patterns and Java programs in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):37–40, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chan:1991:CAP

[Cha91a] Tat-Hung Chan. Computing average path lengths of binary search trees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):10, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chavey:1991:SLC

[Cha91b] Darrah Chavey. A structured laboratory component for the introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):287–295, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chandra:1993:SPP

[Cha93] Kailash Chandra. A set of programming projects for a second programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):51–56, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chavey:1996:SAA

[Cha96] Darrah Chavey. Songs and the analysis of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):4–8, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chiu:1999:IPN

[CHCL99] Chiung-Hui Chiu, Chun-Chieh Huang, Wen-Tsung Chang, and Tsung-Ho Liang. Interaction processes in network supported collaborative concept mapping. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 75–78, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chesnevar:1994:SPA

[Che94a] Carlos Iván Chesñevar. Some problems about English–Spanish translations in computer science literature. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):15–ff., September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chesnevar:1994:SDT

[Che94b] Carlos Iván Chesñevar. Syntactic diagrams as a tool for solving text-processing problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):35–40, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carver:1996:MAS

[CHL96] Curtis A. Carver, Richard A. Howard, and William D. Lane. A methodology for active, student-controlled learning: motivating our weakest students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):195– 199, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Classen:1995:DCU

[CHLS95] Ron Classen, Larry Hughes, Elizabeth Leboffe, and Wayne Smith. Data communications in the undergraduate computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):390– 391, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chmura:1998:WAN

[Chm98] Gail A. Chmura. What abilities are necessary for success in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):55–58, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chung:1999:RAC

[Chu99] Sei-Jong Chung. Recursive algorithm with C++ program for floating-point arithmetic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):28–30, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cigas:1992:AS

[Cig92] John F. Cigas. The art of the state. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (1):153–156, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cizmar:1997:CSR

[Ciz97] Dawn Cizmar. CS student research experience applied to developing instructional technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 120–126, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chu:1991:FPT

[CJ91] I-Ping Chu and Richard Johnsonbaugh. The four-peg Tower of Hanoi puzzle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):2–4, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:1999:GPW

[CJ99] Janet Carter and Tony Jenkins. Gender and programming: what’s going on? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):1–4, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carbone:1998:SMU

[CK98] Angela Carbone and Jens J. Kaasbøll. A survey of methods used to evaluate computer science teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):41–45, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:1999:ULS

[CK99] A. T. Chamillard and Dolores Karolick. Using learning style data in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):291–295, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collofello:1994:ASP

[CKK94] James S. Collofello, Manmahesh Kantipudi, and Mark A. Kanko. Assessing the software process maturity of software engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):16–20, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cordes:1997:APL

[CKM^{+}97] David Cordes, Bruce J. Klein, Renée A. McCauley, Linda Null, Craig E. Wills, and Dorothy Deremer. Application of peer learning to the introductory computer science curriculum (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):373–374, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clancy:1990:FF

[CL90] Michael J. Clancy and Marcia C. Linn. Functional fun. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):63–67, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clancy:1992:CSC

[CL92] M. J. Clancy and M. C. Linn. Case studies in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):220–224, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunha:1998:ICP

[CL98] José C. Cunha and João Lourenço. An integrated course on parallel and distributed processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):217–221, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clancy:1999:PP

[CL99] Michael J. Clancy and Marcia C. Linn. Patterns and pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):37–42, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clarke:1993:PMD

[Cla93] Matthew C. Clarke. Possible models diagrams: a visual alternative to truth tables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):232–236, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clark:1998:EDI

[Cla98a] Martyn Clark. Electronic discussions in information systems (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):273, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clarke:1998:TEA

[Cla98b] Matthew C. Clarke. Teaching the empirical approach to designing human-computer interaction via an experiential group project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):198–201, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1997:CCA

[Cle97a] Tony Clear. Coupling and cohesion among disciplines: some curriculum paradigms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):14–16, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1997:NCA

[Cle97b] Tony Clear. The nature of cognition and action. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):25–29, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1998:GCD

[Cle98a] A. G. Clear. A generic collaborative database-part of a strategy to internationalise the curriculum and develop teamwork and communication capabilities (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):274, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1998:SCB

[Cle98b] Tony Clear. From structure to context-bridging the gap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):13–14, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1998:PCN

[Cle98c] Tony Clear. Programming in context — the next step. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):13–14, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1999:CLT

[Cle99a] A. G. (Tony) Clear. A collaborative learning trial between New Zealand and Sweden — using Lotus Notes Domino in teaching the concepts of Human Computer Interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):111–114, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1999:EOS

[Cle99b] Tony Clear. “Eating Our Seed Corn” or restricting crop diversity? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):15–16, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:1999:OEI

[Cle99c] Tony Clear. Online education — but is it education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):17–18, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clifton:1991:IAS

[Cli91] Joseph M. Clifton. An industry approach to the software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):296–299, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cao:1996:ICC

[CLL^{+}96] Nguyen Vi Cao, Atika Laribi, Michel Léonard, Olivier Parchet, and Catherine Zellweger. Integrating CSCW in a cooperative learning environment to teach information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):125–129, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Castro:1999:PC

[CLV99] Fernando G. Castro, Reynaldo Lezama, and Albeniz Vasquez. Playing click. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):35–36, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Concepcion:1996:DVL

[CM96] Arturo I. Concepcion and David R. Millican. Developing the VLSI laboratory for the computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):47–52, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Camp:1995:MST

[CO95] Tracy Camp and Glen Oberhauser. Microkernels: a submodule for a traditional operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):154–158, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coey:1993:ITS

[Coe93] W. A. Coey. An interactive tutorial system for MC68000 assembly language using HyperCard. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2): 19–23, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohoon:1999:DDC

[Coh99] J. McGrath Cohoon. Departmental differences can point the way to improving female retention in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):198–202, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cole:1991:WLA

[Col91a] Jerald D. Cole. WHILE loops and the analogy of the single stroke engine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):20–22, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collins:1991:EET

[Col91b] William J. Collins. Estimating execution times: a laboratory exercise for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):358–363, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Concepcion:1998:UOO

[Con98a] Arturo I. Concepcion. Using an object-oriented software lifecycle model in the software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):30–34, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Condon:1998:ISI

[Con98b] Jennifer Condon. The Irish software industry and education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):1–4, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cook:1996:CSF

[Coo96a] Curtis R. Cook. A computer science freshman orientation course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):49–55, June 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coon:1996:SLE

[Coo96b] Lawrence A. Coon. A sequence of lab exercises for an introductory compiler construction course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (3):60–64, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cook:1997:CCS

[Coo97] Curtis R. Cook. CS0: computer science orientation course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):87–91, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cooper:1998:DII

[Coo98] Peter A. Cooper. Designing Internet instructional environments (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):275, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cordova:1999:MPG

[Cor99] Jose L. Cordova. Mathematical proofs as graph search problems in theory courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):110–113, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cull:1994:IQP

[CP94] Paul Cull and Rajeev Pandey. Isomorphism and the N-queens problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):29–36, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cross:1996:SIT

[CP96] James H. Cross II and Thomas M. Phillips. Successfully integrating traditional and object-oriented approaches with Ada 95. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):19–23, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clancy:1993:APA

[CPS93] Michael J. Clancy, Richard E. Pattis, and Mark Stehlik. Approaches to programming assignments in CS 1 and CS 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):308, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Curzon:1998:WDS

[CR98] Paul Curzon and Janet Rix. Why do students take programming modules? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):59–63, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Canup:1998:USS

[CS98] Mark J. Canup and Russell L. Shackelford. Using software to solve problems in large computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):135–139, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collins:1999:WLS

[CS99] William J. Collins and Yi Sun. Web labs for the Standard Template library and the Java generic library in a CS2 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):174–178, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cowley:1993:GLI

[CSB93] Bruce Cowley, Greg Scragg, and Doug Baldwin. Gateway laboratories: integrated, interactive learning modules. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):180–184, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clarke:1994:PPG

[CT94] Valerie A. Clarke and G. Joy Teague. A psychological perspective on gender differences in computing participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):258–262, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Culwin:1997:JCC

[Cul97] Fintan Culwin. Java in the C.S. curriculum (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):392, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Culwin:1998:WHA

[Cul98] Fintan Culwin. Web hosted assessment-possibilities and policy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):55–58, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Culwin:1999:OI

[Cul99] Fintan Culwin. Object imperatives! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 31–36, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1991:UIP

[Cun91] Steve Cunningham. User interface programming: a humancomputer communication course for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):330–336, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1994:UCC

[Cun94] Sally Jo Cunningham. Using a computer conferencing system to support writing and research skill development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):5–8, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1995:IRC

[Cun95] Sally Jo Cunningham. An introduction to research and the CS/ IS professional literature for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):5–8, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1996:IAI

[Cun96] Sally Jo Cunningham. Internet-accessible information retrieval tools for advanced DB/IR courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):107–111, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1997:TSC

[Cun97] Sally Jo Cunningham. Teaching students to critically evaluate the quality of Internet research resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2):31–ff., June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:1998:OBC

[Cun98] Steve Cunningham. Outside the box: the changing shape of the computing world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):4–7, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cupper:1998:CSP

[Cup98] Robert D. Cupper. Computer science: a proposed alternative track-applied computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):25–29, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Curl:1992:WAP

[Cur92] L. A. Curl. Writing about programming in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):7–10, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Curtis:1997:WBC

[Cur97] Ronald Curtis. A Web based configuration control system for team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):189–193, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coe:1996:IET

[CWI96] P. S. Coe, L. M. Williams, and R. N. Ibbett. An interactive environment for the teaching of computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):33–35, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cole:1998:UJD

[CWS98] David Cole, Roger Wainwright, and Dale Schoenefeld. Using Java to develop Web based tutorials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 92–96, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cynar:1990:TMS

[Cyn90] Sandra J. Cynar. Teaching modeling and simulation to computing science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):34–37, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Diz:1997:HCS

[DA97] Paloma Díz and Ignacio Aedo. A hypermedia course in Spanish informatics engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):58–61, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dadamo:1990:CQA

[Dad90] Diana T. Dadamo. The correlation quiz: an aid in curbing cheating in programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2): 52–54, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1990:IYW

[Dal90] Nell Dale. If you were lost on a desert island, what one ADT would you like to have with you? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 139–142, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dalbey:1991:AZP

[Dal91a] John Dalbey. Applying zen principles in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):21–23, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1991:PST

[Dal91b] Nell B. Dale, editor. Proceedings of the 22nd SIGCSE Technical Symposium on Computer Science Education, 1991, San Antonio, Texas, USA, March 7–8, 1991, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1991. CODEN SIGSD3. ISBN 0-89791-377-9. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1992:RSM

[Dal92a] Nell Dale. Results of SIGCSE membership survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):1–6, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1992:PST

[Dal92b] Nell B. Dale, editor. Proceedings of the 23rd SIGCSE Technical Symposium on Computer Science Education, 1992, Kansas City, Missouri, USA, March 5–6, 1992, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1992. CO-DEN SIGSD3. ISBN 0-89791-468-6. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1998:PRR

[Dal98a] Nell Dale. Publishing research results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):

17–18, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1998:TTE

[Dal98b] Nell Dale. Two threads from the empirical studies of programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):16–17, December 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daly:1998:PSC

[Dal98c] Charlie Daly. A proposed structure for a computer based learning environment — a pragmatic approach (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):276, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1999:BRE

[Dal99a] Nell Dale. Bibliography relevant to education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):18–19, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1999:RPR

[Dal99b] Nell Dale. Reflections on past research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):20–22, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:1999:DTU

[Dal99c] Nell B. Dale. Distance teaching at Uppsala. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):39–41, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daly:1999:RIC

[Dal99d] Charlie Daly. RoboProf and an introductory computer programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):155–158, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Danilowicz:1992:AC

[Dan92] Ronald L. Danilowicz. Alice in Computerland. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):20–26, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeLoatch:1994:SCS

[DAUZ94] Sandra J. DeLoatch, Ernest C. Ackermann, John Urquhart, and Lynn Ziegler. Scaling up computer science with efficient learning (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):390–391, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1990:TCP

[Dav90] G. L. Davies. Teaching concurrent programming with Pascal-FC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):38–41, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1998:DEO

[Dav98] Gordon Davies. Distance education and the Open University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):27, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1999:GD

[Dav99] Gordon Davies. Global DL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):27, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dershem:1998:TWB

[DB98] Herbert L. Dershem and Peter Brummund. Tools for Webbased sorting animation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):222–226, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DePano:1990:APE

[DBKL90] N. Adlai A. DePano, Farinaz D. Boudreaux, Philip Katner, and Brian Li. Algorthmic paradigms: examples in computational geometry II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):186–191, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

deCampos:1999:MRE

[dC99] Geraldo Lino de Campos. Minimum requirements for effective distance teaching systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):182, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dillon:1997:TEA

[DDG97] Eric Dillon, Carlos Gamboa Dos Santos, and Jacques Guyard. Teaching an engineering approach for network computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):229–232, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daigle:1996:ICP

[DDP96] R. J. Daigle, Michael V. Doran, and J. Harold Pardue. Integrating collaborative problem solving throughout the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):237–241, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeClue:1996:OOP

[DeC96] Tim DeClue. Object-orientation and the principles of learning theory: a new look at problems and benefits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):232–236, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denman:1996:DRA

[Den96] Richard T. Denman. Derivation of recursive algorithms for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):9–13, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:1999:CP

[Den99] Peter J. Denning. Computing the profession. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):1–2, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dietrich:1997:WWB

[DEP97] Suzanne W. Dietrich, Eric Eckert, and Kevin Piscator. Win-RDBI: a Windows-based relational database educational tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):126–130, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deremer:1993:ILE

[Der93] Dorothy Deremer. Improving the learning environment in CS I: experiences with communication strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):31–35, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Detmer:1990:TAL

[Det90] Richard C. Detmer. Teaching assembly language: a comparison of IBM S/360 and Intel 80x86 courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 116–120, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Detmer:1993:PGP

[Det93] Richard C. Detmer. PC graphics packages for modula-2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):52–57, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deveaux:1999:SET

[DFF99] Daniel Deveaux, Regis Fleurquin, and Patrice Frison. Software engineering teaching: a “docware” approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):163–166, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dickey:1998:ADA

[DFH^{+}98] Martin Dickey, Frank Friedman, Max Hailperin, Bill Manaris, and Ursula Wolz. Accommodating diversity of academic preparation in CS1 (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):361–362, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deremer:1995:ATS

[DFZ^{+}95] Dorothy Deremer, Gail Finley, Lynn Ziegler, Doug Baldwin, and Gordon Stegink. Alternative teaching strategies in CS I: supporting diversity (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):368–369, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1999:THE

[DG99] Gordon Davies and Brian J. Garner. Telematics helpdesk for external students of computing at Deakin University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):29–31, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daniels:1996:TCS

[DGEST96] Mats Daniels, Judith Gal-Ezer, Ian Sanders, and G. Joy Teague. Teaching computer science: experience from four continents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):102–106, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:1990:SCC

[DH90] Richard W. Decker and Stuart H. Hirshfield. A survey course in computer science using HyperCard. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 229–235, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:1993:TTO

[DH93] Rick Decker and Stuart Hirshfield. Top-down teaching: objectoriented programming in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):270– 273, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:1994:TRW

[DH94] Rick Decker and Stuart Hirshfield. The top 10 reasons why object-oriented programming can’t be taught in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):51–55, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dawson-Howe:1995:ASA

[DH95a] Kenneth M. Dawson-Howe. Automatic submission and administration of programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4): 51–53, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dean:1995:IFM

[DH95b] Neville Dean and Michael G. Hinchey. Introducing formal methods through Rôle-playing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):302–306, March 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1995.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
SE/Z.bib.

Dawson-Howe:1996:ASA

[DH96a] Kenneth M. Dawson-Howe. Automatic submission and administration of programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2): 40–42, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dawson-Howe:1996:OFY

[DH96b] Kenneth M. Dawson-Howe. Organization of final year projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):2–4, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorf:1996:NWA

[DH96c] Mary Lou Dorf and Gerald R. Heuring. Nightmare on Westwood Avenue: product development laboratory: success, failure, both, neither??? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):214–216, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dankell:1997:UWS

[DH97] Douglas D. Dankell II and James Hearn. The use of the WWW to support distance learning through NTU. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):8–10, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1998:HCC

[DI98] Gordon Davies and John Impagliazzo. History in the computing curriculum (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):277, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dinerstein:1991:SND

[Din91] N. T. Dinerstein. Strategies in network database systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):23–26, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dingle:1997:OOM

[Din97] Adair Dingle. The object-ownership model: a case study for inheritance and operator overloading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 82–86, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dixon:1991:IOO

[Dix91] Chip Dixon. An introduction to object-oriented programming through Turbo Pascal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):33–35, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dixon:1993:OET

[Dix93] Chip Dixon. Object extensibility: a Turbo Pascal graphics example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):38–40, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davy:1999:RLI

[DJ99] John Davy and Tony Jenkins. Research-led innovation in teaching and learning programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):5– 8, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Delahan:1995:AWS

[DKT95] Franz Delahan, William F. Klostermeyer, and George Trapp. Another way to solve nine-tails. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4): 27–28, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dunning:1990:MCP

[DL90] Larry A. Dunning and Ronald L. Lancaster. The millionth computer program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):224–228, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doran:1995:CBA

[DL95] Michael V. Doran and David D. Langan. A cognitive-based approach to introductory computer science courses: lesson learned. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):218–222, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

delaFuente:1999:TCA

[dlFCC99] Santiago Rodríguez de la Fuente, M. Isabel García Clemente, and Rafael Méndez Cavanillas. Teaching computer architecture with a new superscalar processor emulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):99–102, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dey:1992:CTC

[DM92] Sukhen Dey and Lawrence R. Mand. Current trends in computer science curriculum: a survey of four-year programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):9–14, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

deMarneffe:1998:PEQ

[dM98] P. A. de Marneffe. The problem of examination questions in algorithmics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):74–76, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dixon:1997:UNS

[DMK97] Michael W. Dixon, Tanya J. McGill, and Johan M. Karlsson. Using a network simulation package to teach the client-server model. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):71–73, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dezhgosha:1997:CND

[DMW^{+}97] Kamyar Dezhgosha, Ted Mims, Richard Wasniowski, Frances Grodzinsky, Lawrence J. Osborne, Dean Sanders, and Herb Dershem. Computer networks and data communications (panel): a laboratory focus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):365– 366, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daigle:1997:ICS

[DN97] Roy J. Daigle and Marino J. Niccolai. Inter-class synergy by design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):92–95, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denman:1994:DPF

[DNPR94] Richard Denman, David A. Naumann, Walter Potter, and Gary Richter. Derivation of programs for freshmen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):116–120, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:1993:CPL

[dOG93] José de Oliveira Guimarães. Clever programming languages exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):61–64, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:1994:TSM

[dOG94] José de Oliveira Guimarães. Testing student-made compilers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):37–44, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Donaldson:1990:OSA

[Don90] John L. Donaldson. Operating systems from assembler to C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):121–124, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Donaldson:1995:MSC

[Don95] John L. Donaldson. A microprogram simulator and compiler for an enhanced version of Tanenbaum’s MIC-1 machine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):238–242, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorf:1992:BRW

[Dor92] M. L. Dorf. Backtracking the rat way. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (1):272–275, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorin:1997:PPE

[Dor97] Philip M. Dorin. Practice + paradigms: experience with a firstyear programming workshop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):42–44, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Downey:1999:TED

[Dow99] Allen B. Downey. Teaching experimental design in an operating systems class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):316–320, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:1990:CSH

[DP90] Gordon Davies and Jenny Preece. Computer science, home computing and distance learning — the largest computer science course in the world? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):143–146, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dee:1999:ACA

[DR99] Hannah Dee and Peter Reffell. ACOM (“computing for all”): an integrated approach to the teaching and learning of information technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):195, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Drake:1998:CSA

[Dra98] Janet M. Drake. Class to survey application programming topics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):161–165, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Drummond:1998:GSE

[Dru98] Sarah Drummond. Groupware for software engineering student group projects (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):278, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doss:1991:NNN

[DS91] David Doss and Bill Swafford. Networking non-network applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):352–357, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagdilelis:1998:DST

[DS98] V. Dagdilelis and M. Satratzemi. DIDAGRAPH: software for teaching graph theory algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):64–68, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagdilelis:1999:DTO

[DS99] V. Dagdilelis and M. Satratzemi. Didactics too, not only technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):183, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dooley:1994:CV

[DSB94] John F. Dooley, Daniel C. St. Clair, and William E. Bond. Computing χ^{2} values. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):218–222, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dybdahl:1998:AFE

[DST98] Arne Dybdahl, Erkki Sutinen, and Jorma Tarhio. On animation features of Excel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):77–80, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davidovic:1998:OLE

[DT98] Alex Davidovic and Elena Trichina. Open learning environment and instruction system (OLEIS). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 69–73, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dietrich:1996:DTP

[DU96] Suzanne W. Dietrich and Susan D. Urban. Database theory in practice: learning from cooperative group projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):112–116, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Duckworth:1994:IPP

[Duc94] R. James Duckworth. Introducing parallel processing concepts using the MASPAR MP-1 computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):353–356, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dingle:1998:EOO

[DUM98] Adair Dingle, David Umphress, and Russ Moul. Explication of object-oriented databases: a software engineering perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):35–39, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dunstan:1991:BMU

[Dun91] Neil Dunstan. Building monitors with UNIX and C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):7–9, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dershem:1998:JCV

[DV98] Herbert L. Dershem and James Vanderhyde. Java class visualization for teaching object-oriented concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):53–57, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

deVivo:1997:PUO

[dVdVI97] Gabriela O. de Vivo, Marco de Vivo, and Germinal Isern. E pluribus unum: OOPL selection. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 17–ff., June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dyck:1993:EPD

[Dyc93] V. Arnie Dyck. Emphasizing the process in delivering CS-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):127–130, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eaton:1990:WAW

[EBD^{+}90] Virginia Eaton, Sharon Bell, Nell Dale, Susie Gallagher, Helen Gigley, and Cindy Hanchey. Where have all the women gone? (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):251, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eckert:1990:CBC

[Eck90] Richard R. Eckert. Communication between computers and peripheral devices — an analogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):54–62, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epstein:1994:MAD

[EFGR94] Steven Epstein, Marla Fischer, Forouzan Golshani, and Catherine Ricardo. Multimedia across the disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):387–388, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elorriaga:1999:PCS

[EGIU99] Jon Ander Elorriaga, Julián Gutiérrez, Jesús Ibáñez, and Imanol Usandizaga. A proposal for a computer security course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):42–47, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:1999:CSD

[EHS^{+}99] Ainslie Ellis, Dianne Hagan, Judy Sheard, Jason Lowder, Wendy Doube, Angela Carbone, John Robinson, and Sylvia Tucker. A collaborative strategy for developing shared Java teaching resources to support first year programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):84–87, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Egan:1998:FVT

[EKR98] M. A. Egan, M. Krishnamoorthy, and K. Rajan. FCLUST: a visualization tool for fuzzy clustering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):227–231, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Etlinger:1994:PLL

[EL94] Henry A. Etlinger and Michael J. Lutz. Professional literacy: labs for advanced programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):102–105, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erickson:1997:PPB

[EL97] Carl Erickson and Paul Leidig. A pedagogical pattern for bringing service into the curriculum via the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):54–56, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elenborgen:1996:PDA

[Ele96] Bruce S. Elenborgen. Parallel and distributed algorithms: laboratory assignments in Joyce/Linda. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1): 14–18, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elenbogen:1999:CNM

[Ele99] Bruce S. Elenbogen. Computer network management: theory and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):119–121, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:1996:DIT

[Ell96] Ainslie E. Ellis. Design, implementation and testing techniques for multimedia industrial experience projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):119–121, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:1998:GWG

[Ell98a] Ainslie Ellis. Group 1 (working group): development and use of multimedia and Internet resources for a problem based learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):269, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:1998:UMW

[Ell98b] Ainslie Ellis. Use of multimedia and World Wide Web resources in a traditional lecture/tutorial environment (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):279, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elsworth:1992:MCW

[Els92] E. F. Elsworth. The MSL compiler writing project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):41–44, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elrad:1990:ACS

[EN90] Tzilla Elrad and Daniel E. Nohl. The analysis and comparison of scheduling controls in concurrent languages through classification. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):89–93, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:1999:BSC

[Eng99a] John English. Building self-contained Websites on CD-ROM. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):159–162, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:1999:BP

[Eng99b] John English. The BURKS project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):184, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Estell:1995:ECP

[EO95] John K. Estell and Thomas A. Owen. Experiencing the codesign process: Microcomputer systems II laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):34–38, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epp:1991:ECS

[Epp91] Ed C. Epp. An experimental computer science laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):11–14, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epp:1992:YAA

[Epp92] Ed C. Epp. Yet another analysis of algorithms laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):11–14, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epstein:1993:POA

[Eps93] Richard G. Epstein. The power of OOPS (abstract): heterogeneous data structures in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 311, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epstein:1995:LDK

[Eps95] Richard Gary Epstein. Latest developments in the “killer robot” computer ethics scenario. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1): 111–115, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erickson:1993:UTT

[Eri93] Carl Erickson. USENET as a teaching tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):43–47, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erickson:1996:ELE

[Eri96] Carl Erickson. The EOS laboratory environment for a course in operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):353–357, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eskin:1999:GPA

[ES99] Eleazar Eskin and Eric Siegel. Genetic programming applied to Othello: introducing students to machine learning research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):242–246, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Etlinger:1990:RES

[Etl90] Henry A. Etlinger. A retrospective on an early software projects course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):72–77, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Evans:1995:MUA

[Eva95] M. D. Evans. A methodology for a unified ADT course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):50–52, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Evans:1996:NEP

[Eva96] M. D. Evans. A new emphasis & pedagogy for a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):12–16, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

East:1997:PBP

[EW97] J. Philip East and Eugene Wallingford. Pattern-based programming in initial instruction (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):393, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Enright:1994:ACA

[EWC94] Aaron Garth Enright, Linda M. Wilkens, and James T. Canning. An alternative computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):9–12, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ewing:1993:MSC

[Ewi93] Donald J. Ewing. Microcomputer systems I: a computer science and engineering capstone course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (1):155–159, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ezell:1990:CPP

[Eze90] C. L. Ezell. Creating pedagogical programming environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):42–46, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ferri:1998:ACA

[FA98] Francesc J. Ferri and Jesús Albert. Average-case analysis in an elementary course on algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):202–206, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1990:ASC

[FACW90] Michael B. Feldman, Mary Armstrong, Richard Conn, and Philip Wilsey. Ada sources for computer science educators (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):263, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fagin:1994:TYD

[Fag94] Barry Fagin. Two years of “the digital world”: portable courseware for technological literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):97–101, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Faltin:1999:DCA

[Fal99] Nils Faltin. Designing courseware on algorithms for active learning with virtual board games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 135–138, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1997:CPC

[FB97] Michael B. Feldman and Bruce D. Bachus. Concurrent programming CAN be introduced into the lower-level undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):77–79, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Faahraeus:1999:TEC

[FBR^{+}99] Eva R. Fåhræus, Noel Bridgeman, Joze Rugelj, Barbara Chamberlain, and Ursula Fuller. Teaching with electronic collaborative learning groups: Report of the ITiCSE’99 Working Group on Creative Teaching of Electronic Collaborative Learning Groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):121–128, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:1994:NPE

[FC94] David Finkel and Surendar Chandra. NetCp — a project environment for an undergraduate computer networks course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):174–177, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Faulkner:1999:IUI

[FC99a] Xris Faulkner and Fintan Culwin. Integration of usability issues within initial software development education. (It’s all about the user, dummy!). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):296–300, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:1999:WCA

[FC99b] David Finkel and Isabel F. Cruz. Webware: a course about the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):107–110, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fahraeus:1999:TEC

[FCB^{+}99] E. R. Fahraeus, B. Chamberlain, V. Baykov, N. Bridgeman, R. Dumbraveanu, U. Fuller, and J. Rugeli. Teaching with electronic collaborative learning groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):121–128, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Techreports/UKent.bib; http://www.cs.kent.
ac.uk/pubs/1999/1013.

Ford:1994:RGH

[FCR^{+}94] Matthew A. Ford, Elise N. Cassidente, J. Suzanne Rothrock, David W. Brown, and Daniel Miller. The role of gender in high school computer mediated communication (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):410, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:1993:RAP

[Fek93] Alan Fekete. Reasoning about programs: integrating verification and analysis of algorithms into the introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):198–202, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:1995:EGS

[Fek95] Alan Fekete. Enhancing generic skills in the computer organization course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):273–277, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:1996:PRI

[Fek96] Alan Fekete. Preparation for research: instruction in interpreting and evaluating research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):93–97, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1990:TDS

[Fel90] Michael B. Feldman. Teaching data structures with ada: an eight-year perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):21–29, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1992:PDP

[Fel92] M. B. Feldman. The portable dining philosophers: a movable feast of concurrency and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):276–280, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1999:IOU

[Fel99] Michael B. Feldman. Inspiring our undergraduate students’ asperations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):4–7, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fowler:1994:ECI

[FF94] W. A. Lawrence Fowler and R. H. Fowler. An environment for CS integrating hypertext, program design and language facilities (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):407, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:1991:TPP

[FG91] Allan L. Fisher and Thomas Gross. Teaching the programming of parallel computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):102–107, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:1992:TEP

[FG92] Allan L. Fisher and Thomas R. Gross. Teaching empirical performance analysis of parallel programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):309–313, March 1992. CODEN SIGSD3. ISBN 0-89791-468-6. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/1992.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Parallel/benchmark.bib. The Pa-

pers of the Twenty-Third SIGCSE Technical Symposium on Computer Science Education.

Fekete:1996:DCL

[FG96] Alan Fekete and Antony Greening. Designing closed laboratories for a computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):295–299, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Foster:1991:MFR

[FH91] Louis A. Foster and Norman L. Hughes. Making files real with a virtual disk. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):199–204, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:1996:CSF

[FH96] Sue Fitzgerald and Mary Lou Hines. The computer science fair: an alternative to the computer programming contest. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):368–372, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freeman:1997:GIC

[FHL^{+}97] Elizabeth Freeman, Susanne Hupfer, Catherine Lang, Ralph Morelli, Domenick J. Pinto, Frances L. Van Scoy, and Sandra Honda Adams. Gender imbalance in computer science programs, etiology and amelioration (panel): views for U.S. campuses and elsewhere. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):384–385, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:1994:TCH

[FHS^{+}94] David Finkel, Chet Hooker, Scott Salvidio, Mark Sullivan, and Christopher Thomas. Teaching C++ to high school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):286–289, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fienup:1996:RCC

[Fie96] Mark Fienup. Rethinking the CS-2 course with an objectoriented focus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):23–25, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Farthing:1998:PMC

[FJM98] Dave W. Farthing, Dave M. Jones, and Duncan McPhee. Permutational multiple-choice questions: an objective and efficient alternative to essay-type examination questions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):81–85, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Falconer:1992:BIS

[FK92] David R. Falconer and Martin Katz. Building an infrastructure to support writing across the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):34–37, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleury:1991:PPR

[Fle91] Ann E. Fleury. Parameter passing: the rules the students construct. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):283–286, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleury:1993:EDM

[Fle93] Ann E. Fleury. Evaluating discrete mathematics exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):73–77, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1990:SAP

[FLP90] Michael B. Feldman, Arthur Vargas Lopes, and Manuel Perez. SMALL-Ada: personal computer courseware for studying concurrent programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):206–211, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:1997:UWC

[FMM97] Allan Fisher, Jane Margolis, and Faye Miller. Undergraduate women in computer science: experience, motivation and culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):106–110, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Foley:1992:MSC

[Fol92] Dave Foley. Microcode simulation in the computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):57–59, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:1991:BBC

[For91a] Gary Ford. A bibliography of broad-coverage software engineering textbooks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):55–59, June 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:1991:SUC

[For91b] Gary Ford. The SEI undergraduate curriculum in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):375–385, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:1994:PUS

[For94] Gary Ford. The progress of undergraduate software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):51–55, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:1995:TEQ

[FP95] Sue Fitzgerald and Jerry Place. Teaching elementary queueing theory with a computer algebra system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):350–354, March 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Math/maple-
extract.bib;   ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/1995.bib.

Fell:1997:EMP

[FP97] Harriet J. Fell and Viera K. Proulx. Exploring Martian planetary images: C++ exercises for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 30–34, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:1998:BAT

[FP98] Sally Fincher and Marian Petre. Beyond anecdote towards real transfer: using other institutions’ experience of project work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):86–89, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fell:1996:WAC

[FPC96] Harriet J. Fell, Viera K. Proulx, and John Casey. Writing across the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):204–209, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fell:1998:SDP

[FPR98] Harriet J. Fell, Viera K. Proulx, and Richard Rasala. Scaling: a design pattern in introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):326–330, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:1996:PAS

[FR96a] Alan Fekete and John Rosenberg, editors. Proceedings of the ACM SIGCSE 1st Australasian Conference on Computer Science Education, ACSE 1996, Sydney, NSW, Australia, July 1996, volume 1 of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1996. CODEN SIGSD3. ISBN 0-89791-845-2. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freund:1996:TAC

[FR96b] Stephen N. Freund and Eric S. Roberts. Thetis: an ANSI C programming environment designed for introductory use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):300–304, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fernandez:1998:UVI

[FR98] Alejandro Fernández and Gustavo Rossi. Using visualisation to improve object-oriented thinking (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):280, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frailey:1998:SHC

[Fra98] Dennis J. Frailey. Specialization is harmful to computer education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):4–6, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freeman:1994:DST

[Fre94] Geoffrey Freeman. Do students test their software? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):56–57, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fritz:1991:HAT

[Fri91] Jane M. Fritz. HyperCard applications for teaching information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):55–61, March 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fife:1996:GAU

[FRK96] Leslie D. Fife, Gopal Racherla, and Steven E. Killian. Graduate assistant usage in Oklahoma graduate computer science programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):33–35, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frank:1990:ACC

[FS90] Thomas S. Frank and James F. Smith. Ada as a CS1–CS2 language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):47–51, June 1990. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:1998:VSR

[FST98] Ursula Fuller, John Slater, and Gill Tardivel. Virtual seminarsreal networked results? (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):281, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:1991:PSA

[Ful91] Roy Fuller. PDP-11 simulator for Apple Macintosh is designed for instructional use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):17–19, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:1992:MSA

[Ful92] Roy Fuller. Microcode simulator for Apple Macintosh. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):49–51, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fulda:1994:NPI

[Ful94] Joseph S. Fulda. A note on physical implementations of logical entities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):7, September 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:1996:CSP

[FW96] David Finkel and Craig E. Wills. Computer supported peer learning in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):55–56, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1996:QED

[FZ96] Todd J. Feldman and Julie D. Zelenski. The quest for excellence in designing CS1/CS2 assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):319–323, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Golub:1991:CNC

[GA91] Evan B. Golub and Moshe Augenstein. Creation of a new case for LUPSort: ALTERNATING. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):108–111, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gamble:1995:IFS

[Gam95] R. F. Gamble. Integrating a formal specification course with a software projects course via an editing tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):312–316, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldenson:1990:RSE

[GBB^{+}90] Dennis R. Goldenson, Michael Brown, Jane Bruemmer, Nathan Hull, Roy Jones, Bruce McClellan, Joseph Kmoch, Phillip Miller, Mark Stehlik, and Laurie Werth. Roundtable on structure editing (panel session): teachers’ experiences using carnegie Mellon’s GENIE programming environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):267, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:1999:COS

[GBC^{+}99] Michael Goldweber, John Barr, Tracy Camp, John Grahm, and Stephen Hartley. A comparison of operating systems courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):348–349, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:1994:NPT

[GBL94] Michael Goldweber, John Barr, and Chuck Leska. A new perspective on teaching computer literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):131–135, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gray:1998:CLE

[GBS98] John Gray, Tom Boyle, and Colin Smith. A constructivist learning environment implemented in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):94–97, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gonzalez:1998:MEQ

[GC98] Ruben Gonzalez and Greg Cranitch. Multimedia educationquo vadis? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):90–93, September 1998.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1994:ISC

[GCD^{+}94] J. T. Gorgone, J. D. Couger, G. Davis, D. Feinstein, G. Kasper, J. Little, and H. E. Longenecker. Information systems curriculum — a status report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):2–4, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gueraud:1994:SEC

[GCDP94] V. Gueraud, J. P. Cagnat, J. P. David, and J. P. Pernin. Software environments for computer aided education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):19–25, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1997:MUC

[GDC^{+}97] John T. Gorgone, Gordon Davis, J. Daniel Couger, David L. Feinstein, and Herbert E. Longenecker, Jr. Model undergraduate curriculum guidelines for I.S. programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):16–17, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:1994:EUA

[GdLC94] Mário André Mayerhofer Guimarães, Carlos José Pereira de Lucena, and Maurício Roma Cavalcanti. Experience using the ASA algorithm teaching system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (4):45–50, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldman:1996:TDC

[GEC96] David A. Goldman, Richard R. Eckert, and Maxine S. Cohen. Three-dimensional computation visualization for computer graphics rendering algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):358–362, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Geitz:1991:AIC

[Gei91] Robert Geitz. Algorithms and images: computer graphics as an introduction to science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):82–86, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Geitz:1994:CCP

[Gei94] Robert Geitz. Concepts in the classroom, programming in the lab. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):164–168, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1994:SEF

[Ger94] Judith L. Gersting. A software engineering “frosting” on a traditional CS-1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):233–237, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1998:YS

[Ger98] Judith L. Gersting. A Year 2000 story. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2): 70, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gilberg:1996:CSS

[GF96] Richard F. Gilberg and Behrouz A. Forouzan. Comparison of student success in Pascal and C-language curriculums. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):252–255, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Godfrey:1999:JBS

[GG99] Michael Godfrey and Dan Grossman. JDuck: building a software engineering tool in Java as a CS2 project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):48–52, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gelfand:1998:TDS

[GGT98] Natasha Gelfand, Michael T. Goodrich, and Roberto Tamassia. Teaching data structure design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):331–335, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ghafarian:1991:EAC

[Gha91] Ahmad Ghafarian. An experimental approach to a course on parallel and distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):246–253, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gladfelter:1993:SAD

[GHK93] Suzanne E. Gladfelter, William C. Harris, and Karl J. Klee. Successful associate degree programs in the computing sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):299, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gibson:1992:PPP

[Gib92] David E. Gibson. A Pythagorean problem as a programming exercise. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):45–47, June 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gingras:1990:DPR

[Gin90] Armando R. Gingras. Dining philosophers revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):21–ff., September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:1995:LIM

[Gin95] David Ginat. Loop invariants and mathematical games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):263–267, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:1996:EAP

[Gin96] David Ginat. Efficiency of algorithms for programming beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):256–260, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1990:MIS

[GJMS90] Donald Gotterbarn, Deborah Johnson, Keith Miller, and Gene Spafford. Methods of integrating the study of ethics into the computer science curriculum (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):268, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gates:1997:WFD

[GK97] Ann Q. Gates and Vladik Kreinovich. Why is a function defined as set of ordered pairs? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):57, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:1998:GWG

[GK98] Scott Grissom and Deborah Knox. Group 5 (working group): the on-line computer science teaching centre. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):270, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:1995:ATC

[GKB^{+}95] Scott Grissom, Bill Kubitz, Jack Bresenham, G. Scott Owen, and Dino Schweitzer. Approaches to teaching computer graphics (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):382–383, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:1998:DDL

[GKC^{+}98] Scott Grissom, Deborah Knox, Elana Copperman, Wanda Dann, Michael Goldweber, Janet Hartman, Marja Kuittinen, David Mutchler, and Nick Parlante. Developing a digital library of computer science teaching resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):5–17, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goh:1993:LOE

[GL93] Angela Goh and Peng-Chor Leong. Laboratories and other educational experiences based on curricula ’91. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):175–179, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garland:1994:ISC

[GL94] Will Garland and Virginia Levsen. Information systems curricula in AACSB accredited business schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):26–30, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Granger:1996:ICT

[GL96] Mary J. Granger and Joyce Currie Little. Integrating CASE tools into the CS/CIS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):130–132, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Granger:1997:UII

[GL97] Mary J. Granger and Joyce Currie Little. Using IT to integrate societal and ethical issues in the CS/IS curriculum (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):144, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gee:1991:SEC

[GM91] Richard Gee and Rob McArthur. Some experiences with CAI and NATAL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):61–64, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grygiel:1996:PNA

[GMAB96] Terry Grygiel, Aaron Myers, Greg Arnold, and William Breen. Protocols and network architecture: a first course in data communications and computer networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):329–332, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gifford:1999:VWB

[GMF99] Atticus Gifford, Benjamin J. Menasha, and David Finkel. The visible Web browser. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):209, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:1995:UVD

[GNR^{+}95] Scott Grissom, Tom Naps, Rocky Ross, Dalton Hunkins, Susan Rodger, and Dino Schweitzer. Using visual demonstrations to teach computer science (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):370–371, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goelman:1996:ITT

[Goe96] Don Goelman. The INGRES tutorial as a tool in teaching database theory (or, I can’t believe it’s not training). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):117–119, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goh:1992:OSP

[Goh92] Angela Goh. An operating systems project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):29–34, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldberg:1996:CEC

[Gol96a] Murray W. Goldberg. CALOS: an experiment with computeraided learning for operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):175–179, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:1996:PLC

[Gol96b] Michael Goldweber. Proposal for an on-line computer science courseware review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):230, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldberg:1997:CFR

[Gol97a] Murray W. Goldberg. CALOS: first results from an experiment in computer-aided learning for operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):48–52, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldberg:1997:WFY

[Gol97b] Murray W. Goldberg. WebCT and first year: student reaction to and use of a Web-based resource in first year computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):127–129, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldberg:1998:BSV

[Gol98] Adele Goldberg. Building a system in virtual reality with LearningWorks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):5–9, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:1999:RUH

[Gol99] Michael Goldweber. A report on the use of HyperTalk in CS1 within a liberal arts setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):37–41, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1998:FLG

[Gor98a] John T. Gorgone. A fresh look at IS graduate programs is needed. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):15–16, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1998:ISC

[Gor98b] John T. Gorgone. Information systems curricula and accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):15–16, December 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1999:AMN

[Gor99a] John T. Gorgone. Approaching MSIS 2000: a new-fashioned graduate model. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):19–20, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1999:GIS

[Gor99b] John T. Gorgone. Graduate information systems curriculum for the 21st century. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):352–353, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:1999:ISA

[Gor99c] John T. Gorgone. Information systems accreditation: revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):17–18, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1997:OE

[Got97] Don Gotterbarn. An “off by one” error. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):10–11, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1998:ECP

[Got98a] Don Gotterbarn. The ethical computer practitioner-licensing the moral community: a proactive approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):8–10, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1998:RAC

[Got98b] Don Gotterbarn. Reconstructing the ACM code of ethics and teaching computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):9–11, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1999:ICE

[Got99a] Don Gotterbarn. Integration of computer ethics into the CS curriculum: attachment or synthesis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (4):13–14, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:1999:TAC

[Got99b] Don Gotterbarn. Two approaches to computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):11–12, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Granger:1991:ICA

[GP91] Mary J. Granger and Roger A. Pick. The impact of computeraided software engineering on student performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):62–72, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gojenola:1995:IAC

[GPJ95] Koldo Gojenola, Tomás A. Pérez, and Arturo Jaime. Integrative assignments for CS1 and CS2 through libraries of abstract data types. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):47–49, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gerhardt-Powals:1999:DMC

[GPP99] Jill Gerhardt-Powals and Matthew H. Powals. The Digital Millennium Copyright Act: an international assault on fair use? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):191, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gramond:1999:UJI

[GR99] Eric Gramond and Susan H. Rodger. Using JFLAP to interact with theorems in automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):336–340, March 1999. CODEN SIGSD3. ISBN 1-58113-085-6. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/1999.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/Pape.bib.

Gray:1993:EEE

[Gra93] John S. Gray. Is eight enough?: the Eight Queens Problem reexamined. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):39–44, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greenleaf:1990:CDT

[Gre90] Newcomb Greenleaf. Computability and data types. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):219–223, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:1996:CLU

[Gre96a] Tony Greening. Command-line usage in a programming exam. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):55–59, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:1996:TLE

[Gre96b] Tony Greening. Teaching and learning essential computer science skills: the UNIX example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):21–24, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:1997:ESL

[Gre97] Tony Greening. Examining student learning of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):63–66, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:1999:GSC

[Gre99] Tony Greening. Gender stereotyping in a computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):203–207, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grillmeyer:1999:IMT

[Gri99] Oliver Grillmeyer. An interactive multimedia textbook for introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):286– 290, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grodzinsky:1997:CAS

[Gro97a] Frances S. Grodzinsky. Computer access for students with disabilities: an adaptive technology laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):292–295, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grove:1997:PIS

[Gro97b] Ralph F. Grove. A proposal for integrated software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2):11–ff., June 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grove:1998:UPS

[Gro98] Ralph F. Grove. Using the personal software process to motivate good programming practices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 98–101, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Galpin:1993:GIC

[GS93] Vashti Galpin and Ian Sanders. Gender imbalances in computer science at the University of the Witwatersrand. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):2–4, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goczyla:1996:SQL

[GS96] Krzysztof Goczyla and Stanisław Szejko. Software quality lab in the informatics option of Gdañsk TU electronics faculty. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):133–135, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:1999:TRP

[GS99] David Ginat and Eyal Shifroni. Teaching recursion in a procedural environment — how much should we emphasize the computing model? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):127–131, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gschwind:1994:RHE

[Gsc94] Michael Gschwind. Reprogrammable hardware for educational purposes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):183–187, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grout:1996:EC

[GSH96] Jarrell C. Grout, Robert G. Strader, and John B. Hanks. Essential C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):3–14, June 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:1995:TLT

[GSK^{+}95] David Gries, Fred B. Schneider, Joan Krone, Stan Warford, and J. Peter Weston. Teaching as a logic tool (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):384–385, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gould:1999:GDI

[GSvD99] Daniel L. Gould, Rosemary M. Simpson, and Andries van Dam. Granularity in the design of interactive illustrations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):306–310, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goelman:1998:TMD

[GSW^{+}98] Don Goelman, Roberta Evans Sabin, Marty J. Wolf, Pete Knoke, and Mike Murphy. The terminal Master’s degree (panel): does it need to be cured? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):355–356, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goodrich:1998:TAA

[GT98] Michael T. Goodrich and Roberto Tamassia. Teaching the analysis of algorithms with visual proofs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):207–211, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gitchell:1999:SUD

[GT99a] David Gitchell and Nicholas Tran. Sim: a utility for detecting similarity in computer programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):266–270, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goodrich:1999:URT

[GT99b] Michael T. Goodrich and Roberto Tamassia. Using randomization in the teaching of data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):53–57, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:1999:ETI

[Gui99] Mario A. M. Guimaraes. Experience teaching an introduction to DBMS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):48–49, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gurwitz:1995:AUI

[Gur95] Chaya Gurwitz. Achieving a uniform interface for binary tree implementations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):66–70, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gurwitz:1998:IMT

[Gur98] Chaya Gurwitz. The Internet as a motivating theme in a math/ computer core course for nonmajors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 68–72, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guvenir:1995:OOT

[Güv95] H. Altay Güvenir. An object-oriented tutoring system for teaching sets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):39–46, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guzdial:1995:CMS

[Guz95] Mark Guzdial. Centralized mindset: a student problem with object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):182– 185, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Granger:1994:CCE

[GVLZ94] Mary J. Granger, Mary R. Vorgert, Joyce Currie Little, and Lois Zells. Creating an I-CASE environment (abstract): pitfalls and promises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):376–377, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1997:CEC

[GY97a] Judith Gersting and Frank H. Young. Content + experiences = curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):325–329, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1997:PAT

[GY97b] Judith L. Gersting and Frank H. Young. Project-after they are finished. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):24, December 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1998:CWS

[GY98a] Judith L. Gersting and Frank H. Young. Contributions of the working student. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):26–27, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1998:SLC

[GY98b] Judith L. Gersting and Frank H. Young. Service learning via the computer science club. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):25–26, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1999:SSE

[GY99a] Judith L. Gersting and Frank H. Young. Sharpening subjective evaluation skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):26, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1999:SM

[GY99b] Judith L. Gersting and Frank H. Young. Student mentors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):28–29, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:1993:BIV

[GYT^{+}93] Judith L. Gersting, Frank Young, Eric Tinsley, Helen Hays, and Anthony Schaeffer. Bringing the industry view of software engineering to the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 307, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Green:1994:TTS

[GZ94] Jim Green and Stacey B. Zaremba. A target tracking system applied to swimming rats: an interdisciplinary project in computer science and psychology (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):408, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grundke:1996:SVD

[GZ96] E. W. Grundke and Zhengya Zhu. Simulator for visualizing data link layer protocols. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):235, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harlan:1995:PTP

[HA95] Robert M. Harlan and Joseph G. Akulis. Parallel threads: parallel computation labs for CS 3 and CS 4. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):141–145, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hirshfield:1994:OOP

[HAB^{+}94] Stuart Hirshfield, Owen Astrachan, John Barr, Karen Donnelly, David Levine, and Mark McGinn. Object-oriented programming (abstract): how to “scale up” CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):396, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadjerrouit:1998:CFI

[Had98a] Said Hadjerrouit. A constructivist framework for integrating the Java paradigm into the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):105–107, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadjerrouit:1998:JFP

[Had98b] Said Hadjerrouit. Java as first programming language: a critical evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):43–47, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadjerrouit:1999:CAO

[Had99] Said Hadjerrouit. A constructivist approach to object-oriented design and programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):171–174, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hafen:1994:DWS

[Haf94] Marguerite Hafen. Developing writing skills in computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):268–270, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haggard:1991:PDS

[Hag91] Gary Haggard. A project for data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):140–145, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hosny:1992:MPI

[HAG92] H. M. Hosny, M. S. Akabawy, and T. G. Gough. Manpower profiling for information technology: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):1–10, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hainline:1996:DPT

[Hai96] Douglas Hainline. DATATUTOR: a package for teaching data representation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):234, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hailperin:1997:IFP

[Hai97] Max Hailperin. Introducing fixed-point iteration early in a compiler course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):258–261, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hannay:1992:HAS

[Han92] David G. Hannay. Hypercard automata simulation: finitestate, pushdown and Turing machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2): 55–58, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hunkins:1993:CGA

[HANWN93] Daltin R. Hunkins, Steven K. Andrianoff, Rosalee Nerheim-Wolfe, and Thomas L. Naps. Computer graphics across the CS curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):295, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1990:EMO

[Har90] Stephen J. Hartley. Experience with MINIX in an operating systems lab. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):34–38, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harlan:1991:SPC

[Har91] Robert M. Harlan. Searching in parallel: a case study with the single-source shortest path algorithm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):254–259, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1992:ELS

[Har92] Stephen J. Hartley. Experience with the language SR in an undergraduate operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (1):176–180, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harrison:1993:COS

[Har93] William S. Harrison. Co-oping at the super collider. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):50–60, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harlan:1994:ASA

[Har94a] Robert M. Harlan. The automated student advisor: a large project for expert systems courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 31–35, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1994:AOS

[Har94b] Stephen J. Hartley. Animating operating systems algorithms with XTANGO. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):344–348, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1996:HLM

[Har96] Stephen J. Hartley. A hypermedia lab manual for operating systems: using the network to teach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):8–10, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1997:UEN

[Har97] Janet Hartman. Undergraduate education and the national science foundation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):18–19, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1998:AYJ

[Har98a] Stephen J. Hartley. “Alfonse, your Java is ready!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):247–251, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1998:ERN

[Har98b] Janet Hartman. Educational research: a new arena for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):18–19, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1998:RVY

[Har98c] Janet Hartman. The reviewer’s view of your proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):19, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1999:AWH

[Har99a] Stephen J. Hartley. “Alfonse, wait here for my signal!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):58–62, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1999:GBW

[Har99b] Janet Hartman. Grant bytes on the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):20–21, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1999:PUN

[Har99c] Janet Hartman. Program updates from the national science foundation division of undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):22–23, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hastings:1994:SAI

[Has94] David Hastings. A scalable approach to integrating object oriented programming into the undergraduate liberal arts curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):407, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haynes:1995:ERU

[Hay95] S. M. Haynes. Explaining recursion to the unsophisticated. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):3–6, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haynes:1997:CHL

[Hay97] Christopher T. Haynes. Compiling: a high-level introduction using scheme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):253–257, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haynes:1998:EAA

[Hay98] Christopher T. Haynes. Experience with an analytic approach to teaching programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):350–354, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hendrix:1997:VDE

[HBC97] T. Dean Hendrix, Larry A. Barowski, and James H. Cross II. A visual development environment for multi-lingual curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):20–24, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hart:1998:FPF

[HCD^{+}98] Hal Hart, Jim Caristi, Robert Dewar, Mark Gerhardt, Drew Hamilton, Christopher Haynes, and Sam Rebelsky. The future of programming-are fundamental changes in computer science programs coming? (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):370– 371, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hitchner:1999:CGI

[HCGW99] Lew Hitchner, Steve Cunningham, Scott Grissom, and Rosalee Wolfe. Computer graphics: the introductory course grows up. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):341–342, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howard:1996:FLS

[HCL96] Richard A. Howard, Curtis A. Carver, and William D. Lane. Felder’s learning styles, Bloom’s taxonomy, and the kolb learning cycle: tying it all together in the CS2 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):227–231, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hewson:1999:MAS

[HDC99] John Hewson, Wendy Doube, and Michael Calagaz. A multimedia animated simulation generator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 131–134, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Headington:1995:RID

[Hea95] Mark R. Headington. Removing implementation details from C++ class declarations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):24–28, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heck:1994:DPP

[Hec94] Patrick Heck. Dynamic programming for pennies a day. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):213–217, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hedman:1999:CDL

[Hed99] Anders Hedman. Creating digital libraries together-collaboration, multimodality, and plurality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 147–150, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hein:1993:DLA

[Hei93] James L. Hein. A declarative laboratory approach for discrete structures, logic, and computability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (3):19–25, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heines:1999:ECW

[Hei99] Jesse M. Heines. Evaluating course Web sites and student performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):143–146, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:1990:DMP

[Hen90] Peter B. Henderson. Discrete mathematics as a precursor to programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):17–21, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hext:1994:MWP

[Hex94] Jan Hext. Mr Waring’s problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (4):13–16, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haller:1998:RWC

[HF98] S. M. Haller and T. V. Fossum. Retaining women in CS with accessible role models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):73–76, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hampel:1998:HTM

[HFKSM98] Thorsten Hampel, Ferdinand Ferber, Reinhard Keil-Slawik, and Wolfgang H. Müller. Hypermedia teaching of mechanics-MechANIma. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):112–116, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hall:1998:VLE

[HG98] Lynne Hall and Adrian Gordon. A virtual learning environment for entity relationship modelling. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):345–349, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haberman:1999:DLM

[HG99] Bruria Haberman and David Ginat. Distance learning model with local workshop sessions applied to in-service teacher training. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):64–67, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:1996:ETU

[HGPJ^{+}96] Stephen Hartley, Jill Gerhardt-Powals, David Jones, Colin Mc-Cormack, M. Dee Medley, Blaine Price, Margaret Reek, and Marguerite K. Summers. Enhancing teaching using the Internet: report of the working group on the World Wide Web as an interactive teaching resource. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI): 218–228, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holland:1997:AOM

[HGW97] Simon Holland, Robert Griffiths, and Mark Woodman. Avoiding object misconceptions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):131–134, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harrington:1995:JFY

[HH95a] Jan L. Harrington and Helen M. Hayes. A joint first year program for computer science and information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):121–125, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartel:1995:PLC

[HH95b] Pieter H. Hartel and L. O. Hertzberger. Paradigms and laboratories in the core computer science curriculum: An overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):13–20, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Misc/TUBScsd/1980.bib;   ftp://ftp.wins.
uva.nl/pub/computer-systems/functional/reports/ACMSIGCSE_
paradigms.ps.Z.

Hitz:1995:MVC

[HH95c] Martin Hitz and Marcus Hudec. Modula-2 versus C++ as a first programming language — some empirical results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):317–321, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hirmanpour:1995:DCC

[HHK95] Iraj Hirmanpour, Thomas B. Hilburn, and Andrew Kornecki. A domain centered curriculum: an alternative approach to computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):126–130, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Higginbotham:1993:ISR

[Hig93] T. F. Higginbotham. The integer square root of N via a binary search. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):41–45, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hillam:1990:IAP

[Hil90] Bruce P. Hillam. Integrating an array processor into a “hands on” computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2): 11–14, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilzer:1992:CS

[Hil92] Ralph C. Hilzer, Jr. Concurrency with semaphores. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):45–50, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:1993:TAT

[Hil93] Thomas B. Hilburn. A top-down approach to teaching an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 58–62, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:1996:IFS

[Hil96] Thomas B. Hilburn. Inspections of formal specifications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):150–154, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hitchner:1999:ATG

[Hit99] Lewis E. Hitchner. An automatic testing and grading method for a C++ list class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):48–50, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harris:1992:RDC

[HJ92] William C. Harris and Leon P. Johnson. Redesign of an A.S. degree in computer science to meet emerging national standards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):2–6, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hutchens:1996:UIE

[HK96] David H. Hutchens and Elizabeth E. Katz. Using iterative enhancement in undergraduate software engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):266–270, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hitz:1997:TCW

[HK97] Martin Hitz and Stefan Kögeler. Teaching C++ on the WWW. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):11–13, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harrington:1993:CES

[HKH^{+}93] Susan Marie Harrington, Charles Kelemen, Rachelle Heller, Sandoval Melim, and Ellen Spertus. Creating an environment for the success of women students in undergraduate, co-ed computer science programs (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 290, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howard:1997:LSM

[HKSD97] Richard A. Howard, Lisa Kaczmarczyk, Frederick Springsteel, and Nell Dale. Learning style models and computer science education (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):383, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

House:1994:ASC

[HL94] Donald House and David Levine. The art and science of computer graphics: a very depth-first approach to the nonmajors course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):334–338, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hoggarth:1998:ASD

[HL98] Gil Hoggarth and Mike Lockyer. An automated student diagram assessment system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):122–124, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Higginbotham:1991:STC

[HM91] C. William Higginbotham and Ralph Morelli. A system for teaching concurrent programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):309–316, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:1998:SEA

[HMBO98] Thomas B. Hilburn, Susan Mengel, Donald J. Bagert, and Dale Oexmann. Software engineering across computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):117–121, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henry:1999:USD

[HML^{+}99] Sallie Henry, Nancy Miller, Wei Li, Joseph Chase, and Todd Stevens. Using software development teams in a classroom environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):356–357, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hays:1990:SPR

[HMOS90] James H. Hays, Leland Miller, Bobbie Othmer, and Mohammad Saeed. Simulation of process and resource management in a multiprogramming operating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):125–128, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ho:1993:ACC

[HMS93] Chi Fai Ho, Christopher L. Morgan, and Istvan Simon. An advanced classroom computing environment and its applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):228–231, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hoft:1991:IND

[Höf91] Hartmut Höft. Implementation of a non-deterministic loop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):24–28, June 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Math/maple-extract.bib.

Holt:1994:IUO

[Hol94] Richard C. Holt. Introducing undergraduates to object orientation using the Turing language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):324–328, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holliday:1997:SCI

[Hol97] Mark A. Holliday. System calls and interrupt vectors in an operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):53–57, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holliday:1998:WBI

[Hol98a] Mark A. Holliday. A Web-based introduction to computer networks (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):282, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hollinworth:1998:DMC

[Hol98b] Nick D. P. Hollinworth. The design of multimedia courseware (poster): preliminary survey results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):283, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holmboe:1999:CFK

[Hol99] Christian Holmboe. A cognitive framework for knowledge in informatics: the case of object-orientation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):17–20, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hosch:1996:JST

[Hos96a] F. Hosch. Java as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):45–50, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hosch:1996:JFL

[Hos96b] Frederick Hosch. Java as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):45–50, September 1, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Compiler/FORTRAN/fortran3.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/JAVA/java.bib;
 ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Misc/Pape.bib.

Houle:1997:EPV

[Hou97] Michael E. Houle. Ethics, programming, and virtual environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):91–93, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howell:1993:EWU

[How93] Kathy Howell. The experience of women in undergraduate computer science: what does the research say? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):1–8, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howatt:1994:CGS

[How94] James W. Howatt. On criteria for grading student programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):3–7, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harlan:1993:PSE

[HP93] Robert M. Harlan and David M. Patrone. Parsing as search: an easy-to-understand RTN interpreter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (3):26–30, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herrmann:1994:ISB

[HP94] Nira Herrmann and Jeffrey L. Popyack. An integrated, software-based approach to teaching introductory computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):92–96, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herrmann:1995:CAL

[HP95] Nira Herrmann and Jeffrey L. Popyack. Creating an authentic learning experience in introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):199–203, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Homkes:1993:CC

[HR93] Rick Homkes and John Minor Ross. Clarifying ’C’. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):45–51, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howerton:1999:AYP

[HRBK99] Charles P. Howerton, Mary Ann Robbert, Carl E. Bredlau, and Peter J. Knoke. Addressing the Y2K problem in the computing classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):364–365, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1991:TCP

[HS91] Janet Hartman and Dean Sanders. Teaching a course in parallel processing with limited resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):97–101, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1993:DPP

[HS93] Janet Hartman and Dean Sanders. Data parallel programming: a transition from serial to parallel computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):96–100, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holmes:1997:ASS

[HS97] Goefrey Holmes and Tony C. Smith. Adding some spice to CS1 curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):204–208, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hagan:1998:VDC

[HS98] Dianne Hagan and Judy Sheard. The value of discussion classes for teaching introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):108–111, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanchey:1990:MST

[HSC^{+}90] Cindy Hanchey, Marguerite K. Summers, Carol Chrisman, Joyce Currie Little, and Richard A. Lejk. Managing student teams in information systems courses (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):265, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hagan:1997:MER

[HSM97] Dianne Hagan, Judy Sheard, and Ian Macdonald. Monitoring and evaluating a redesigned first year programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):37–39, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haberman:1998:ADT

[HSS98] Bruria Haberman, Zahava Scherz, and Ehud Shapiro. Abstract data types as a project development organiser. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):102–104, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:1997:DQW

[HT97] Thomas B. Hilburn and Massood Towhidnejad. Doing quality work: the role of software process definition in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):277–281, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hou:1998:APS

[HT98] Lily Hou and James Tomayko. Applying the personal software process in CS1: an experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):322– 325, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haddad:1997:ME

[HTW97] Hisham Haddad, Herbert Tesser, and Steven P. Wartik. Megaprogramming education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):282–286, March 1997. CODEN SIGSD3. ISBN 0-89791-889-4. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1997.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
SE/cbsd.bib.

Haliburton:1998:GDP

[HTW98] William Haliburton, Mack Thweatt, and Nancy J. Wahl. Gender differences in personality components of computer science students: a test of Holland’s congruence hypothesis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):77–81, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hughes:1990:ULC

[Hug90] Larry Hughes. Using a low-cost communications tool in data communications courses (tutorial session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):259, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hughes:1992:TOS

[Hug92] Larry Hughes. Teaching operating systems using Turbo C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):181–186, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huss:1995:LPP

[Hus95] Janice E. Huss. Laboratory projects for promoting hands-on learning in a computer security course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2): 2–6, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hvorecky:1990:CBP

[Hvo90] Jozef Hvorecký. On a connection between programming and mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):53–54, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:1990:RWS

[HW90] Janet Hartman and Curt M. White. “real world” skills vs. “school taught” skills for the undergraduate computer major. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):216–218, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hill:1991:CAP

[HW91] Jane C. Hill and Andrew Wayne. A CYK approach to parsing in parallel: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):240–245, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartfield:1992:LHD

[HWB92] B. Hartfield, Terry Winograd, and John Bennett. Learning HCI design: mentoring project groups in a course on humancomputer interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):246–251, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hwang:1992:MIS

[HWS92] Betty W. Hwang, Andrew B. Whinston, and Wilhelmina Savenye. Management information systems curricula in the United States and the Republic of China: a comparative study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):113–122, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1996:PST

[IAK96] John Impagliazzo, Elizabeth S. Adams, and Karl J. Klee, editors. Proceedings of the 27th SIGCSE Technical Symposium on Computer Science Education, 1996, Philadelphia, Pennsylvania, USA, February 15–17, 1996, volume 28(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1996. CO-DEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86.

Impagliazzo:1994:IPC

[IBF^{+}94] John Impagliazzo, J. Dennis Bjornson, Dennis J. Frailey, Jeanette Horan, and Gerald H. Thomas. An industry perspective on computer science accreditation (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):379–380, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1990:ICS

[ICD^{+}90] John Impagliazzo, Helene Chlopan, Ronald M. Davis, David M. Hata, and Karl Klee. Issues in the computing sciences at twoyear colleges (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):258, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1997:PIC

[IG97] John Impagliazzo and Michael Goldweber. Perspectives on innovations in the computing curriculum (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):146, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ingargiola:1994:RST

[IHA^{+}94] Giorgio Ingargiola, Nathan Hoskin, Robert Aiken, Rajeev Dubey, Judith Wilson, Mary-Angela Papalaskari, Margaret Christensen, and Roger Webster. A repository that supports teaching and cooperation in the introductory AI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):36–40, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1998:RPD

[IMAM98] John Impagliazzo, Mark Mandelbaum, Robert Aiken, and James E. Miller. Refereed papers, the digital library, and the future of the SIGCSE Bulletin (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):359–360, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1999:PAS

[Imp99a] John Impagliazzo, editor. Proceedings of the 4th Annual SIGCSE/SIGCUE Conference on Innovation and Technology in Computer Science Education (ITICSE-99), 27-June–1-July 1999, Cracow, Poland, volume 31(3) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:1999:TVL

[Imp99b] John Impagliazzo. Teaching very large instruction word architectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):208, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:1991:MCC

[Jac91] David Jackson. A mini-course on concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):92–96, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jacker:1996:TST

[Jac96] Kenneth H. Jacker. Tkined/scotty: tools for network documentation, monitoring and troubleshooting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):233, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Janser:1997:ILS

[Jan97] Achim W. Janser. An interactive learning system visualizing computer graphics algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):21–23, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jarc:1994:DSU

[Jar94] Duane J. Jarc. Data structures: a unified view. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):2–4, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1998:CSA

[JBCL98] Lawrence G. Jones, Keith Barker, Susan Conry, and Doris Lidtke. Computer science accreditation (panel): current status and future directions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):363–364, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jawary:1997:II

[JBV97] Anita Jawary, Christiana Birchak, and Susan Strack Vargo. International initiative. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):85–87, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1993:ICP

[JCF^{+}93] Rhys Price Jones, Doug Cooper, Dan Friedman, Ric Holt, and Peter Robinson. Issues in the choice of programming language for CS 1 (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):301, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jerinic:1997:OME

[JD97a] Ljubomir Jerinic and Vladan Devedzic. OBOA model of explanation module in intelligent tutoring shell. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):133–135, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jovanovic-Dolecek:1997:FDP

[JD97b] Gordana Jovanovic-Dolecek. FASRS-demo package for Fourier analysis of stationary random signals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 46–53, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jeffrey:1991:UPN

[Jef91] John M. Jeffrey. Using Petri nets to introduce operating system concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):324–329, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jenkins:1998:PAT

[Jen98] Tony Jenkins. A participative approach to teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):125–129, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jessup:1994:TPC

[Jes94] Peter G. Jessup. Teaching parallel computing with multi-pascal (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):408, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:1997:UAE

[JF97] David Jackson and Andrew Fovargue. The use of animation to explain genetic algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):243– 247, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnsonbaugh:1991:GGS

[JK91] Richard Johnsonbaugh and Martin Kalin. A graph generation software package. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):151–154, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johansen:1996:ILG

[JKB96] Mary Johansen, Jason Kapusta, and Doug Baldwin. Interactive learning with gateway labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):232, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joy:1998:EEM

[JL98] Mike Joy and Michael Luck. Effective electronic marking for online assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):134–138, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:1998:VOH

[JLH^{+}98] Jeremy Johnson, Y. N. Lakshman, Thomas T. Hewett, Tim Souder, Tom Fitzgerald, Sara Donegan, and Paul Morgovsky. Virtual office hours using TechTalk, a Web-based mathematical collaboration tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):130–133, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:1999:PTS

[JLP99] Daniel T. Joyce, John Lewis, and Jane Prey, editors. The proceedings of the Thirtieth SIGCSE Technical Symposium on Computer Science Education: March 24-March 28, 1999, New Orleans, Louisiana, volume 31(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27 .A34 1999.

Johnsonbaugh:1990:CPL

[JM90] Richard Johnsonbaugh and David P. Miller. Converses of pumping lemmas. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):27–30, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1991:TCS

[JM91] Warren T. Jones and F. Mike McGuirt. Telecommunications and computer science: two merging paradigms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):13–22, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:1996:AAN

[JM96] David Jackson and Ian G. Morton. Algorithm animation of neural networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):39–41, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/Pape.bib.

Jackson:1997:SEI

[JMM97] Ursula Jackson, Bill Z. Manaris, and Renée A. McCauley. Strategies for effective integration of software engineering concepts and techniques into the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):360–364, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

John:1992:IPC

[Joh92] D. J. John. Integration of parallel computation into introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):281–285, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

John:1994:NSP

[Joh94] David J. John. NSF supported projects: parallel computation as an integrated component in the undergraduate curriculum in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):357–361, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:1997:ISE

[Joh97] Hubert A. Johnson. Integrating software engineering into the traditional computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2):39–ff., June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johansson:1998:PEP

[Joh98] Per-Gunnar Johansson. Programming by example (poster): an instructional approach allowing introductory students to quickly grasp the power and excitement of programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):284, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johann:1999:FTH

[Joh99a] Patricia Johann. A funny thing happened on the way to the formula: demonstrating equality of functions and programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):32–34, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:1999:EST

[Joh99b] Hubert A. Johnson. Evaluation of the SEPA in teaching undergraduate software engineering in the traditional computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):81–83, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1996:CDE

[Jon96a] David Jones. Computing by distance education: problems and solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):139–146, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1996:RJA

[Jon96b] David Jones. RCOS.java: an animated operating system for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):233, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:1990:VLA

[Joy90] Daniel T. Joyce. A virtual lab to accompany CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):40–43, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:1998:CPS

[Joy98] Daniel Joyce. The computer as a problem solving tool: a unifying view for a non-majors course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 63–67, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1999:ABC

[JPKPM99] Ricardo Jiménez-Peris, Sami Khuri, and Marta Patiño-Martínez. Adding breadth to CS1 and CS2 courses through visual and interactive programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):252–256, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1996:GVE

[JPPFPMVI96] Ricardo Jiménez-Peris, Cristóbal Pareja-Flores, Marta Patiño-Martínez, and J. Ángel Velázquez-Iturbide. Graphical visualization of the evaluation of functional programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):36–38, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1997:LMT

[JPPFPMVI97] Richardo Jiménez-Peris, Cristóbal Pareja-Flores, Marta Patiño-Martínez, and J. Ángel Velázquez-Iturbide. The locker metaphor to teach dynamic memory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):169–173, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1996:MIV

[JPPM96] Ricardo Jiménez-Peris and Marta Patiño-Martínez. A Modula-2 interpreter/visualizer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):232, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Peris:1996:DML

[JPPMMF96] Ricardo Jiménez-Peris, Marta Patiño-Martínez, and Jesús M. Milán-Franco. DD-Mod: a library for teaching distributed programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):229, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackowitz:1990:TWR

[JPS90] Paul M. Jackowitz, Richard M. Plishka, and James R. Sidbury. Teaching writing and research skills in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):212–215, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jarvinen:1999:BTF

[JPT^{+}99] Kimmo Järvinen, Tuukka Pienimäki, Tommi Teräsvirta, John Joel Kyaruzi, and Erkki Sutinen. Between Tanzania and Finland: learning Java over the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):217–221, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:1996:WBL

[JRS96] Rhys Price Jones, Fritz Ruehr, and Richard Salter. Webbased laboratories in the introductory curriculum enhance formal methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):160–164, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jipping:1990:CDP

[JTW90] Michael J. Jipping, Jeffrey R. Toppen, and Stephen Weeber. Concurrent distributed Pascal: a hands-on introduction to parallelism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):94–99, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:1997:GSP

[JU97] David Jackson and Michelle Usher. Grading student programs using ASSYST. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):335–339, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Juhasz:1999:USS

[Juh99] Zoltan Juhasz. Using spreadsheets as a simple and effective teaching tool for predicting and visualizing parallel program performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):51–54, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Julstrom:1992:SSW

[Jul92] Bryant A. Julstrom. Slow sorting: a whimsical inquiry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):11–13, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jin:1995:LTP

[JY95] Lan Jin and Lan Yang. A laboratory for teaching parallel computing on parallel structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):71–75, March 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1995.bib.

Kolesar:1995:TCS

[KA95] Mary Veronica Kolesar and Vicki H. Allan. Teaching computer science concepts and problem solving with a spreadsheet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):10–13, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:1994:AWN

[KACL94] Karl Klee, Richard Austing, Robert Campbell, and Joyce Currie Little. Articulation: who needs it? your students do! (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):403, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khera:1993:IPC

[KAK93] Vivek Khera, Owen Astrachan, and David Kotz. The Internet Programming Contest: a report and philosophy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):48–52, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kwiatkowski:1996:TPP

[KAL^{+}96] Jan Kwiatkowski, Marek Andruszkiewicz, Emilio Luque, Tomas Margalef, Jose Cunha, João Lourenço, Henryk Krawczyk, and Stanisław Szejko. Teaching parallel processing: development of curriculum and software tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):159–161, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaminski:1992:IFP

[Kam92] D. M. Kaminski. Introducing the fuzzy paradigm using Prolog. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):202–206, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Katsinis:1994:DMP

[Kat94] Constantine Katsinis. The development of a multi-processor personal computer in a senior computer design laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):349–352, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Katz:1995:PSH

[Kat95] Kaila Katz. The present state of historical content in computer science texts: a concern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):43–50, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1992:CCL

[Kay92a] D. G. Kay. A course in computer law. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (1):252–254, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1992:BAF

[Kay92b] David G. Kay. A balanced approach to first-year computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):15–18, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1993:HCS

[Kay93] David G. Kay. An honors computer science seminar for undergraduate non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):141–144, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1995:TCS

[Kay95] David G. Kay. Training computer science teaching assistants: a seminar for new TAs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):53–55, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1996:BCH

[Kay96] David G. Kay. Bandwagons considered harmful, or the past as prologue in curriculum change. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):55–58, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1998:CSC

[Kay98a] David G. Kay. Computer scientists can teach writing: an upper division course for computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):117–120, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1998:LICa

[Kay98b] David G. Kay. Large introductory computer science classes: strategies for effective course management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):131–134, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

King:1997:CSA

[KB97] L. A. Smith King and John Barr. Computer science for the artist. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):150–153, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:1999:AGA

[KBCP99] Karl J. Klee, Nancy Burns, Fay Cover, and Judith Porter. ACM guidelines for associate and certificate level programs in computer information technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):

359, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kopec:1999:HSD

[KCA99] Danny Kopec, Richard Close, and Jim Aman. How should data structures and algorithms be taught. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 175–176, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1998:LICb

[KCC^{+}98] David G. Kay, Jacobo Carrasquel, Michael J. Clancy, Eric Roberts, and Joseph Zachary. Large introductory courses in research computer science departments (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):374–375, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kiper:1996:PAT

[KCD^{+}96] James D. Kiper, Valerie Cross, Diane Delisio, Ann Sobel, and Douglas Troy. Perspectives on assessment through teaching portfolios in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):200– 203, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1998:CSD

[KCPC98] Barry L. Kurtz, Hong Cai, Chris Plock, and Xijia Chen. A concurrency simulator designed for sophomore-level instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):237–241, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koneva:1990:EVN

[KD90] Luisa Koneva and Jordan Denev. EASY/VI — a new instructional computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):55–58, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1993:EEC

[KDE^{+}93] Barry L. Kurtz, Nell Dale, Jerry Engel, Jim Miller, Keith Barker, and Harriet Taylor. Evaluating effectiveness in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):293, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Keating:1993:HNN

[Kea93] John G. Keating. Hopfield networks, neural data structures and the nine flies problem: neural network programming projects for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):33–37, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kelsh:1993:LAC

[Kel93] James P. Kelsh. Levels of abstraction in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):35–37, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1999:PRP

[KFD^{+}99] Deborah Knox, Sally Fincher, Nell Dale, Elizabeth Adams, Don Goelman, James Hightower, Ken Loose, and Fred Springsteel. The peer review process of teaching materials: Report of the ITiCSE’99 Working Group on Validation of the Quality of Teaching Materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):87–100, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Techreports/UKent.bib; http://www.cs.kent.
ac.uk/pubs/1999/962.

Kallman:1994:TEC

[KG94] Ernest A. Kallman and John P. Grillo. Teaching ethics in IS courses (abstract): everything you always wanted to know but were afraid to ask. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):389, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1999:DTR

[KG99] Deborah Knox and Scott Grissom. Developing teaching resources for reuse and publishing in the CSTC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):370, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kuras:1999:CCM

[KGZ99] Marian Kuraś, Mariusz Grabowski, and Agnieszka Zajac. Changing IS curriculum and methods of instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):36–39, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kearsley:1992:TIS

[KH92] Greg Kearsley and Rachelle Heller. Teaching instructional software design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):24–28, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Konstam:1994:TCS

[KH94] Aaron Konstam and John E. Howland. Teaching computer science principles to liberal arts students using scheme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):29–ff., December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:1999:VCS

[KH99] Sami Khuri and Hsiu-Chin Hsu. Visualizing the CPU scheduler and page replacement algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):227–231, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khosraviyani:1990:UBS

[Kho90] Firooz Khosraviyani. Using binary search on a linked list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):7–10, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:1994:IGR

[Khu94] Sami Khuri. Intractability: a geometric representation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):228–232, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:1993:FRA

[KI93] Karl J. Klee and John Impagliazzo. Final report (abstract): curricula for two-year degree programs in computer sciences, and computing and engineering technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):302, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

King:1992:EPL

[Kin92] K. N. King. The evolution of the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):213–219, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

King:1998:EST

[Kin98] C. King. An evaluation of strategies for teaching technical computing topics to students at different levels (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):285, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kirsch:1996:TOA

[Kir96] Raymond P. Kirsch. Teaching OLE automation: a problembased learning approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):68–72, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1992:USG

[KJ92] B. L. Kurtz and J. B. Johnston. Using the synthesizergenerator to teach principles of programming language semantics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):207–212, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1997:DLM

[KJ97] Deborah L. Knox and Daniel T. Joyce. Designing laboratory materials for computing courses (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):143, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Karpouzis:1998:RPC

[KK98a] Kostas Karpouzis and Stefanos Kollias. The rendering pipeline in the classroom: a diversified approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):139–142, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1998:PBI

[KK98b] Judy Kay and Bob Kummerfeld. A problem-based interface design and programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):194– 197, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1998:PCU

[KKA98] Barry L. Kurtz, Chinhyun Kim, and Jamal Alsabbagh. Parallel computing in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):212–216, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kula:1990:IPE

[KKC90] Timothy S. Kula, Raymond Konopka, Jr., and John A. Cicero. Image processing experiments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):167– 170, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koshelev:1999:EAM

[KKL99] Misha Koshelev, Vladik Kreinovich, and Luc Longpré. Encryption algorithms made natural. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):50–51, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:1995:RFY

[KKR95] Michael Kölling, Bett Koch, and John Rosenberg. Requirements for a first year object-oriented teaching language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):173–177, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kiper:1992:USE

[KLE92] James Kiper, Michael J. Lutz, and Henry A. Etlinger. Undergraduate software engineering laboratories: a progress report from two universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):57–62, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:1999:RAA

[Kle99a] Karl J. Klee. Recent activities of the ACM two-year college education committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):31, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klein:1999:MSC

[Kle99b] Bruce J. Klein. Message from the SIGCSE chair. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):2, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:1998:ALP

[KLL^{+}98] Karl J. Klee, Joyce Currie Little, John Lawlor, Pamela Matthiesen, T. S. Pennington, Josephine Freedman, and Karen Richards. Associate-level programs for preparation of computer support personnel (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 373, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klemetti:1995:PPT

[KLMS95] Harri Klemetti, Ismo Lapinleimu, Erkki Mäkinen, and Mika Sieranta. A programming project: trimming the spring algorithm for drawing hypergraphs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3): 34–38, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klein:1993:PST

[KLY93] Bruce J. Klein, Cary Laxer, and Frank H. Young, editors. Proceedings of the 24th SIGCSE Technical Symposium on Computer Science Education, 1993, Indianapolis, Indiana, USA, February 18–19, 1993, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1993. CODEN SIGSD3. ISBN 0-89791-565-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1998:RLT

[KM98] Deepak Kumar and Lisa Meeden. A robot laboratory for teaching artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):341–344, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khosraviyani:1991:TDA

[KMH91] Firooz Khosraviyani, Mohammad H. Moadab, and Douglas F. Hale. Time distribution analysis for binary search of a linked list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):7–12, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1995:DIG

[KMO95] Barry L. Kurtz, Unmesh S. Mayekar, and Michael B. O’Neal. Design and implementation of a generalized problem solving assistant for algorithm development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):97–101, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1996:HPC

[Kno96] Deborah L. Knox. Historical perspectives of computing: an introductory lab for browsing on the World Wide Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):20–22, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1997:EAL

[Kno97a] Deborah L. Knox. Enhancing accessibility of lab materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):20–21, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1997:IDS

[Kno97b] Deborah L. Knox. Integrating design and simulation into a computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):42–44, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1997:LPC

[Kno97c] Deborah L. Knox. On-line publication of CS laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):340–344, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1998:IYB

[Kno98a] Deborah L. Knox. If you build it, they will come. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):22–23, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1998:RGN

[Kno98b] Deborah L. Knox. Repository growth and the near future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):21–23, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1999:CST

[Kno99a] Deborah L. Knox. The computer science teaching center. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):22–23, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1999:PGC

[Kno99b] Deborah L. Knox. Planned growth for the CSTC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):25–26, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1991:DIU

[KOC91] Barry L. Kurtz, Richard L. Oliver, and Edward M. Collins. The design, implementation, and use of DSTutor: a tutoring system for denotational semantics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):169–177, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolikant:1998:HMD

[Kol98] Yifat Ben-David Kolikant. “how much did you get?” (poster): the influence of algebraic knowledge of computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):287, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kooshesh:1995:SPU

[Koo95] Ali A. Kooshesh. Solving a problem using cooperating data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):61–65, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kortright:1994:SPI

[Kor94] Lisa M. Levy Kortright. From specific problem instances to algorithms in the introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 71–75, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kosa:1999:BPW

[Kos99] Martha J. Kosa. Beginners program Web page builders and verifiers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):185, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kotz:1995:DPP

[Kot95] David Kotz. A data-parallel programming library for education (DAPPLE). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):76–81, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kotze:1998:WHM

[Kot98] Paula Kotzé. Why the hypermedia model is inadequate for computer-based instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):148– 152, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kourik:1997:DCS

[Kou97] Janet Kourik. Developing critical skills in technical courses (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):390, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kourik:1998:PAI

[Kou98] Janet L. Kourik. A practical approach to internationalizing information systems & computer science courses (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):382, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1990:ISC

[KP90] Barry L. Kurtz and Thomas H. Puckett. Implementing a single classwide project in software engineering using Ada tasking for synchronization and communication. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22 (1):6–11, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Katz:1991:HOC

[KP91] Elizabeth E. Katz and Hayden S. Porter. HyperTalk as an overture to CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):48–54, March 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koldehofe:1998:BAD

[KPT98] Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. Building animations of distributed algorithms for educational purposes (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):286, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koldehofe:1999:DAV

[KPT99] Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. Distributed algorithms visualisation for educational purposes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):103–106, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knight:1994:UCS

[KPW94] John C. Knight, Jane C. Prey, and Wm. A. Wulf. Undergraduate computer science education: a new curriculum philosophy & overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):155–159, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:1996:BLT

[KR96a] Michael Kölling and John Rosenberg. Blue — a language for teaching object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):190–194, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:1996:OOP

[KR96b] Michael Kölling and John Rosenberg. An object-oriented program development environment for the first programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):83–87, March 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:1999:TTT

[KR99] Michael Kölling and John Rosenberg. Tools and techniques for teaching objects first in a Java course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 368, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krone:1992:SDM

[Kro92] Joan Krone. Student designed machines for a theory of computation course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):51–52, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krone:1996:USCa

[Kro96a] Joan Krone. Using symbolic computation for teaching data structures and algorithm analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):230, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krone:1996:USCb

[Kro96b] Joan Krone. Using symbolic computation for teaching data structures and algorithm analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (4):19–24, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kmoch:1994:CUS

[KS94] Joseph W. Kmoch and Mark Stehlik. The creation and use of scoring standards (rubrics): experiences with the advanced placement computer science exam (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):362–363, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:1998:APA

[KS98] Sami Khuri and Yanti Sugono. Animating parsing algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):232–236, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:1994:AGA

[KSIR94] David G. Kay, Terry Scott, Peter Isaacson, and Kenneth A. Reek. Automated grading assistance for student programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):381–382, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kitchen:1992:GPT

[KST92] Andrew T. Kitchen, Nan C. Schaller, and Paul T. Tymann. Game playing as a technique for teaching parallel computing concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):35–38, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Ai/minimax.bib.

Kluit:1998:VPJ

[KSW98] Peter G. Kluit, Marleen Sint, and Frank Wester. Visual programming with Java: evaluation of an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):143–147, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kubilus:1999:MTD

[Kub99] Norbert J. Kubilus. Multi-team development project experience in a core computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):186, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1996:FDT

[Kum96] Amruth N. Kumar. Fork diagrams for teaching selection in CS I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):348–352, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1998:EPC

[Kum98] Amruth N. Kumar. Evaluating the pedagogy of computer science courseware delivered over the Web (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):288, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1999:CWL

[Kum99] Amruth N. Kumar. On changing from written to on-line tests in Computer Science I: an assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 25–28, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:1991:LAS

[Kur91] Barry L. Kurtz. Laboratory activities for studying the formal semantics of programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):162–168, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kushan:1994:PPT

[Kus94] Barbara Kushan. Preparing programming teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):248–252, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kerner:1994:WCW

[KV94] Janet T. Kerner and Kathy Vargas. Women and computers: what we can learn from science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2): 52–56, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kick:1993:WCS

[KW93] Russell C. Kick, Jr. and F. Stuart Wells. Women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):203–207, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:1994:UBS

[KW94] Sami Khuri and Jason Williams. Understanding the bottomup SLR parser. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):339–343, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:1996:NAN

[KW96] Sami Khuri and Jason Williams. Neuralis: an artificial neural network package. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):25–27, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/Pape.bib.

Koffman:1999:CUJ

[KW99] Elliot Koffman and Ursula Wolz. CS1 using Java language features gently. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):40–43, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:1996:ULC

[KWJ^{+}96] Deborah Knox, Ursula Wolz, Daniel Joyce, Elliot Koffman, Joan Krone, Atika Laribi, J. Paul Myers, Viera K. Proulx, and Kenneth A. Reek. Use of laboratories in computer science education: guidelines for good practice: report of the working group on computing laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):167–181, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liffick:1996:NPS

[LA96] Blaise W. Liffick and Robert Aiken. A novice programmer’s support environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):49–51, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lodgher:1994:PAT

[LAH94] Akhtar Lodgher and Hisham Al-Haddad. A practical approach for teaching reuse in a data structures course using Ada (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):407, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lambert:1990:CTC

[Lam90] Kenneth A. Lambert. Compiling Tinyturing in a compiler construction course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):2–6, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lambert:1992:AOC

[Lam92] Kenneth A. Lambert. An ALGOL object code interpreter in scheme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):39–44, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lambert:1995:RIC

[Lam95] Joseph M. Lambert. Restructuring the introductory computer science course for engineers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):204– 208, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Langan:1993:MPD

[Lan93] David D. Langan. A multi-purpose dataflow simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):87–90, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lankewicz:1998:RTC

[Lan98] Linda B. Lankewicz. Resources for teaching computer networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):112–116, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LaRusch:1993:TAI

[LaR93] Michele R. LaRusch. Teaching artificial intelligence as the year 2000 approaches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):38–42, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Larmour:1997:SRA

[Lar97] R. Larmour. A survey into the relevance and adequacy of training in systems analysis and design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2): 54–64, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawhead:1997:MCO

[Law97a] Pamela B. Lawhead. A model for the creation of online courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):31–36, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawhead:1997:WDL

[Law97b] Pamela B. Lawhead. The Web and distance learning (panel): what is appropriate and what is not. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (3):144, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:1996:RFC

[LB96] Mei-Ling Liu and Lori Blanc. On the retention of female computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):32–36, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lau:1994:TIF

[LBJ94] Kung-Kiu Lau, Vicky J. Bush, and Pete J. Jinks. Towards an introductory formal programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):121–125, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leska:1996:MPC

[LBK96] Chuck Leska, John Barr, and L. A. Smith King. Multiple paradigms in CS I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):343–347, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lake:1990:SAP

[LC90] Al Lake and Curtis Cook. Style: an automated program style analyzer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):29–33, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lees:1996:ANL

[LC96a] B. Lees and J. Cowie. Applying natural language technology to the learning of operating systems functions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):11–13, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lemone:1996:ECE

[LC96b] Karen A. Lemone and Walter Ching. Easing into C++: experiences with RoBOTL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):45–49, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leach:1990:AOS

[Lea90] Ronald J. Leach. An advanced operating systems project using concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):39–ff., September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leach:1995:UME

[Lea95] Ronald J. Leach. Using metrics to evaluate student programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):41–43, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leach:1997:SEU

[Lea97] Ronald J. Leach. Some experiences using the Internet for a software development project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):51–53, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lupton:1990:CSR

[LEB^{+}90] William L. Lupton, Mary Ellis, Andrew Bernat, Benjamin Martin, Surrendar Pulusani, and Leroy Roquemore. Computer science research and instuction at institutions with large minority enrollments (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 262, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lebeck:1999:CCP

[Leb99] Alvin R. Lebeck. Cache conscious programming in undergraduate computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):247–251, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lopez:1993:REU

[LEDW93] Antonio M. Lopez, Jr., Gerald L. Engel, Herbert L. Dershem, and Santa Wiltz. Research experience for undergraduates (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):303, March 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leeper:1990:PCD

[Lee90] Robert Leeper. A project course in database. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):78–80, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1996:HCS

[Lee96] John A. N. Lee. History in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):15–20, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1997:HCS

[Lee97] John A. N. Lee. History in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):12–13, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1998:HCS

[Lee98a] John A. N. Lee. History in the computer science curriculum: part II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):11–13, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1998:NCC

[Lee98b] John A. N. Lee. Napier’s chessboard calculator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):11–12, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1999:FCE

[Lee99a] John A. N. Lee. The first computer — an ethical concern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):15–16, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1999:MC

[Lee99b] John A. N. Lee. Multiplying on the chessboard. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):13–14, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lelewer:1994:SCC

[Lel94] Debra A. Lelewer. A seminar course in computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):253–257, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leonard:1991:USE

[Leo91] James R. Leonard. Using a software engineering approach to CS1: a comparative study of student performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):23–26, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levine:1993:DDL

[Lev93] David B. Levine. Dealing with different levels of abstraction in a data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):261– 264, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levy:1995:CLU

[Lev95] Suzanne Pawlan Levy. Computer language usage in CS1: survey results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):21–26, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levin:1997:PDC

[Lev97] Martin H. Levin. A prototype for a data communications laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):179–183, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levin:1999:IDC

[Lev99a] Martin H. Levin. The implementation of a data communications laboratory in small to medium sized universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):187, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levin:1999:UST

[Lev99b] Martin H. Levin. Use of a soundcard in teaching audio frequency and analog modem communications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):79–83, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levitin:1999:DWT

[Lev99c] Anany Levitin. Do we teach the right algorithm design techniques? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):179–183, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawhead:1996:MDW

[LG96] Pamela B. Lawhead and Kathryn F. Gates. Managing the development of a Web-based project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):87–89, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Little:1999:IPW

[LGB^{+}99] Joyce Currie Little, Mary J. Granger, Roger Boyle, Jill Gerhardt-Powals, John Impagliazzo, Carol Janik, Norbert J. Kubilus, Susan K. Lippert, W. Michael McCracken, Grazyna Paliwoda, and Piotr Soja. Integrating professionalism and workplace: Issues into the computing and information technology curriculum: Report of the ITiCSE’99, Working Group on Professionalism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):106–120, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lidtke:1999:NDA

[LGBM99] Doris K. Lidtke, John Gorgone, Della Bonnette, and Michael C. Mulder. New developments in accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):345, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leidig:1993:ISC

[LGK^{+}93] Paul M. Leidig, Mary J. Granger, Asad Khailany, Joan Pierson, and Dean Sanders. Information systems curriculum (abstract): where we should be going? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 300, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lim:1992:DGD

[LH92] Billy B. L. Lim and Richard Hunter. DBTool: a graphical database design tool for an introductory database course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):24–27, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lowder:1999:WBS

[LH99] Jason Lowder and Dianne Hagan. Web-based student feedback to improve learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):151–154, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Logan:1990:ARA

[LHA90] Paula M. Logan, Clifford W. Hale, and John M. Atkins. An academic relational algebra database management system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):25–28, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Li:1998:THD

[Li98] Wing-Ning Li. Towards a hierarchical design and integration of programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):121–125, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liang:1995:TDP

[Lia95] Y. Daniel Liang. Teaching dynamic programming techniques using permutation graphs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):56–60, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lim:1998:TWD

[Lim98] Billy B. L. Lim. Teaching Web development technologies in CS/ IS curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):107–111, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Little:1990:CRH

[Lit90] Joyce Currie Little. Curriculum ’68 revisited — an historical perspective (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 264, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Livesey:1997:SSM

[Liv97] Mike Livesey. STAMPS: a state-machine based processor simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):267–271, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lovato:1995:PVD

[LK95] Mona E. Lovato and Michael F. Kleyn. Parser visualizations for developing grammars with yacc. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):345–349, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lambrix:1998:CSI

[LK98] Patrick Lambrix and Mariam Kamkar. Computer science as an integrated part of engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):153–156, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lloyd:1994:EBG

[Llo94] William S. Lloyd. Exploring the Byzantine generals problem with beginning computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):21–24, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Laribi:1995:ELA

[LLP95] Atika Laribi, Michel Léonard, and Olivier Parchet. An experiment in learning about basic information system concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):248–252, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LaSalle:1991:NFR

[LM91] A. J. La Salle and L. R. Medsker. An NSF funded research experience for undergraduates: intelligent systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):39–44, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lopez:1994:URP

[LM94] Antonio M. Lopez, Jr. and Kenneth C. Messa, Jr. An undergraduate research program in multi-paradigm software design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):271–275, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewandowski:1998:CST

[LM98a] Gary Lewandowski and Amy Morehead. Computer science through the eyes of dead monkeys: learning styles and interaction in CS I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):312–316, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:1998:CBN

[LM98b] Stuart Lewis and Gaius Mulley. A comparison between novice and experienced compiler users in a learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):157–161, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lidtke:1998:ICC

[LM98c] Doris K. Lidtke and Michael C. Mulder. Information centric curriculum (ISC’98) (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 380, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lelewer:1991:HCD

[LN91] Debra A. Lelewer and Cheng Ng. An honors course in data compression. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):146–150, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lutz:1997:RLT

[LN97] Michael J. Lutz and J. Fernando Naveda. The road less traveled: a baccalaureate degree in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):287–291, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leung:1998:DWW

[LN98] Raymond M. W. Leung and Eugenia M. W. Ng. Does World Wide Web provide better resources than library for learning — a case study (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):290, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Locklair:1991:ICS

[Loc91] Gary H. Locklair. The introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):235–239, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1998:PVD

[LP98] Pete Lee and Chris Phillips. Programming versus design (poster): teaching first year students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):289, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lomic:1999:DEC

[LP99] Marijana Lomic and Zoran Putnik. On distance education courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):194, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:1998:PST

[LPJI98] John Lewis, Jane Prey, Daniel T. Joyce, and John Impagliazzo, editors. Proceedings of the 29th SIGCSE Technical Symposium on Computer Science Education, 1998, Atlanta, Georgia, USA, February 26–March 1, 1998, volume 30(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1998. CO-DEN SIGSD3. ISBN 0-89791-994-7. ISSN 0097-8418 (print), 2331-3927 (electronic).

Luker:1994:PAI

[LR94] Paul A. Luker and Dennis Rothermel. The philosophy of artificial intelligence: a general studies course with particular benefits to computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 41–45, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lang:1993:SSL

[LS93] Joseph E. Lang and Barbara A. Smith. Scheduled supervised laboratories in CS1: a comparative analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):185–188, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Luque:1992:QAT

[LSS92] E. Luque, R. Suppi, and J. Sorribes. A quantitative approach for teaching parallel computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1): 286–298, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Luque:1996:PSD

[LSS^{+}96] E. Luque, J. Sorribes, R. Suppi, E. Cesar, J. L. Falguera, and M. Serrano. Parallel systems development in education: a guided method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):156–158, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Luker:1994:TMO

[Luk94] Paul A. Luker. There’s more to OOP than syntax! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):56–60, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:1997:TCL

[LW97] Greg C. Lee and Chen-Chih Wu. On teaching computer literacy to future secondary school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (2):2–6, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Long:1998:PIF

[LWB^{+}98] Timothy J. Long, Bruce W. Weide, Paolo Bucci, David S. Gibson, Joe Hollingsworth, Murali Sitaraman, and Steve Edwards. Providing intellectual focus to CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):252–256, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Long:1999:CVF

[LWBS99] Timothy Long, Bruce Weide, Paolo Bucci, and Murali Sitaraman. Client view first: an exodus from implementation-biased teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):136–140, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lin:1996:CCD

[LWC96] Janet Mei-Chuen Lin, Cheng-Chih Wu, and Guey-Fa Chiou. Critical concepts in the development of courseware for CS closed laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):14–19, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Laxer:1995:PST

[LWMG95] Cary Laxer, Curt M. White, James E. Miller, and Judith L. Gersting, editors. Proceedings of the 26th SIGCSE Technical Symposium on Computer Science Education, 1995, Nashville, Tennessee, USA, March 2–4, 1995, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lidtke:1998:TCT

[LZ98] Doris K. Lidtke and Harry H. Zhou. A top-down, collaborative teaching approach of introductory courses in computer sciences (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):291, September 1998.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mader:1990:ASP

[MAC^{+}90] David Mader, E. Robert Anderson, Robert Cupper, James Leone, and Ralph Meeker. Accreditation and the small, private college (panel session): problems and opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):261, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Macpherson:1997:TSP

[Mac97] P. A. Macpherson. A technique for student program submission on UNIX systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):54–56, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1995:EIS

[MAD^{+}95] Renée A. McCauley, Clark Archer, Nell Dale, Rym Mili, James Robergé, and Harriet Taylor. The effective integration of software engineering principles throughout the undergraduate computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):364–365, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Magagnosc:1994:SCO

[Mag94a] David Magagnosc. Simulation in computer organization: a goals based study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):178–182, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Magno:1994:NAC

[Mag94b] Dominic Magno. A new approach to computer science in the community college: negotiated teaching and learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):47–51, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1994:OIT

[MAG^{+}94c] Renée A. McCauley, Evans J. Adams, Donald Gotterbarn, Linda M. Northrop, Hossein Saiedian, and Stuart Zweben. Organizational issues in teaching project-oriented software engineering courses (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):392–393, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Makinen:1996:PPC

[Mäk96] Erkki Mäkinen. Programming projects on chess. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):41–44, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Makkonen:1997:DCH

[Mak97] Pekka Makkonen. Does collaborative hypertext support better engagement in learning of the basics in informatics? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):130–132, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Makkonen:1998:WBP

[Mak98] Pekka Makkonen. WWW-based presentations as a complementary part of conventional lectures in the basic of informatics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):162–165, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mutchler:1999:CCL

[MAK^{+}99] David Mutchler, Claude Anderson, Andrew Kinley, Cary Laxer, Dale Oexmann, Rimli Sengupta, and Frank H. Young. CS 1 closed laboratories = multimedia materials + human interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):210, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mallozzi:1997:BTS

[Mal97] John S. Mallozzi. Binary trees á la STL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):159–163, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1995:SMS

[Mar95] Edward C. Martin. Solid modeling with scheme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):336–339, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1997:WCE

[Mar97] C. Dianne Martin. What is computer ethics? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):8–9, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marshall:1998:UIT

[Mar98a] David Marshall. Using the Internet to teach the Internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):171–174, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1998:DAC

[Mar98b] C. Dianne Martin. Deconstructing the ACM code of ethics and professional conduct. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):8–9, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1998:CSP

[Mar98c] C. Dianne Martin. Is computer science a profession? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):7–8, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marion:1999:CWS

[Mar99a] William Marion. CS1: what should we be teaching? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):35–38, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1999:ARAa

[Mar99b] C. Dianne Martin. From awareness to responsible action (part 1): defining learning objectives and necessary skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):8–10, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1999:ARAb

[Mar99c] C. Dianne Martin. From awareness to responsible action (part 2): developing a curriculum with progressive integration of ethics and social impact. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):10–12, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mattis:1995:AMC

[Mat95] W. E. Mattis. An advanced microprocessor course with a design component. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):60–64, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maxim:1993:PLC

[Max93] Bruce R. Maxim. Programming languages-comparatively speaking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):25–29, March 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mazaitis:1993:OOP

[Maz93] Dorothy Mazaitis. The object-oriented paradigm in the undergraduate curriculum: a survey of implementations and issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):58–64, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCloskey:1995:AAL

[MB95] Robert McCloskey and John Beidler. An analysis of algorithms laboratory utilizing the maximum segment sum problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):21–26, December 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McManus:1996:CUG

[MBC^{+}96] Margaret M. McManus, Catherine M. Beise, John A. Cross, John Saunders, and Michael B. Twidale. Challenges of using groupware to teach groupware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):165– 166, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Merritt:1993:AMH

[MBE^{+}93] Susan M. Merritt, Charles J. Bruen, J. Philip East, Darlene Grantham, Charles Rice, Viera K. Proulx, Gerry Segal, and Carol E. Wolf. ACM model high school computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):309, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maxim:1990:IPA

[MBJS90] Bruce R. Maxim, Gregory Bachelis, David James, and Quentin Stout. Introducing parallel algorithms in undergraduate computer science courses (tutorial session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):255, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1998:PFC

[MBW^{+}98] Daniel D. McCracken, Michael Berman, Ursula Wolz, Owen Astrachan, and Nell Dale. Possible futures for CS2 (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):357–358, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McMillan:1990:IAD

[MC90a] Thomas McMillan and William Collins. Implementing abstract data types in Turbo Pascal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):134– 138, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Micco:1990:LPD

[MC90b] Mary Micco and Phillip Cumpston. A large project for demonstrating knowledge engineering techniques including applications of neural networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):245–250, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCormick:1991:LTD

[McC91] John A. McCormick. A laboratory for teaching the development of real-time software systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):260–264, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1992:PLC

[McC92] Daniel D. McCracken. Programming languages in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):1–4, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McConnell:1996:ALU

[McC96] Jeffrey J. McConnell. Active learning and its use in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):52–54, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1997:CSE

[McC97] Renée A. McCauley. Computer science education links. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):19–20, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1998:GIR

[McC98a] Renée McCauley. General interest resources of use to computer science educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):20–21, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1998:WIA

[McC98b] Renée McCauley. Warning! instructional animation tools abound on the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):19–20, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1999:RTT

[McC99a] Renée McCauley. A road trip through online resources for introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (4):24–25, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1999:TAI

[McC99b] Renée McCauley. Teaching the artificial intelligence course: online resources that can help a lot! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2): 21–22, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1999:IAT

[McC99c] Daniel D. McCracken. An inductive approach to teaching object-oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):184–188, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:1991:NSL

[McD91] Chris McDonald. A network specification language and execution environment for undergraduate teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):25–34, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:1992:TCJ

[McD92] Chris McDonald. Teaching concurrency with Joyce and Linda. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):46–52, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCormick:1994:TAB

[MCD^{+}94a] John W. McCormick, Fintan Culwin, Nicholas J. DeLillo, Michael B. Feldman, Richard Pattis, and Walter J. Savitch. Teaching Ada by the book: the pedagogy of Ada in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):385–386, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:1994:FPI

[MCd^{+}94b] Philip Miller, Michael J. Clancy, Andrea A. diSessa, Jeremy Roschelle, Michael Eisenberg, Mark Guzdial, Elliot Soloway, and Mitchel Resnick. The future of programming instruction (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):400, March 1994. CODEN SIGSD3. ISBN 0-89791-646-8. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Compiler/common-lisp.bib;   ftp://ftp.
math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/1994.bib.

McDonald:1996:ULD

[McD96] Chris McDonald. User-level distributed file systems projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):333–337, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matocha:1998:EAA

[MCH98] Jeff Matocha, Tracy Camp, and Ralph Hooper. Extended analogy: an alternative lecture method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):262–266, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKenzie:1995:UCG

[McK95] Peter McKenzie. Using CASE to get to maintenance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):57–59, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McManus:1996:GTT

[McM96] Margaret M. McManus. Gathering together through groupware in an MA program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):136–138, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McConnell:1990:CGD

[MCMHS90] Jeffrey J. McConnell, Steve Cunningham, Barbara Mones-Hattal, and Deborah Sokolove. Computer graphics as a discipline (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):257, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McNally:1994:VSA

[McN94] Myles F. McNally. Visualization of sorting algorithms (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):408, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McNally:1996:UVT

[McN96] Myles F. McNally. Using visualization in teaching data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):231, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McNally:1998:UIP

[McN98] Myles F. McNally. Using image processing as a laboratory theme in CS1 and CS2 (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):292, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKinley:1999:MIG

[MCW99] P. K. McKinley, B. H. C. Cheng, and J. J. Weng. Moving industry-guided multimedia technology into the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):160–164, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McFall:1994:FSM

[MD94] Ryan McFall and Herbert L. Dershem. Finite state machine simulation in an introductory lab. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):126–130, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morsiani:1999:LOS

[MD99] Mauro Morsiani and Renzo Davoli. Learning operating systems structure and implementation through the MPS computer system simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):63–67, March 1999.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mulder:1995:PCI

[MDG^{+}95] Michael C. Mulder, Gordon Davis, John Gorgone, David Feinstein, and Doris K. Lidtke. A proposed curriculum in information science (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):380– 381, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mayerhofer:1992:DAS

[MdL92] Mário André Mayerhofer and Carlos José Pereira de Lucena. Design of an algorithm simulation and animation environment (ASA). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):7–14, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Magee:1992:ICD

[ME92] Michael Magee and Sue Englert. An interdisciplinary course in digital image processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):197– 201, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medley:1998:LFC

[Med98] M. Dee Medley. On-line finals for CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):178–180, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meredith:1990:ELP

[Mer90] Marsha Meredith. An effective Lisp project for a programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):45–48, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meredith:1992:IPC

[Mer92] Marsha J. Meredith. Introducing parallel computing into the undergraduate computer science curriculum: a progress report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):187–191, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Messersmith:1991:SCU

[Mes91] Kenneth G. Messersmith. Student created user manuals for a course on programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (4):53–54, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meyer:1996:CLG

[Mey96] R. Mark Meyer. CANLOGS: a logic gate simulator for all seasons. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):58–62, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1998:CAC

[MF98] Daniel D. McCracken and Dennis J. Frailey. A conversation about computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2): 36–39, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mansouri:1998:PPA

[MGH98] Fatima Z. Mansouri, Cleveland A. Gibbon, and Colin A. Higgins. PRAM: Prolog automatic marker. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):166–170, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mims:1991:UTL

[MH91] Ted Mims and Andrzej Hoppe. Utilizing a transputer laboratory and Occam2 in an undergraduate operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):317–323, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marshall:1996:IHC

[MH96a] A. D. Marshall and S. Hurley. Interactive hypermedia courseware for the World Wide Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):1–5, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927

(electronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/MPG/sigcse99.bib.

McGill:1996:SPD

[MH96b] Tanya McGill and Valerie Hobbs. A supplementary package for distance education students studying introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):73–77, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McLoughlin:1996:TFP

[MH96c] Henry McLoughlin and Kevin Hely. Teaching formal programming to first year computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):155–159, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mead:1998:EYE

[MHB^{+}98] Nancy R. Mead, Thomas B. Hilburn, Donald J. Bagert, Michael Ryan, Oddur Benediktsson, and Frances Van Scoy. Everything you ever wanted to know about software engineering education, but were afraid to ask (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):260–263, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murphy:1998:CCS

[MHN98] Mike Murphy and Richard Halstead-Nussloch. A challenge for computer science educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):41, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mancoridis:1993:CCE

[MHP93] Spiros Mancoridis, Richard C. Holt, and David A. Penny. A “curriculum-cycle” environment for teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):15–19, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mercuri:1998:UHJ

[MHP98] Rebecca Mercuri, Nira Herrmann, and Jeffrey Popyack. Using HTML and JavaScript in introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):176–180, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moses:1994:TFW

[MIG^{+}94] Louise E. Moses, Rachelle D. Isles, Frances Grundy, Danielle R. Bernstein, Valerie A. Clarke, and G. Joy Teague. Too few women! too few minorities! what can we do? (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):367–368, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:1993:MDS

[Mil93] Diane M. Miller. MIS as a discipline: a structured definition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):27–32, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:1996:MAI

[Mil96] James E. Miller. Mistakes to avoid in the introduction of compressed video. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):150–152, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitzenmacher:1996:DSP

[Mit96] Michael Mitzenmacher. Designing stimulating programming assignments for an algorithms course: a collection of exercises based on random graphs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):29–36, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitrovic:1998:LSC

[Mit98] Antonija Mitrovic. Learning SQL with a computerized tutor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):307–311, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1999:TSE

[MJ99] Renée McCauley and Ursula Jackson. Teaching software engineering early: experiences and results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2): 86–91, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1996:DSU

[MJM96] Renée A. McCauley, Ursula Jackson, and Bill Manaris. Documentation standards in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):242–246, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:1997:IPT

[MK97] Chris McDonald and Kamran Kazemi. Improving the PVM teaching environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):219–223, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mayo:1998:SNL

[MK98] Jean Mayo and Phil Kearns. A secure networked laboratory for kernel programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):175–177, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mayo:1999:SUA

[MK99] Jean Mayo and Phil Kearns. A secure unrestricted advanced systems laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):165–169, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mynatt:1990:ECB

[ML90] Barbee T. Mynatt and Laura Marie Leventhal. An evaluation of a CASE-based approach to teaching undergraduate software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):48–52, February 1990.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mutchler:1996:UMG

[ML96] David Mutchler and Cary Laxer. Using multimedia and GUI programming in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):63–65, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mulder:1997:EEE

[MLS97] Michael C. Mulder, Doris Lidtke, and Gordon E. Stokes. Enterprise enhanced education: an information technology enabled extension of traditional learning environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):355–359, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:1993:DOC

[MM93] Gary McDonald and Merry McDonald. Developing oral communication skills of computer science undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):279–282, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meter:1994:EST

[MM94a] Glenn Meter and Philip Miller. Engaging students and teaching modern concepts: Literate, situated, object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):329–333, March 1994. CO-DEN SIGSD3. ISBN 0-89791-646-8. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/1994.
bib.

Mili:1994:TFC

[MM94b] Rym Mili and Ali Mili. Teaching a first course on data structures: a software engineering approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):21–25, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1997:RAS

[MM97] Renée A. McCauley and Bill Z. Manaris. Report on the annual survey of departments offering CSAC/CSAB-accredited computer science degree programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):301–305, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:1998:CSD

[MM98] Renée McCauley and Bill Manaris. Computer science degree programs: what do they look like? A report on the annual survey of accredited programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):15–19, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:1999:CSC

[MM99] Merry McDonald and Gary McDonald. Computer science curriculum assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):194–197, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKim:1991:ACE

[MMB^{+}91] James C. McKim, Jr., Timothy O. Martyn, Roger H. Brown, Michael M. Danchak, Kathleen L. Farrell, C. William Higginbotham, Irina S. Ilovic, Brian J. McCartin, and J. Peter Matelski. An alternative culminating experience for Master’s students in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):210– 218, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meyerowitz:1998:SDM

[MMK98] Jane Meyerowitz, Maurice Mars, and Steve Knight. Student development of multimedia CBE (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):294, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Macek:1999:GCP

[MMKW99] Tomas Macek, Bozena Mannova, Josef Kolar, and Barbara Williams. Global cooperation project in computer programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):208–211, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Merritt:1990:INS

[MN90] Susan M. Merritt and Cecilia Y. Nauck. Inventing a new sorting algorithm: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):181–185, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morton:1993:SPL

[MN93] Larry Morton and Nicholas Norgaard. A survey of programming languages in CS programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2): 9–11, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1999:MDD

[MNC99] Michael McCracken, Wendy Newstetter, and Jeff Chastine. Misconceptions of designing: a descriptive study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):48–51, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marsden:1996:MMB

[MO96] Paul Marsden and Mark O’Connell. MuPMoTT: a multimedia based tool supporting the teaching of process modelling within a framework of structured system analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):116–118, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mody:1991:CES

[Mod91] R. P. Mody. C in education and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):45–56, September 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/leavens.bib.

Moen:1995:LTI

[Moe95] Selmer Moen. A low-tech introduction to operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):149–153, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mooney:1992:CSP

[Moo92] James D. Mooney. A course in software portability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):53–56, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moore:1993:SIB

[Moo93] Thomas K. Moore. Scientific investigation in a breadth-first approach to introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):63–67, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moore:1995:AUH

[Moo95] Thomas Moore. Active use of hypertext to aid learning and classroom instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):297–301, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morris:1992:EIC

[Mor92] J. M. Morris. The effects of an introductory computer course on the attitudes of older adults towards computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):72–75, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morris:1995:EML

[Mor95] J. Morgan Morris. Experiences with mosaic for legacy projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):292–296, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moses:1993:OCS

[Mos93] L. E. Moses. Our computer science classrooms: are they “friendly” to female students? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3):3–12, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moser:1997:FAG

[Mos97] Robert Moser. A fantasy adventure game as a learning environment: why learning to program is so difficult and what can be done about it. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):114–116, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moses:1998:ACE

[Mos98] L. E. Moses. ART199/CS199 the electronic medium (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):296, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Motil:1991:BBA

[Mot91] John Motil. Begin-BIG an approach to the introductory computing course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):226–230, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Motil:1996:DTS

[Mot96] John Motil. Desktop teaching: on the students desk. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):231, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miranda:1996:UIT

[MP96] José Eduardo Pina Miranda and Jorge Sousa Pinto. Using Internet technology for course support. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):96–100, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marti:1996:PTN

[MPH96] Willis F. Marti, Udo W. Pooch, and John A. Hamilton, Jr. PACKET tracing: a new paradigm for teaching computer network courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):162–164, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mulder:1995:CLU

[MPHL95] Michael C. Mulder, Jane Chu Prey, J. E. Haines, and Doris K. Lidtke. Collaborative learning in undergraduate information science education (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):400– 401, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mazlack:1992:DEE

[MPTW92] Lawrence J. Mazlack, Roger A. Pick, Paul Tudor, and Wallace R. Wood. Developing expertise in expert system development by developing prototypes for actual commercial applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):81–86, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miao:1999:CMI

[MPW99] Yongwu Miao, Hans-Rüdiger Pfister, and Martin Wessner. Combining the metaphors of an institute and of networked computers for building collaborative learning environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):188, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medley:1998:GWG

[MR98] M. Dee Medley and Rebecca H. Rutherfoord. Group 2 (working group): ethical issues related to Internet development and research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):269, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:1999:NSF

[MR99] C. Dianne Martin and Margaret M. Reek. National Science Foundation Course, Curriculum and Laboratory Improvement program: adaptation and implementation track. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):366, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medley:1998:EIR

[MRA^{+}98] M. Dee Medley, Rebecca H. Rutherfoord, G. Ernest Anderson, R. Waldo Roth, and Stuart A. Varden. Ethical issues related to Internet development and research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4): 61–76, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mueller:1993:IFY

[MRS93] Conrad Mueller, Sheila Rock, and Ian Sanders. An improved first year course taking into account third world students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):213–217, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murthy:1990:MPA

[MS90] Narayan Murthy and Allen Stix. Multiple precision arithmetic: a programming assignment in CS2 applying linked lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):129–133, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McFall:1997:ICS

[MS97] Ryan McFall and Gordon Stegink. Introductory computer science for general education: laboratories, textbooks, and the internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):96–100, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moore:1998:MCP

[MS98] Simon Moore and Martyn Spink. A multimedia C programming course that supports different learning situations (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):295, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin-Sanchez:1995:GIA

[MSPF95] Óscar Martín-Sánchez and Cristóbal Pareja-Flores. A gentle introduction to algorithm complexity for CS1 with nine variations on a theme by Fibonacci. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):49–56, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medley:1995:EPI

[MSRT95] Mary Dee Medley, Kay G. Schulze, Bob Riser, and Rebekah L. Tidwell. Ethical and professional issues in computing (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):366–367, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:1997:ECA

[MSSZ97] Philip Miller, Kris Stewart, Klaus Sutner, and Josepyh L. Zachary. Exploiting computer algebra systems in computer science courses (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):377–378, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meisalo:1997:CTD

[MST97] V. Meisalo, E. Sutinen, and J. Tarhio. CLAP: teaching data structures in a creative way. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 117–119, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mukundan:1999:TCG

[Muk99] R. Mukundan. Teaching computer graphics using Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):66–69, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mullins:1991:PIP

[Mul91] Paul M. Mullins. Pascal with Internet primitives for pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):51–54, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murphy:1991:EAU

[Mur91] Michael G. Murphy. Evolution of an approach to undergraduate ADA and software engineering instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):55–57, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGill:1995:IRB

[MV95] Tanya McGill and Simone Volet. An investigation of the relationship between student algorithm quality and program quality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):44–48, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meyer:1998:SFC

[MV98] Jeanine Meyer and Stuart A. Varden. A survey of first computing course students (poster): new findings and their implications for the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):293, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marion:1998:IRR

[MVS^{+}98] Bill Marion, Keith Vander Linden, Roberta Sabin, Judy Cushing, and Penny Anderson. Integrating recent research results into undergraduate curricula (panel): initial steps. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):378–379, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mason:1998:ITC

[MW98a] David V. Mason and Denise M. Woit. Integrating technology into computer science examinations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):140–144, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:1998:SAH

[MW98b] J. Paul Myers, Jr. and Henry M. Walker. The state of academic hiring in computer science: and interim review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):32–36, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Madej:1999:IIS

[MW99a] Jan Madej and Tadeusz Wilusz. IntraED — an intranet solution for EDucation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):203, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mason:1999:PMF

[MW99b] David V. Mason and Denise M. Woit. Providing mark-up and feedback to students with online marking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):3–6, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:1999:WWO

[MW99c] Michael McCracken and Robert Waters. Why? when an otherwise successful intervention fails. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):9–12, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maly:1996:IRI

[MWO^{+}96] K. Maly, C. Wild, C. M. Overstreet, H. Abdel-Wahab, A. Gupta, A. Youssef, E. Stoica, R. Talla, and A. Prabhu. Interactive remote instruction: initial experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):153–155, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moore:1994:ASC

[MWP94] Thomas Moore, Michael Wick, and Blaine Peden. Assessing student’s critical thinking skills and attitudes toward computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):263–267, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mengel:1999:CSS

[MY99] Susan A. Mengel and Vinay Yerramilli. A case study of the static analysis of the quality of novice student programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):78–82, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:1990:CRM

[Mye90] J. Paul Myers, Jr. The central role of mathematical logic in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):22–26, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:1992:MSW

[Mye92] J. Paul Myers, Jr. Men supporting women computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):63–66, March 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:1994:IUD

[Mye94] William Myers. Instructional uses of demonstration disks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):76–79, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:1998:EMH

[Mye98] J. Paul Myers, Jr. The eight-minute halting problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):53–56, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marchioro:1995:CSP

[MZS^{+}95] Thomas L. Marchioro II, Joseph Zachary, D. E. Stevenson, Ignatios Vakalis, and Leon Tabak. Computational science, parallel and high performance computing in undergraduate education (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):362–363, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nachbar:1998:BRW

[Nac98] Daniel Nachbar. Bringing real-world software development into the classroom: a proposed role for public software in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):171–175, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nadi:1996:SPS

[Nad96] Najib Nadi. Simulation programs for studying computer organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):233, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nakhimovsky:1998:IBS

[Nak98] Alexander Nakhimovsky. An intranet-based software engineering course (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):297, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1990:AVC

[Nap90] Thomas L. Naps. Algorithm visualization in computer science laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):105–110, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1996:AVS

[Nap96] Thomas L. Naps. Algorithm visualization served off the World Wide Web: why and how. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):66–71, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927

(electronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/Pape.bib.

Naps:1997:UWD

[Nap97a] Thomas Naps. Using the WWW as the delivery mechanism for interactive, visualization-based instructional modules (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):143, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1997:AVW

[Nap97b] Thomas L. Naps. Algorithm visualization on the World Wide Web — the difference Java makes! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (3):59–61, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1998:JVC

[Nap98] Thomas L. Naps. A Java visualiser class: incorporating algorithm visualisations into students’ programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):181–184, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Narayanan:1991:PME

[Nar91] V. Narayanan. Pattern matching for everyone. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):27–30, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Navrat:1994:HPC

[Náv94] Pavol Návrat. Hierarchies of programming concepts: abstraction, generality, and beyond. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):17–21, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1998:MWE

[NB98] Thomas L. Naps and Eric Bressler. A multi-windowed environment for simultaneous visualization of related algorithms on the world wide Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):277–281, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Navrat:1999:SEE

[NB99] Pavol Návrat and Mária Bieliková. Software engineering education: different contexts, similar contents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):55–59, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1999:UVT

[NC99a] Thomas L. Naps and Eric E. Chan. Using visualization to teach parallel algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):232–236, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nielsen:1999:CTC

[NC99b] S. Travis Nielsen and Douglas M. Campbell. Current trends in computer science graduate admissions: a survey of the top 108 programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):31–34, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neff:1993:LPE

[Nef93] Norman Neff. A logic programming environment for teaching mathematical concepts of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):20–24, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neff:1999:ODC

[Nef99] Norman Neff. OO design in compiling an OO language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):326–330, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nevison:1994:TPC

[Nev94] Chris Nevison. Teaching parallel computing on a messagepassing architecture (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 399, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nevison:1999:SSC

[Nev99] Christopher H. Nevison. Seminar: safe concurrent programming in Java with CSP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):367, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newmarch:1993:PBA

[New93] Jan Newmarch. A plan-based approach to Prolog recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):12–18, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nebash:1997:UHL

[NF97] Bohdan Nebash and Michael B. Feldman. Using HTML linking to help novice programmers to reuse components. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):315–319, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Najera:1994:ITC

[NG94] Jerónimo Nájera and Luis Gómez. An instructed tool for cache coherence protocols. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):61–64, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:1998:DPD

[Ngu98] Dung Nguyen. Design patterns for data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):336–340, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neal:1998:IPU

[NI98] L. R. Neal and A. D. Irons. Integrating professionalism into undergraduate degree courses in computing (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):264–267, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nevison:1995:CAP

[NKN^{+}95] Christopher H. Nevison, Joe Kmoch, Robert E. Noonan, Tim Corica, Sarah Fix, and David G. Kay. Changes in the advanced placement computer science course: case studies and C++ (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):374–375, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nohl:1994:UPA

[Noh94] Daniel E. Nohl. Using profiling to analyze algorithms (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):409, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noyes:1993:TAB

[Noy93] James L. Noyes. Teaching AI: a breadth-first approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):33–37, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noyes:1995:TMN

[Noy95] James L. Noyes. Teaching a modern numerical analysis course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):355–359, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newsome:1992:GCS

[NP92] Mark Newsome and Cherri M. Pancake. A graphical computer simulator for systems programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):157–162, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nute:1993:RPA

[NPP^{+}93] Tom Nute, Lavon Page, Jo Perry, Richard Rinewalt, Pat Ryan, Laurie White, and Stan Wileman. Results and problems from the 1992 ACM scholastic programming contest finals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):43–50, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nute:1990:AES

[NRR^{+}90] T. C. Nute, J. R. Rienwalt, D. J. Rodjak, William H. Golaz, and J. R. Comer. Academic experiences in software project management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):47–53, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1994:OOA

[NS94] Thomas L. Naps and Brian Swander. An object-oriented approach to algorithm visualization-easy, extensible, and dynamic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):46–50, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1996:TVE

[NS96] Thomas L. Naps and Jeremy Stenglein. Tools for visual exploration of scope and parameter passing in a programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):305–309, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Norris:1998:ETY

[NSS98] Cathleen Norris, Jennifer Smolka, and Elliot Soloway. Educational technology over 25 years (panel): understanding the conditions when it works. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):268, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nishida:1996:LSS

[NSTT96] Tomohiro Nishida, Akinori Saitoh, Yoshihiro Tsujino, and Nobuki Tokura. Lecture supporting system by using e-mail and WWW. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):280–284, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Null:1996:ATC

[Nul96] Linda Null. Applying TQM in the computer science classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):120–124, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nulden:1998:EPA

[Nul98] Urban Nuldén. The ExCon project: advocating continuous examination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):126–130, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nerheim-Wolfe:1992:PLI

[NW92a] Rosalee Nerheim-Wolfe. Providing a laboratory for instruction set design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):163–167, March 1992.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neumann:1992:MDA

[NW92b] William T. Neumann and Marvin C. Woodfill. A multidisciplinary approach for digital systems design curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):168–172, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:1999:PDD

[NW99a] Dung (‘Zung’) Nguyen and Stephen B. Wong. Patterns for decoupling data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):87–91, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Norris:1999:CSC

[NW99b] Cindy Norris and James Wilkes. Computer systems “conference” for teaching communication skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):189–193, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:1995:NTS

[OBH^{+}95] Barbara Boucher Owens, Gene Bailey, Shelly Heller, Ted Mims, and Laurie White. The non-traditional student in computing: characteristics, needs and experiences (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):372–373, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:1997:CCS

[OBPB97] Barbara Boucher Owens, Shirley Booth, Marian Petre, and Anders Berglund. Changing computer science curricula (panel): planning for the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):139–140, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:1994:NMC

[OCH^{+}94] Barbara Boucher Owens, Robert D. Cupper, Stuart Hirshfield, Walter Potter, and Richard Salter. New models for the CS1 course: what are they and are they leading to the same place? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):364–365, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliver:1994:SDP

[OD94] S. Ron Oliver and John Dalbey. A software development process laboratory for CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):169–173, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ODwyer:1998:ESE

[O’D98] Tom O’Dwyer. Educational software — a European perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):10–13, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

OGorman:1998:TOS

[O’G98] John O’Gorman. Teaching operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):61–63, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osborne:1993:OUC

[OJ93] Martin Osborne and James L. Johnson. An only undergraduate course in object-oriented technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (1):101–106, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ONeal:1995:WMS

[OK95] Michael B. O’Neal and Barry L. Kurtz. Watson: a modular software environment for introductory computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):87–91, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olagunju:1991:PMC

[Ola91] Amos O. Olagunju. The plight of a minority in computer science: an educational manifesto. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):219–225, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliver:1994:SEP

[Oli94] Dave Oliver. Software engineering project work in combined distance and on campus modes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2): 31–35, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olszewski:1993:CL

[Ols93] Jacek Olszewski. CSP laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1): 91–95, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olsson:1995:TTC

[OM95] Ronald A. Olsson and Carole M. McNamee. Tools for teaching CCRs, monitors, and CSP concurrent programming concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):31–40, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oh:1999:TRT

[OM99] Jae C. Oh and Daniel Mossé. Teaching real time OSs with DORITOS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):68–72, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Omland:1999:ESA

[Oml99] Hans Olav Omland. Educating systems analysts emphasising the human factor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):44–47, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ONeill:1995:CPM

[O’N95] Edward J. O’Neill. Classroom presentation of memory usage with constants, variables, and subprograms in Pascal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):18–24, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ODubhchair:1998:CCI

[OQ98] Kate O’Dubhchair and Sarah Quilty. The changing classroom-ICTs in 21st century education (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):298, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osborne:1991:OSG

[Osb91] Martin Osborne. OZ: a syntax generating editor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):51–52, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osborne:1992:TCU

[Osb92a] Lawrence Osborne. Teaching C with UNIX for college credit to professional programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):43–48, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osborne:1992:ATT

[Osb92b] M. Osborne. APPGEN: a tool for teaching systems analysis and design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):259–263, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osborne:1995:SEC

[Osb95] Martin Osborne. Software engineering, C++, and windows. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):243–247, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owen:1990:ETA

[Owe90] G. Scott Owen. Experiences in teaching an advanced computer graphics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):162–166, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owen:1992:TCG

[Owe92] G. Scott Owen. Teaching computer graphics using Render-Man. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):304–308, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owen:1996:IWW

[Owe96]  G. Scott Owen.  Integrating World Wide Web tech-

nology into undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):101–103, ???? 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Compiler/java.bib;   ftp://ftp.math.utah.
edu/pub/mirrors/ftp.ira.uka.de/bibliography/Compiler/
JAVA/java.bib.

Penny:1990:LST

[PA90] J. P. Penny and P. J. Ashton. Laboratory-style teaching of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):192–196, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:1999:NAP

[PAC^{+}99] Nick Parlante, Owen Astrachan, Mike Clancy, Richard E. Pattis, Julie Zelenski, and Stuart Reges. Nifty assignments panel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):354–355, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pandey:1990:GLP

[Pan90] Rajeev Pandey. Getting the languages for a programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):11–14, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parker:1995:SDC

[Par95] David L. Parker. Structured design for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):258–262, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:1997:TOO

[Par97] Nick Parlante. Teaching with object oriented libraries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):140–144, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattis:1990:PEC

[Pat90] Richard E. Pattis. A philosophy and example of CS-1 programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):34–39, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattis:1993:PEA

[Pat93] Richard E. Pattis. The “procedures early” approach in CS 1: a heresy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):122–126, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattis:1994:TEF

[Pat94] Richard E. Pattis. Teaching EBNF first in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):300–303, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattinson:1996:NMT

[Pat96] Colin Pattinson. Network management teaching-managing a simulated network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):233, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattis:1997:TOC

[Pat97] Richard E. Pattis. Teaching OOP in C++ using an artificial life framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):39–43, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paxton:1995:NAT

[Pax95] John T. Paxton. A novel approach to teaching artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):283–286, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paxton:1996:WUW

[Pax96] John T. Paxton. Webucation: using the Web as a classroom tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):285–289, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Persky:1998:REC

[PBA98] Yakov Persky and Mordechai Ben-Ari. Re-engineering a concurrency simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):185–188, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perl:1993:ATF

[PBEC93] Teri Perl, Dennis Bybee, Carol E. Edwards, and Coco Conn. ACM task force report on K-12 education and technology (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):289, March 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pilgrim:1998:SLE

[PC98] C. J. Pilgrim and M. J. Creek. The Swinburne On-Line Education Project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):189–192, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parrish:1997:ACU

[PCLM97] Allen Parrish, David Cordes, Cynthia Lester, and Deanne Moore. Assessing computer usage patterns in a software development course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):58–62, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poucher:1990:ASP

[PCRR90] William B. Poucher, James Comer, Richard Rinewalt, and Patrick Ryan. The ACM scholastic programming contest-1977 to 1990 (special panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 256, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perez-Davila:1995:BBA

[PD95] Alfredo Pérez-Dávila. O.S. bridge between academia and reality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):146–148, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parker:1996:SBS

[PD96] Brenda C. Parker and Peter G. Drexel. A system-based sequence of closed labs for computer systems organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):53–57, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perry:1999:HF

[Per99] Nigel Perry. The humble fraction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (4):61–65, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pesante:1991:IWC

[Pes91] Linda H. Pesante. Integrating writing into computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):205–209, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petre:1998:AITa

[Pet98a] Marian Petre. Assessing innovation in teaching: an example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):40–42, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petre:1998:AITb

[Pet98b] Marian Petre. Assessing innovation in teaching: an example — part 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):28–31, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pfeiffer:1999:WEW

[Pfe99] Phil Pfeiffer. What employers want from students: a report from OOPSLA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):69–70, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:1994:UGP

[PFR^{+}94] Viera Proulx, Harriet Fell, Richard Rasala, Johannes A. G. M. Koomen, and Carol W. Wilson. Using graphical presentation techniques in closed laboratory exercises (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):366, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Popyack:1993:MMF

[PH93] Jeffrey L. Popyack and Nira Herrmann. Mail merge as a first programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):136–140, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pierce:1995:CRJ

[PH95] Margaret Anne Pierce and John W. Henry. Computer-related judgements of computer professionals and students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):106–110, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pierce:1996:CES

[PH96] Margaret Anne Pierce and John W. Henry. Computer ethics and social issues case analysis: an environmental algorithm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):50–54, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pfeiffer:1997:MSP

[PH97] Phil Pfeiffer and Matt Heintzelman. Machines, statues, and people: strategies for promoting RSI awareness in computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):296–300, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parker:1999:CBI

[PH99] Helen Parker and Mike Holcombe. Campus-based industrial software projects: risks and rewards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):189, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pigford:1992:DET

[Pig92] D. V. Pigford. The documentation and evaluation of teamoriented database projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):28–33, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pilgrim:1993:DCV

[Pil93] Robert A. Pilgrim. Design and construction of the very simple computer (VSC): a laboratory project for an undergraduate computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):151– 154, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pilgrim:1995:TTT

[Pil95] Robert A. Pilgrim. TIC-TAC-TOE: introducing expert systems to middle school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):340– 344, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Piotrowski:1992:TIE

[Pio92] Jerzy A. Piotrowski. Translation — an introductory exercise. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):20–28, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pilgrim:1996:AUI

[PL96] C. J. Pilgrim and Y. K. Leung. Appropriate use of the Internet in computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):81–86, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pilgrim:1997:CEM

[PLG97] C. J. Pilgrim, Y. K. Leung, and D. D. Grant. Cost effective multimedia courseware development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (3):45–50, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plimmer:1998:MIW

[Pli98] Beryl Plimmer. Machines invented for WW II code breaking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):37–40, December 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pargas:1997:TPC

[PLU97] Roy P. Pargas, Joe C. Lundy, and John N. Underwood. Tournament play in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):214–218, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parker:1995:GOA

[PM95] Brenda C. Parker and John D. McGregor. A goal-oriented approach to laboratory development and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):92–96, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poole:1996:ISG

[PM96] Bernard John Poole and Timothy S. Meyer. Implementing a set of guidelines for CS majors in the production of program code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):43–48, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Palme:1997:UCC

[PM97] Jacob Palme and Sirkku Männikö. Use of computer conferencing to teach a course on humans and computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):88–90, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Palakal:1998:ILE

[PMB98] Mathew J. Palakal, Frederick W. Myers, and Carla L. Boyd. An interactive learning environment for breadth-first computing science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):1–5, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Patino-Martinez:1997:AAC

[PMCGJP97] Marta Patiño-Martínez, J. Ignacio Castelló-Gómez, and Ricardo Jiménez-Peris. AnLex and AnSin: a compiler generator system for beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):248–252, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prey:1999:PST

[PN99] Jane Prey and Robert E. Noonan, editors. Proceedings of the 30th SIGCSE Technical Symposium on Computer Science Education, 1999, New Orleans, Louisiana, USA, March 14–28, 1999, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1999. CODEN SIGSD3. ISBN 1-58113-085-6. ISSN 0097-8418 (print), 2331-3927 (electronic).

Polak:1999:TCC

[Pol99] Przemyslaw Polak. Teaching computer communication skills using case study method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):190, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poole:1991:USE

[Poo91] Bernard John Poole. Undergraduate software engineering: the fruits of experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):15–20, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pournaghshband:1990:SPC

[Pou90] Hassan Pournaghshband. The students’ problems in courses with team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):44–47, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Power:1999:DLD

[Pow99] Chris Power. Designer — a logic diagram design tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):211, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Palmer:1995:ATT

[PP95] Thomas V. Palmer and James C. Pleasant. Attitudes toward the teaching of formal methods of software development in the undergraduate computer science curriculum: a survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):53–59, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Price:1997:TPT

[PP97] Blaine Price and Marian Petre. Teaching programming through paperless assignments: an empirical evaluation of instructor feedback. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):94–99, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plouzeau:1992:ECD

[PR92] N. Plouzeau and M. Raynal. Elements for a course on the design of distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):35–40, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pierson:1998:WBA

[PR98a] Willard C. Pierson and Susan H. Rodger. Web-based animation of data structures using JAWAA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):267–271, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prim:1998:SES

[PR98b] Marta Prim and Jordi Roig. Self-evaluation system for digital systems subject (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):299, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pratt:1990:UCA

[Pra90] Terrence W. Pratt. Upgrading CS1: an alternative to the proposed COCS survey course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):68–71, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prather:1992:CSU

[Pra92] Ronald E. Prather. Computer science in an undergraduate liberal arts and sciences setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):59–64, June 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prechelt:1994:MEP

[Pre94] Lutz Prechelt. A motivating example problem for teaching adaptive systems design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):25–34, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prey:1996:CLC

[Pre96] Jane Chu Prey. Cooperative learning and closed laboratories in an undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):23–24, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Preiss:1999:DPD

[Pre99] Bruno R. Preiss. Design patterns for the data structures and algorithms course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):95–99, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:1996:FCS

[PRF96] Viera K. Proulx, Richard Rasala, and Harriet Fell. Foundations of computer science: what are they and how do we teach them? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):42–48, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:1997:RGC

[Pro97] Viera K. Proulx. Recursion and grammars for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):74–76, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:1998:TSC

[Pro98] Viera K. Proulx. Traffic simulation: a case study for teaching object oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):48–52, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:1999:HER

[Pro99] Viera K. Proulx. Hospital emergency room simulation: object oriented design issues for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):92–94, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paxton:1993:IBF

[PRS93] John Paxton, Rockford J. Ross, and Denbigh Starkey. An integrated, breadth-first computer science curriculum based on computing curricula 1991. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):68–72, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paxton:1994:MTI

[PRS94] John T. Paxton, Rockford J. Ross, and J. Denbigh Starkey. A methodology for teaching an integrated computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):1–5, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pruhs:1993:SP

[Pru93] Kirk Pruhs. The SPIN-OUT puzzle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (3):36–38, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pyott:1991:AAT

[PS91] Sean Pyott and Ian Sanders. ALEX: an aid to teaching algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):36–44, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petre:1997:HTE

[PS97] Marian Petre and Douglas Siviter. Harnessing technology for effective inter/intra-institutional collaboration (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):145, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Preston:1998:SID

[PS98] Jon Anderson Preston and Russell Shackelford. A system for improving distance and large-scale classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):193–198, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Preston:1999:ILA

[PS99] Jon A. Preston and Russell Shackelford. Improving on-line assessment: an investigation of existing marking methodologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):29–32, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Put:1999:SCS

[PSZ99a] Dariusz Put, Janusz Stal, and Marek Zurowicz. The system of checking students’ knowledge with the use of wide area networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):192, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Put:1999:UIT

[PSZ99b] Dariusz Put, Janusz Stal, and Marek Zurowicz. The use of the Internet in the teaching process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):193, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Polack-Wahl:1999:ICR

[PW99] Jennifer A. Polack-Wahl. Incorporating the client’s role in a software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):73–77, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prince:1994:GGT

[PWST94] Charles Prince, Roger L. Wainwright, Dale A. Schoenefeld, and Travis Tull. GATutor: a graphical tutorial system for genetic algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):203–207, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paprzycki:1996:SFU

[PZ96] Marcin Paprzycki and Janusz Zalewski. Shaping the focus of the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):37–44, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rabung:1994:ICC

[Rab94] John Rabung. Introducing computer concepts to novices by “practical” immersion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):88–91, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radensky:1990:CAU

[Rad90] Atanas Radensky. Can Ada be used as a primary programming language?: major problems and their solutions by means of subsets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):201–205, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:1993:VO

[Rad93] Atanas Radenski. A voyage to oberon. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(3): 13–18, September 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:1997:DIC

[Ras97a] Richard Rasala. Design issues in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):4–7, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:1997:FOF

[Ras97b] Richard Rasala. Function objects, function templates, and passage by behavior in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):35–38, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:1997:MCT

[Ras97c] Richard Rasala. A model C++ tree iterator class for binary search trees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):72–76, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:1999:AAA

[Ras99] Richard Rasala. Automatic array algorithm animation in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):257–260, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosel:1992:IFS

[RB92] Andreas Rösel and Paul Bailes. Identifying foundations of software engineering expertise. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):52–64, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:1997:WUI

[RBG^{+}97] Rockford J. Ross, Christopher M. Boroni, Frances W. Goosey, Michael Grinder, and Paul Wissenbach. WebLab! A universal and interactive teaching, learning, and laboratory environment for the World Wide Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):199–203, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:1997:CTM

[RBL^{+}97] Susan H. Rodger, Anna O. Bilska, Kenneth H. Leider, Magdalena Procopiuc, Octavian Procopiuc, Jason R. Salemme, and Edwin Tsang. A collection of tools for making automata theory and formal languages come alive. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 15–19, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberge:1997:BCS

[RC97] James Robergé and C. R. Carlson. Broadening the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):320–324, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Resler:1998:VVT

[RD98] R. Daniel Resler and Dean M. Deaver. VCOCO: a visualisation tool for teaching compilers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):199–202, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rebelsky:1998:DBC

[Reb98] Samuel A. Rebelsky. Designing and building course webs with site-level authoring tools (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 384, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:1990:UOS

[Ree90] Margaret M. Reek. An undergraduate operating systems lab course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):171–175, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:1995:TAT

[Ree95] Margaret M. Reek. A top-down approach to teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):6–9, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:1996:SIS

[Ree96a] Kenneth A. Reek. A software infrastructure to support introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):125–129, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:1996:TIV

[Ree96b] Kenneth A. Reek. Teaching inheritance versus inclusion to first year computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1): 24–26, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reeves:1997:CMT

[Ree97] Douglas S. Reeves. A course on multimedia technology for computer science and computer engineering students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):174–178, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:1998:IPS

[Ree98] David Reed. Incorporating problem-solving patterns in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):6–9, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:1990:TIC

[Rei90] Richard J. Reid. A toolkit for individualized compiler-writing projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):81–85, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:1991:OOP

[Rei91] Richard J. Reid. Object-oriented programming in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):9–14, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:1992:LBC

[Rei92] Richard J. Reid. A laboratory for building computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):192–196, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:1993:OOP

[Rei93] Richard J. Reid. The object oriented paradigm in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):265–269, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:1995:OOS

[Rei95a] Richard J. Reid. Object-oriented simulation of computer architectures using C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):19–23, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reinfelds:1995:TPF

[Rei95b] Juris Reinfelds. A three paradigm first course for CS majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):223–227, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Resler:1996:PDT

[Res96] Dan Resler. The Prisoner’s Dilemma Tournament revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):31–36, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Revesz:1998:ETR

[Rev98] Gyorgy E. Revesz. Experimenting with tail-recursion and simulating coroutines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):51–54, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reynolds:1996:RCS

[RF96] Charles Reynolds and Christopher Fox. Requirements for a computer science curriculum emphasizing information technology: subject area curriculum issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):247–251, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Riehl:1993:USI

[RFH^{+}93] Arthur M. Riehl, Daniel P. Friedman, Brian Harvey, Simon Kaplan, Richard Salter, and George Springer. Using SCHEME in the introductory computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):288, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robles:1997:UMC

[RFPA97] Tomás Robles, David Fernández, Encarna Pastor, and Santiago Alamillo. Using multimedia communication technologies in distance learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):6–7, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Riser:1998:LJT

[RG98a] Robert Riser and Donald Gotterbarn. On-line journal: a tool for enhancing student journals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 203–205, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:1998:JPA

[RG98b] Susan H. Rodger and Eric Gramond. JFLAP (poster): an aid to studying theorems in automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):302, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ruckert:1993:EC

[RH93] Martin Ruckert and Richard Halpern. Educational C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):6–9, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robinson:1999:JBT

[RHND99] Matthew B. Robinson, Jason A. Hamshar, Jorge E. Novillo, and Andrew T. Duchowski. A Java-based tool for reasoning about models of computation through simulating finite automata and Turing machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):105– 109, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rucinski:1990:PDR

[RHP90] Andrzej Rucinski, Frank Hludik, and John L. Pokoski. A partnership in domestication of rapid prototyping technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):1–5, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richardson:1998:FYI

[Ric98] A. S. Richardson. First year information systems papers — optimising learning — minimising administration (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):301, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rifkin:1994:EIE

[Rif94a] Adam Rifkin. eText: an interactive environment for learning parallel programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):281–285, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rifkin:1994:TPP

[Rif94b] Adam Rifkin. Teaching parallel programming and software engineering concepts to high school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):26–30, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosenberg:1997:TOO

[RK97] John Rosenberg and Michael Kölling. Testing object-oriented programs: making it simple. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):77–81, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roantree:1998:ACC

[RK98] Mark Roantree and Tia E. Keyes. Automated collection of coursework using the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):206–208, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramakrishnan:1993:OSP

[RL93] Sub Ramakrishnan and Ann-Marie Lancaster. Operating systems projects: linking theory, practice and use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):256–260, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robertson:1995:ASN

[RL95] Stephanie A. Robertson and Martin P. Lee. The application of second natural language acquisition pedagogy to the teaching of programming languages — a research agenda. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):9–12, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1995:UUT

[RLR95] Eric Roberts, John Lilly, and Bryan Rollins. Using undergraduates as teaching assistants in introductory programming courses: an update on the Stanford experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):48–52, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosbottom:1996:ITW

[RM96] John Rosbottom and Adrian Meads. Interactive tutorials on the World Wide Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):231, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosbottom:1998:UIA

[RM98] John Rosbottom and Claude Moulin. Using intelligent agents to change the delivery of education (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):303, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberge:1992:CPP

[Rob92] James Robergé. Creating programming projects with visual impact. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):230–234, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1993:UCC

[Rob93] Eric S. Roberts. Using C in CS1: evaluating the Stanford experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):117–121, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1995:CBG

[Rob95a] Eric S. Roberts. A C-based graphics library for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):163–167, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1995:LES

[Rob95b] Eric S. Roberts. Loop exits and structured programming: reopening the debate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):268–272, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1996:TCP

[Rob96] Eric Roberts. Tools for creating portable demonstration programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):78–80, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1998:SUT

[Rob98] Eric Roberts. Strategies for using technology in the teaching of ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):209–212, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1999:CSC

[Rob99a] Eric Roberts. Conserving the seed corn: reflections on the academic hiring crisis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):4–9, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robertson:1999:HMR

[Rob99b] John S. Robertson. How many recursive calls does a recursive function make? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):60–61, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:1995:ILA

[Rod95] Susan H. Rodger. An interactive lecture approach to teaching computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):278–282, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:1996:IAC

[Rod96] Susan H. Rodger. Integrating animations into courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):72–74, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roggio:1990:UPE

[Rog90] Robert F. Roggio. Upsilon Pi Epsilon (UPE): the role of the computing science honor society in computer science programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):59–62, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:1991:EDP

[Ros91a] Rockford J. Ross. Experience with the DYNAMOD program animator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):35–42, March 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossa:1991:MFU

[Ros91b] Robert F. Rossa. The master file update problem in AWK. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):58–60, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:1996:LTC

[Ros96] John Minor Ross. Learning to teach C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(2):25–30, June 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosbottom:1997:CMO

[Ros97] John Rosbottom. Computer managed, open question, open book assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):100–102, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossi:1999:KIB

[Ros99] Alfred J. Rossi. KPPCDL: an Internet based shared environment for introductory programming education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):196, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rout:1992:CQS

[Rou92] T. P. Rout. The culture of quality and software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):29–34, June 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rozanski:1994:ADI

[Roz94] Evelyn P. Rozanski. Accreditation: does it enhance quality? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):145–149, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1998:DJG

[RP98] Eric Roberts and Antoine Picard. Designing a Java graphics library for CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):213–218, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:1994:AAI

[RPF94] Richard Rasala, Viera K. Proulx, and Harriet J. Fell. From animation to analysis in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):61–65, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:1996:ECP

[RR96] Margaret M. Reek and Kenneth A. Reek. An electronic class photo album. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):15–18, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramakrishnan:1997:CPD

[RR97] Sub Ramakrishnan and B. Madhu Rao. Classroom projects on database connectivity and the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):116–120, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:1999:EEU

[RR99] Steven Robbins and Kay A. Robbins. Empirical exploration in undergraduate operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):311–315, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rybacki:1993:ESA

[RRR93] Richard Rybacki, Kay A. Robbins, and Steven Robbins. Ethercom: a study of audio processes and synchronization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):218–222, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberge:1991:ELW

[RS91] James Robergé and Candice Suriano. Embedding laboratories within the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):6–10, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberge:1994:ULT

[RS94] James Robergé and Candice Suriano. Using laboratories to teach software engineering principles in the introductory computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):106–110, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rada:1997:NSE

[RS97] Roy Rada and James Schoening. New standards for educational technology relevant to multiple ACM SIGs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(2):28–30, June 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosen:1998:AIH

[RSB98] C. Rosen, D. Simper, and J. I. Bostock. An augmented induction for HND computer studies students through an early integrated, inter-disciplinary assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 219–222, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ragonis:1998:DIE

[RSBAS98] Noa Ragonis, Ehud Shapiro, Mordechai Ben-Ari, and Zahava Scherz. Development, implementation and evaluation of a course in expert systems for high-school students (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):300, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1999:CIR

[RSLD99] Eric Roberts, Russell L. Shackelford, Richard J. LeBlanc, and Peter J. Denning. Curriculum 2001: interim report from the ACM/IEEE-CS task force. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):343–344, March 1999. CODEN SIGSD3. ISBN 1-58113-085-6. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1999.bib;   ftp:/

/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/leavens.bib.

Rautama:1997:EAA

[RST97] E. Rautama, E. Sutinen, and J. Tarhio. Excel as an algorithm animation environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):24–26, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rowe:1998:VLT

[RT98] Glenn Rowe and Gareth Thorburn. VINCE-an on-line tutorial tool for teaching introductory programming (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):304, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rubio:1992:FCC

[Rub92] Juan Alvarez Rubio. A first computing course based on curricula 1991. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):5–8, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rubinstein:1995:CSP

[Rub95] Roy S. Rubinstein. Computer science projects with music. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):287–291, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rugelj:1999:IAK

[Rug99] Joze Rugelj. Intelligent agent for knowledge tree administration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):197, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:1996:AAH

[RW96] Susan H. Rodger and Ellen L. Walker. Activities to attract high school girls to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):373–377, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schweitzer:1995:IIC

[SA95] Dino Schweitzer and Tom Appolloni. Integrating introductory courses in computer graphics and animation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):186–190, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sabin:1990:UCP

[Sab90] Roberta E. Sabin. Using CS2 projects to introduce computer science concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):25–28, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Saiedian:1993:TMF

[Sai93] Hossein Saiedian. Towards more formalism in software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):193– 197, March 1993. CODEN SIGSD3. ISBN 0-89791-565-8. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/1993.bib;   ftp://ftp.math.utah.
edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/Pape.
bib.

Saiedian:1997:DFS

[Sai97] Hossein Saiedian. Developing formal specifications via Z (workshop). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):396, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shub:1990:FFM

[SAL^{+}90] Charles M. Shub, Owen Astrachan, David Levine, Stuart Reges, and Henry Walker. Faster, fairer, and more consistent grading, using techniques from the advanced placement reading (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):266, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sapega:1991:IUE

[Sap91] August E. Sapega. Introducing undergraduate engineering students to laboratory automation using high-level application programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):1–5, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schaller:1997:UJC

[SBB^{+}97] Nan C. Schaller, Michael Berman, Judith Bishop, Paddy Nixon, Evelyn Rozanski, and Peter Welch. Using Java in computer science education (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 140–142, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stone:1999:CML

[SBB^{+}99] Don C. Stone, Seth Bergmann, Ganesh Baliga, A. Michael Berman, and John Schmalzel. A CS1 maze lab, using joysticks and MIPPETs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):170–173, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Starkey:1997:AIC

[SBD97] J. Denbigh Starkey, Ray S. Babcock, and Anne S. DeFrance. An accelerated introductory computer science course sequence for non-traditional Master’s students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):145–149, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Seffah:1999:AOO

[SBD99] Ahmed Seffah, Moncef Bari, and Michel Desmarais. Assessing object-oriented technology skills using an Internet-based system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):71–74, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scragg:1994:CSN

[SBK94] Greg Scragg, Doug Baldwin, and Hans Koomen. Computer science needs an insight-based curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):150–154, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:1994:HII

[SBTC94] Thomas J. Scott, Ralph B. Bisland, Jr., Lee H. Tichenor, and James H. Cross II. Handling interpersonal issues for student team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):397–398, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shum:1994:ULP

[SC94] Stephen Shum and Curtis Cook. Using literate programming to teach good programming practices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 66–70, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scafide:1996:FCC

[Sca96] Jean Scafide. A first course in C++ and object oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4):25–32, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shiflet:1993:ATP

[SCF^{+}93] Angela B. Shiflet, Scott Cannon, Terry J. Frederick, Janet Hartman, Marsha Meredith, Chris Nevison, and Dean Sanders. Approaches to teaching parallel processing on the undergraduate level (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):298, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schwietzer:1990:RTM

[Sch90] Dino Schwietzer. Ray tracing: a means to motivate students in an introductory graphics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 157–161, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schaller:1991:CGP

[Sch91] Nan C. Schaller. Computer graphics and parallelism, an interdisciplinary fable. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):61–64, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schweitzer:1992:DIV

[Sch92] D. Schweitzer. Designing interactive visualization tools for the graphics classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):299–303, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schorsch:1995:CAS

[Sch95] Tom Schorsch. CAP: an automated self-assessment tool to check Pascal programs for syntax, logic and style errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):168–172, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schill:1996:ITB

[Sch96a] Alexander Schill. Interdisciplinary teleteaching based on WWW and internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):230, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schollmeyer:1996:CPH

[Sch96b] Martina Schollmeyer. Computer programming in high school vs. college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):378–382, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schoenefeld:1997:OOD

[Sch97] Dale A. Schoenefeld. Object-oriented design and programming: an Eiffel, C++, and Java course for C programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):135–139, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schneider:1999:CSI

[Sch99] G. Michael Schneider. Computational science as an interdisciplinary bridge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):141–145, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:1991:MCS

[Sco91] Thomas J. Scott. Mathematics and computer science at odds over real numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):130– 139, March 1991. CODEN SIGSD3. ISBN 0-89791-377-9. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Math/fparith.bib;   ftp://ftp.math.utah.
edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

1991.bib. 22nd SIGCSE Technical Symposium on Computer Science Education.

Scragg:1991:MCO

[Scr91] Greg W. Scragg. Most computer organization courses are built upside down. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):341–346, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanchez-Calle:1991:FRP

[SCVI91] Angel Sánchez-Calle and J. Angel Velázquez-Iturbide. Fun, rigour and pragmatism in functional programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):11–16, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Syang:1993:CAT

[SD93] Angel Syang and Nell B. Dale. Computerized adaptive testing in computer science: assessing student programming abilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):53–56, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaffer:1994:OIT

[SDJ^{+}94] Dale Shaffer, George J. Davis, Keith Jolly, Martyn Roberts, and Miriam Roy. Opportunities in international teacher exchanges (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):372, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Searls:1993:IHS

[Sea93] Delmar E. Searls. An integrated hardware simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2):24–28, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sebern:1997:IDC

[Seb97] Mark J. Sebern. Iterative development and commercial tools in an undergraduate software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):306–309, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Saiedian:1991:TOP

[SF91] Hossein Saiedian and Hassan Farhat. A team-oriented, projectintensive database course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):192–198, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stehlik:1998:APT

[SFR^{+}98] Mark Stehlik, Sarah Fix, Susan Rodger, Chris Nevison, and Mark Weiss. Advanced placement transition to C++ (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):372, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:1991:AAA

[SG91] Ian Sanders and Harshila Gopal. AAPT: algorithm animator and programming toolbox. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):41–47, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schulze:1996:TEI

[SG96] Kay G. Schulze and Frances S. Grodzinsky. Teaching ethical issues in computer science: what worked and what didn’t. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):98–101, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schulze:1997:TES

[SG97a] Kay G. Schulze and Frances S. Grodzinsky. Teaching ethical and social issues in CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 6–9, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shifroni:1997:SGT

[SG97b] Eyal Shifroni and David Ginat. Simulation game for teaching communications protocols. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):184–188, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:1999:CIG

[SG99a] Ann E. Kelley Sobel and Mario Guimaraes. Current issues in graduate student research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 351, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:1999:CIU

[SG99b] Ann E. Kelley Sobel and Mario Guimaraes. Current issues in undergraduate student research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 350, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:1990:GSP

[SH90] Dean Sanders and Janet Hartman. Getting started with parallel programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):86–88, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skrien:1991:MSR

[SH91] Dale Skrien and John Hosack. A multilevel simulator at the register transfer level for use in an introductory machine organization class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):347–351, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scherz:1995:LPB

[SH95] Zahava Scherz and Bruria Haberman. Logic programming based curriculum for high school students: the use of abstract data types. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):331–335, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sharp:1996:MMA

[SH96] Helen Sharp and Pat Hall. A multi-media approach to providing software project experience for postgraduate students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):109–115, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sondergaard:1997:PAS

[SH97] Harald Søndergaard and A. John Hurst, editors. Proceedings of the ACM SIGCSE 2nd Australasian Conference on Computer Science Education, ACSE 1997, Melbourne, Victoria, Australia, 1997, volume 2 of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1997. CODEN SIGSD3. ISBN 0-89791-958-0. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:1998:OFS

[SH98] Judy Sheard and Dianne Hagan. Our failing students: a study of a repeat group. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):223–227, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:1999:CMT

[SH99a] Dean Sanders and Janet Hartman. A compendium of multimedia tools for all budgets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):369, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:1999:SLE

[SH99b] Judy Sheard and Dianne Hagan. A special learning environment for repeat students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):56–59, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:1999:ELE

[SH99c] Ann E. Kelley Sobel and Joseph C. Hill. Enhancing the learning environment of deaf students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):71–72, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shay:1991:SPD

[Sha91] William A. Shay. A software project for a data communication course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):15–20, March 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shannon:1998:LPI

[Sha98] Christine Shannon. LongNbrs and palindromes: an interesting application of a class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):190–193, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shannon:1999:SLP

[Sha99] Christine Shannon. A semester-long project in CS 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):198, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sherry:1995:MCS

[She95] Lorraine Sherry. A model computer simulation as an epistemic game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):59–64, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shene:1998:MPI

[She98] Chin-Kuang Shene. Multithreaded programming in an introduction to operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):242–246, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shiflet:1992:PPC

[Shi92] Angela B. Shiflet. The publication process for computer science textbooks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(2):15–19, June 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shiau:1993:OOP

[Shi93a] Lie June Shiau. Object-oriented programming for numerical methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):5–8, December 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shiau:1993:UHA

[Shi93b] LieJune Shiau. Using a heuristic algorithm to help analyze graph problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):18–20, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shub:1990:SUE

[Shu90] Charles M. Shub. Should undergraduates explore internals of workstation operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1): 111–115, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaffer:1996:USD

[SHY96] Clifford A. Shaffer, Lenwood S. Heath, and Jun Yang. Using the Swan data structure visualization system for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):140–144, March 1996. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1996.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/Pape.bib.

Siegel:1999:WDF

[Sie99] Eric V. Siegel. Why do fools fall into infinite loops: singing to your computer science class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):

167–170, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sigle:1990:DDC

[Sig90] John Sigle. Dynamic display of computing processes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):2–4, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Silver:1990:BQD

[Sil90] Charles L. Silver. A better queue design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):63–64, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Silver:1991:UAS

[Sil91] James L. Silver. Using Ada to specify and evaluate projects in a data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):337– 340, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Silver:1994:MAD

[Sil94] Charles L. Silver. Mathematizing abstract data types. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):57–60, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Siviter:1996:CAM

[Siv96] D. Siviter. Computer assistance for managing educational resources and managing collaborative educational processes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):104–106, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shinners-Kennedy:1994:GDC

[SK94] Dermot Shinners-Kennedy. A graduate diploma in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):58–64, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schaller:1995:ETP

[SK95] Nan C. Schaller and Andrew T. Kitchen. Experiences in teaching parallel computing-five years later. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3): 15–20, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sajaniemi:1998:SML

[SK98] Jorma Sajaniemi and Marja Kuittinen. Systems for multi-level teaching materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):59–61, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skelton:1993:ITQ

[Ske93] Gordon W. Skelton. Integrating total quality management with software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(2): 29–30, June 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sangwan:1998:SPV

[SKL98] Raghvinder S. Sangwan, James F. Korsh, and Paul S. LaFollette, Jr. A system for program visualization in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):272–276, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shackelford:1994:IDF

[SL94] Russell L. Shackelford and Richard J. LeBlanc. Integrating “depth first” and “breadth first” models of computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):6–10, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Szpiniak:1998:OET

[SLM98] Ariel Ferreira Szpiniak, Carlos D. Luna, and Ricardo H. Medel. Our experiences teaching functional programming at University of Río Cuarto (Argentina). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2): 28–30, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Slonneger:1991:EDS

[Slo91] Ken Slonneger. An exercise in denotational semantics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):178–183, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Slonneger:1995:ESM

[Slo95] Ken Slonneger. Executing an SECD machine using logic programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):228–232, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1991:DIU

[Smi91] Wayne D. Smith. The design of an inexpensive undergraduate data communications laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (1):273–276, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1994:PPA

[Smi94a] Harry F. Smith. In pursuit of parallelism (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):409, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1994:LSF

[Smi94b] Wayne D. Smith. A laboratory to support a first course in data communications using personal computers and Turbo Pascal (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):404, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1998:STC

[Smi98] Matt Smith. Structuring team communications with targeted use of groupware (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):306, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:1996:EIF

[Sob96] Ann E. Kelley Sobel. Experience integrating a formal method into a software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):271–274, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Song:1996:TSE

[Son96] Ki-Sang Song. Teaching software engineering through real-life projects to bridge school and industry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(4): 59–64, December 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sackrowitz:1996:UPF

[SP96] Marian Gunsher Sackrowitz and Ann Parker Parelius. An unlevel playing field: women in the introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):37–41, March 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scherz:1999:OPB

[SP99] Zahava Scherz and Sarah Polak. An organizer for project-based learning and instruction in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):88–90, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1996:IPE

[SPSS96] Harry F. Smith, Patrick Plusnick, Mark Sarojak, and William Seitz. Image processing as an exemplar of parallelism applied to graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):363–367, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stegink:1999:CSG

[SPV99] Gordon Stegink, Jon Pater, and Daron Vroon. Computer science and general education: Java, graphics, and the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):146–149, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stanchev:1991:TSM

[SR91] Svetlin Stanchev and Atanas Radensky. Teaching some modern functional programming concepts: an approach based on an extended FP-like language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):31–36, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sivalingam:1999:EWA

[SR99] Krishna M. Sivalingam and V. Rajaravivarma. Education of wireless and ATM networking concepts using hands-on laboratory experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):114–118, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:1999:DVE

[SRJ99] Todd Smith, Anthony Ruocco, and Bernard Jansen. Digital video in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):122–126, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stehlik:1999:CFD

[SRL^{+}99] Mark Stehlik, Susan H. Rodger, Kathy Larson, Alyce Brady, and Chris Nevison. Current and future direction of the advanced placement exam. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):358, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Selby:1998:WCW

[SRY98] Linda Selby, Ken Ryba, and Alison Young. Women in computing: what does the data show? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):62–67, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sabin:1994:CLI

[SS94] Roberta Evans Sabin and Edward P. Sabin. Collaborative learning in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):304–308, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Simeonov:1995:MIM

[SS95] Simeon Simeonov and G. Michael Schneider. MSIM: an improved microcode simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):13–17, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spooner:1997:SEC

[SS97] David L. Spooner and Michael Skolnick. Science and engineering case studies in introductory computing courses for non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):154–158, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scragg:1998:SBW

[SS98] Greg Scragg and Jesse Smith. A study of barriers to women in undergraduate computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 82–86, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schneider:1998:TCS

[SSH98] G. Michael Schneider, Daniel Schwalbe, and Thomas M. Halverson. Teaching computational science in a liberal arts environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):57–60, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stern:1999:SMC

[SSN99] Linda Stern, Harald Søndergaard, and Lee Naish. A strategy for managing content complexity in algorithm animation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):127–130, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Staehr:1993:DUT

[Sta93] Lorraine Staehr. Debating: its use in teaching social aspects of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):46–49, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stasko:1997:USB

[Sta97] John T. Stasko. Using student-built algorithm animations as learning aids. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):25–29, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:1994:TDS

[STBC94] Thomas J. Scott, Lee H. Tichenor, Ralph B. Bisland, Jr., and James H. Cross II. Team dynamics in student programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):111–115, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Steidley:1991:RCL

[Ste91] Carl W. Steidley. Robotics: a closer look at microprocessor systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):265–272, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stewart:1994:DSU

[Ste94] Carolee Stewart. Distributed systems in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):17–20, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Studer:1995:YSI

[STM95] Scott D. Studer, James Taylor, and Ken Macie. Youngster: a simplified introduction to computing: removing the details so that a child may program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):102– 105, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Struble:1991:EHH

[Str91] George Struble. Experience hosting a high-school level programming contest. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2):36–38, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stratton:1999:TNF

[Str99] David Stratton. Teaching network fundamentals using a simulated network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):62–64, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stuart:1994:CPA

[Stu94] Brian L. Stuart. Connect 4 as a problem in artificial intelligence and robotics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):41–46, June 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Styer:1994:DUS

[Sty94] Eugene Styer. On the design and use of a simulator for teaching computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):45–ff., September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sullivan:1993:SPM

[Sul93] Sarah L. Sullivan. A software project management course role-play-team-project approach emphasizing written and oral communication skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):283–287, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sullivan:1994:RPR

[Sul94] Sarah L. Sullivan. Reciprocal peer reviews. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):314–318, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanchez:1996:VVL

[SV96] Javier Sanchez and Jordi Vitria. ViLi (Vision LISP): a software environment for teaching image processing and analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):234, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheets:1994:CUV

[SVFB94] Michael A. Sheets, Michael A. Vance, Jeremy A. Freed, and David W. Brown. Compuquake: upheaval of a virtual community (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):410, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:1993:TEC

[SVP93] Thomas J. Scott, Richard B. Voss, and Cherri M. Pancake. Teaching an ethics component to computer science majors (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):304, March 1993. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skublics:1991:TSF

[SW91] Suzanne Skublics and Paul White. Teaching Smalltalk as a first programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):231– 234, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schoenefeld:1993:IDM

[SW93] Dale A. Schoenefeld and Roger L. Wainwright. Integration of discrete mathematics topics into the secondary mathematics curriculum using Mathematica: a summer institute for high school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):78–82, March 1993. CODEN SIGSD3. ISBN 0-89791-565-8. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic).  URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Math/mathematica.bib;
ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/1993.bib.

Sears:1995:VAA

[SW95] Andrew Sears and Rosalee Wolfe. Visual analysis: adding breadth to a computer graphics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):195–198, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skrzypek:1999:KVL

[SW99] Jerzy Skrzypek and Tadeusz Wilusz. KOMBIZ97 — virtual lab experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):204, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spies:1997:VPT

[SWWB97] Wayne Spies, Lee Wittenberg, Robert Workman, and Baarry Burd. Visual programming tools in the C.S. curriculum (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):388–389, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Selby:1998:TPP

[SY98] Linda Selby and Alison Young. The teaching of programming (poster): changing the delivery to meet all students’ needs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):305, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Szejko:1999:EES

[Sze99] Stanisław Szejko. An exercise in evaluating significance of software quality criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):199, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Szuecs:1996:CWA

[Szu96] Laszlo Szuecs. Creating windows applications using Borland’s OWL classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):145–149, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Towhidnejad:1996:SEE

[TA96] Massood Towhidnejad and James R. Aman. Software engineering emphasis in advanced courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):210–213, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taffe:1991:SMS

[Taf91] William J. Taffe. Simulation and modelling with stella: a general education course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):87–91, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tam:1992:TLI

[Tam92] Wing C. Tam. Teaching loop invariants to beginners by examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):92–96, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tangorra:1990:RCA

[Tan90] Filippo Tangorra. The role of the computer architecture simulator in the laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):5–10, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tang:1992:CSS

[Tan92] L. S. Tang. A CPU scheduling simulation from structured programming to object-oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (1):129–133, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tao:1995:UES

[Tao95] Yonglei Tao. Using expert systems to understand objectoriented behavior. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):14–18, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1996:IMT

[Tay96] Harriet G. Taylor. Integrating multimedia and telecommunications laboratory experiences into teacher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):230, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1997:ESA

[Tay97a] Harriet Taylor. The evolution of standards for accrediting computer science teacher preparation programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):67–71, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1997:UOS

[Tay97b] Jacqueline Taylor. Using online seminars to demonstrate the social psychological impacts of computer-mediated communication systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):80–84, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1998:MCT

[Tay98] R. Gregory Taylor. Motivating the Church–Turing thesis in the twenty-first century. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):228–231, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1999:INE

[Tay99a] Harriet G. Taylor. Implementing national educational technology standards for student in the United States. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):200, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1999:MLL

[Tay99b] Harriet G. Taylor. Math link: linking curriculum, instructional strategies, and technology to enhance teaching and learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):201, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tinoco:1997:OEW

[TBF97] Lucio C. Tinoco, N. Dwight Barnette, and Edward A. Fox. Online evaluation in WWW-based courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):194–198, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tomer:1998:IMT

[TBF98] D. S. Tomer, Doug Baldwin, and Christopher J. Fox. Integration of mathematical topics in CS1 and CS2 (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):365, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Teague:1993:AWT

[TC93] G. Joy Teague and Val A. Clarke. Attracting women to tertiary computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):208–212, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1996:DEI

[TCP^{+}96] Pete Thomas, Linda Carswell, Marian Petre, Barbara Poniatowska, Blaine Price, and Judy Emms. Distance education over the Internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):147–149, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Toal:1990:SEG

[TD90] Raymond J. Toal and Philip M. Dorin. Software engineering and the game of Monopoly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):2–4, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Teague:1992:RSC

[Tea92] Joy Teague. Raising the self confidence and self esteem of final year female students prior to job interviews. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):67–71, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Temte:1991:LBI

[Tem91] Mark C. Temte. Let’s begin introducing the object-oriented paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):73–77, March 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Temte:1992:CCP

[Tem92] Mark C. Temte. A compiler construction project for an objectoriented language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):138–141, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Terry:1995:UIP

[Ter95] P. D. Terry. Umbriel-imperative programming for unsophisticated students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):7–14, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tewari:1995:SRO

[Tew95] Rajiv Tewari. Software reuse and object-oriented software engineering in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):253–257, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tewari:1994:OOL

[TG94] Rajiv Tewari and David Gitlin. On object-oriented libraries in the undergraduate curriculum: importance and effectiveness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):319–323, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Theotokis:1997:VVI

[TGGP97] Dimitrios Theotokis, George Gyftodimos, Panagiotis Georgiadis, and George Philokyprou. VIBDaST: a virtual interactive book as a data structures teaching aid. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):111–113, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tiwari:1998:LGT

[TH98] Ashok Tiwari and Clive Holtham. Learning groupware through using groupware-computer supported collaborative learning with face to face students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):236–238, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1990:MMC

[Tho90] Paul A. V. Thomas. Micron: a microprogrammed computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):11–20, September 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1998:OSE

[Tho98a] Pete G. Thomas. Observing students electronically as they learn (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):307, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1998:GWG

[Tho98b] Stan Thomas. Group 4 (working group): the impact of campus-wide portable computing on computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):270, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thurente:1990:SUC

[Thu90] David J. Thurente. Simulation in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):53–57, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thweatt:1994:CCL

[Thw94] Mack Thweatt. CSI closed lab vs. open lab experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):80–82, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tichenor:1995:IAG

[Tic95] Lee H. Tichenor. Inexpensive advanced graphics applications for the C.S. majors graphics class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):191–194, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tu:1990:CCP

[TJ90] Jho-Ju Tu and John R. Johnson. Can computer programming improve problem-solving ability? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2): 30–33, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tjaden:1996:HVS

[Tja96] Bunny J. Tjaden. How visual software influences learning in college students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):229, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tjaden:1998:DLM

[Tja98] Bunny J. Tjaden. Do lab modules in CS actually help students?: an empirical study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):282– 286, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1998:ICW

[TLNS98] Stan J. Thomas, Cary Laxer, Tomohiro Nishida, and Helen Sherlock. The impact of campus-wide portable computing on computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):39–44, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:1994:DUS

[TLR94] Paul T. Tymann, Douglas Lea, and Rajendra K. Raj. Developing an undergraduate software engineering program in a liberal arts college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):276–280, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1992:INA

[TM92] Harriet G. Taylor and C. Dianne Martin. The impact of new accreditation and certification standards for secondary computer science teachers on university computer science departments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):235–239, March 1992.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Toll:1995:DPI

[Tol95a] William E. Toll. Decision points in the introduction of parallel processing into the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):136–140, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Toll:1995:SPD

[Tol95b] William E. Toll. Socket programming in the data communications laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):39–43, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Toll:1998:DCC

[Tol98] Bill Toll. The distributed course — a curriculum design paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):20–24, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tomayko:1991:TSD

[Tom91] James E. Tomayko. Teaching software development in a studio environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):300–303, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tosten:1993:UMR

[Tos93] Rodney S. Tosten. Using a model railroad system in an artificial intelligence and operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):30–32, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:1996:VVT

[Tow96] Gloria Childress Townsend. Viewing video-taped role models improves female attitudes toward computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):42–46, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:1998:TLA

[Tow98] Gloria Childress Townsend. Turning liabilities into assets in a general education course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):58–62, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:1999:TES

[Tow99] Gloria Childress Townsend. ThetenthStrand == 3 * ethicaldebates + solution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):17–21, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1993:IAD

[TP93] Harriet G. Taylor and Katharine M. Paine. An interdisciplinary approach to the development of writing skills in computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):274–278, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tinkham:1994:SOT

[TP94] Nancy Lynn Tinkham and Darren F. Provine. The stage one Turing test as an artificial intelligence class exercise. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2):5–8, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trahan:1993:SVT

[TR93] Robin Trahan and Susan Rodger. Simulation and visualization tools for teaching parallel merge sort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25 (1):237–241, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Traxler:1998:HDC

[Tra98] John Traxler. How do computer science lecturers create modules? (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):308, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tremblay:1998:UCF

[Tre98] Guy Tremblay. An undergraduate course in formal methods: “description is our business”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):166– 170, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tremblay:1999:ETL

[Tre99] Guy Tremblay. Exercises for teaching logic in a formal methods course: formalizing ERDs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):100– 104, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trichina:1999:DIT

[Tri99] Elena Trichina. Didactic instructional tool for topics in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):95–98, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trono:1991:ACA

[Tro91] John A. Trono. Average case analysis when merging two ordered lists of different length. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(3):5– 6, September 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trono:1992:UPC

[Tro92] John A. Trono. An undergraduate project to compute minimal perfect hashing functions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):53–56, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Troeger:1994:TFM

[Tro94a] Douglas R. Troeger. Tutorial-formal methods in the first course (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):378, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trono:1994:NEC

[Tro94b] John A. Trono. A new exercise in concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(3):8–10, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Trono:1994:TR

[Tro94c] John A. Trono. Taxman revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (4):56–58, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tillison:1995:GWC

[TS95] Julie Tillison and Ching-Kuang Shene. On generating worstcases for the insertion sort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):57–58, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Toothman:1998:EPI

[TS98] Brian Toothman and Russell Shackelford. The effects of partially-individualized assignments on subsequent student performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):287–291, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

TenEyck:1998:SAD

[TSG98] James Ten Eyck, G. Sampath, and Richard Goldstone. Specification of an algorithm design system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 232–235, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tureman:1994:CLS

[Tur94] Robert L. Tureman, Jr. Computing laboratories and the small community college: defining the directed computing laboratory in the small college computing environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):160–163, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:1992:DAI

[TW92] Pamela A. Taylor and Dana L. Wyatt. Database and artificial intelligence integration: a challenge to academia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(4):35–42, December 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tucker:1994:NDI

[TW94] Allen B. Tucker and Peter Wegner. New directions in the introductory computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1): 11–15, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turk:1997:TSE

[TW97] Jane Turk and Samuel J. Wiley. Teaching social and ethical issues in the literacy course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):10–14, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:1991:VTT

[Tym91] Paul Tymann. VNET: a tool for teaching computer networking to undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):21–24, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:1999:UCL

[TZ99] Joseph A. Turner and Joseph L. Zachary. Using course-long programming projects in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):43–47, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Uckan:1991:ILP

[Uck91] Yuksel Uckan. Integrating logic programming into a data base course: views as rules in deductive relational data bases. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):184–191, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urban:1997:IPU

[UD97] Susan D. Urban and Suzanne W. Dietrich. Integrating the practical use of a database product into a theoretical curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):121–125, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urban-Lurain:1999:DUM

[ULW99] Mark Urban-Lurain and Donald J. Weinshank. “I do and I understand”: mastery model learning for a large non-major course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):150–154, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Utting:1999:GDG

[Utt99] Ian Utting. Gathering and disseminating good practice at teaching and learning conferences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):202, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanWyk:1995:PWU

[Van95] Christopher J. Van Wyk. Programming as writing: using portfolios. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):39–42, December 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanScoy:1998:UBC

[Van98] Frances L. Van Scoy. Using the BURKS 2 CD-ROM in a principles of programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 239–242, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Villarreal:1998:GCS

[VB98] E. E. Villarreal and Dennis Butler. Giving computer science students real-world experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):40–44, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vest:1998:RWC

[VBB^{+}98] Sharon N. Vest, Robert Biddle, Christina Bjorkman, Linda Null, Eric Roberts, and Greg Scragg. The retention of women in the computing sciences (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):366–367, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Villadeamigo:1997:EES

[VdlFCC97] José M. Pérez Villadeamigo, Santiago Rodríguez de la Fuente, Rafael Méndez Cavanillas, and M. Isabel García Clemente. The em88110: emulating a superscalar processor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):45–50, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:1994:MMD

[VI94] J. Ángel Velázquez-Iturbide. A methodology for monitor development in concurrent programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (3):22–28, September 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:1996:HTL

[VI96] J. Ángel Velázquez-Iturbide. HIPE: a Turbo-like environment for functional programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):234, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:1996:IPT

[VIRF96] J. Ángel Velázquez-Iturbide, Ana Maria Cosias Rincon, and Nuria Llorente Fernandez. An inexpensive programming toolset for algorithm development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):231, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Varden:1990:FIC

[VL90] Stuart A. Varden and Frank J. LoSacco. Facilitating intracorporate cooperation: a university creates the environment.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):152–156, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanMeer:1992:ISA

[VS92] Gretchen L. Van Meer and Charles D. Sigwart. Introducing students to adaptive computer devices: for compliance with the Americans with Disabilities Act of 1990. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):19–23, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Veraart:1996:SSE

[VW96a] V. E. Veraart and S. L. Wright. Supporting software engineering education with a local Web site. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (1):275–279, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Veraart:1996:UCR

[VW96b] V. E. Veraart and S. L. Wright. Using CD-ROMs and local Web pages to provide course materials for distance-students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):90–92, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wadkins:1995:RPP

[Wad95] J. R. Jefferson Wadkins. Rigorous proofs of program correctness without formal logic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):307–311, March 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/1995.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/Pape.bib; ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/SE/alspaugh.bib.

Wahl:1999:YYA

[Wah99] Nancy J. Wahl. YAATCE-yet another approach to teaching computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):22–26, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wainwright:1992:IFP

[Wai92] Roger L. Wainwright. Introducing functional programming in discrete mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):147–152, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilson:1996:RAS

[WAK96] Judith Wilson, Robert Aiken, and Irvin Katz. Review of animation systems for algorithm understanding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):75–77, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1990:PVT

[Wal90] Henry M. Walker. Program verification (tutorial session): techniques and uses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):254, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waller:1994:FCC

[Wal94] William A. Waller. A framework for CS1 and CS2 laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):198–202, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1995:MSS

[Wal95] Henry M. Walker. Message from the SIGCSE Secretary/ Treasurer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(4):1–4, December 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wallingford:1996:TFC

[Wal96] Eugene Wallingford. Toward a first course based on objectoriented patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):27–31, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1997:CLC

[Wal97a] Henry M. Walker. Collaborative learning: a case study for CS1 at Grinnell College and Austin. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):209–213, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1997:RVS

[Wal97b] Henry M. Walker. A racquetball or volleyball simulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):22–23, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1998:BBP

[Wal98a] Henry M. Walker. The balance between programming and other assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):23–25, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1998:MIA

[Wal98b] Henry M. Walker. Modules to introduce assertions and loop invariants informally within CS1: experiences and observations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):31–35, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1998:WWC

[Wal98c] Henry M. Walker. Writing within the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):24–25, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:FIE

[Wal99a] Henry M. Walker. Finding interesting examples and assignments for CS1/2 assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):24–25, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:GRS

[Wal99b] Henry M. Walker. Guided reading and seminar classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4):27–28, December 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:RSS

[Wal99c] Henry M. Walker. Report from the SIGCSE secretary/ treasurer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):92–93, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:POS

[Wal99d] Robert A. Walker. A practical one-semester “VLSI design” course for computer science (and other) majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):237–241, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wans:1998:IML

[Wan98] Claudia Wans. An interactive multimedia learning system for the postlingually deaf (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):309, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Warford:1992:GPR

[War92] J. S. Warford. Good pedagogical random number generators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):142–146, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Warford:1995:ETF

[War95] J. Stanley Warford. An experience teaching formal methods in discrete mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):60–64, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Warford:1999:BNO

[War99] J. Stanley Warford. BlackBox: a new object-oriented framework for CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):271–275, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Withers:1992:EOS

[WB92] James M. Withers and Mark B. Bilodeau. An examination of operating systems laboratory techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):60–64, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Witten:1993:GRS

[WB93] Ian H. Witten and Timothy C. Bell. Getting research students started: a tale of two courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):165– 169, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:TWD

[WB99] Ellen L. Walker and Logan Browne. Teaching Web development with limited resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):12–16, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1994:DCF

[WBBW94] Henry M. Walker, Kim Bruce, James Bradley, and Tom Whaley. Describing the CS forest to undergraduates (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):369–370, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1998:SPA

[WBC^{+}98] Henry M. Walker, Laura Baker, Gail Chapman, Kathleen Larson, and Joseph Kmoch. SIGCSE panel: approaches for encouraging high school/college faculty interaction (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):368–369, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1995:CSM

[WBCS95] Henry M. Walker, Nancy Baxter, Robert Cupper, and G. Michael Schneider. The computer science major within a liberal arts environment (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):378–379, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1994:IMP

[WC94] Ursula Wolz and Edward Conjura. Integrating mathematics and programming into a three tiered model for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):223–227, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1992:LDS

[WCC^{+}92] D. K. Walker, H. Chahryar, J. Chaudri, A. Lodgher, E. Majdalani, and David S. Tucker. Lectureroom demonstration of serial data communications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):264– 267, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1995:FMU

[WCDT95] Henry M. Walker, H. Conrad Cunningham, Ruth Davis, and Douglas Troeger. Formal methods in the undergraduate computer science curriculum (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):398–399, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1994:CSL

[WCMT94] David K. Walker, Hamid Chahryar, James W. Moore, and David S. Tucker. Computer science lecture room demonstration experiment in event counting (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):408, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woodworth:1999:ICE

[WD99] Pat Woodworth and Wanda Dann. Integrating console and event-driven models in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):132– 135, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wu:1998:CMC

[WDB98] Cheng-Chih Wu, Nell B. Dale, and Lowell J. Bethel. Conceptual models and cognitive learning styles in teaching recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):292–296, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woodman:1996:JSS

[WDH96] Mark Woodman, Gordon Davies, and Simon Holland. The joy of software-starting with objects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1): 88–92, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1997:MMI

[WDM97] Ursula Wolz, Daniel Domen, and Michael McAuliffe. Multimedia integrated into CS 2: an interactive children’s story as a unifying class project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):103–110, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Webber:1996:PT

[Web96] Adam Brooks Webber. The Pascal trainer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):261–265, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weiss:1997:ETD

[Wei97] Mark Allen Weiss. Experiences teaching data structures with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):164–168, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

White:1997:PST

[WEKM97] Curt M. White, Carl Erickson, Bruce J. Klein, and James E. Miller, editors. Proceedings of the 28th SIGCSE Technical Symposium on Computer Science Education, 1997, San Jose, California, USA, February 27 — March 1, 1997, volume 29(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1997. CODEN SIGSD3. ISBN 0-89791-889-4. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wenner:1990:LCC

[Wen90] Patricia Wenner. The laboratory component of a computer organization course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):197–200, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Werth:1990:GIS

[Wer90] Laurie Honour Werth. Graphical interfaces as software engineering projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):12–16, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Werth:1997:GSC

[Wer97] Laurie H. Werth. Getting started with computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):1–5, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Werth:1999:LSP

[Wer99] Laurie Honour Werth. Licensing software professionals: where are we? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):27–30, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wills:1994:PLI

[WFGW94] Craig E. Wills, David Finkel, Michael A. Gennert, and Matthew O. Ward. Peer learning in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):309–313, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weis:1993:CPU

[WG93] Richard L. Weis and Judith L. Gersting. A course on professionalism in the undergraduate CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):170–174, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Werth:1995:PJA

[WGD^{+}95] John Werth, John Gorgone, Gordon Davis, David Feinstein, Bart Longenecker, and George Kasper. Proposed joint ACM/ DPMA/AIS undergraduate information systems degree curriculum model (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):392–393, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woodman:1997:OSU

[WGHL97] Mark Woodman, Robert Griffiths, Simon Holland, and Andrew Law. The object shop-using CD-ROM multimedia to introduce object concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):345–349, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woodman:1996:SUS

[WH96] Mark Woodman and Simon Holland. From software user to software author: an initial pedagogy for introductory objectoriented computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):60–62, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Whaley:1991:FPV

[Wha91] Tom Whaley. A framework for program verification in the context of linked structures and pointer variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):119–123, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Whitworth:1998:WSP

[Whi98] Brian Whitworth. The web of system properties: a general view of systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):46–50, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilson:1993:BCS

[WHN93] Judith D. Wilson, Nathan Hoskin, and John T. Nosek. The benefits of collaboration for student programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):160–164, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:1995:UCO

[Wic95] Michael R. Wick. On using C++ and object-orientation in CS1: the message is still more important than the medium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):322–326, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wildenberg:1990:USA

[Wil90] Gerald Wildenberg. Using a stack assembler language in a compiler course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):43–44, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Willshire:1995:ODN

[Wil95a] Mary Jane Willshire. Old dogs, new tricks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):178–181, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilson:1995:IBF

[Wil95b] Ronald E. Wilson. Integrating a breadth-first curriculum with relevant programming projects in CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):214–217, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Williams:1997:ENG

[Wil97] Kathleen A. Williams. Educating the next generation of information specialists: industry and university collaborative learning pilot project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):350–354, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Winslow:1993:PSA

[Win93] Leon E. Winslow. Problem specification with action machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(1):189–192, March 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Winslow:1996:PPP

[Win96] Leon E. Winslow. Programming pedagogy — a psychological overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):17–22, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wise:1992:DSS

[Wis92] M. J. Wise. Detection of similarities in student programs: YAP’ing may be preferable to plague’ing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):268–271, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wise:1996:YID

[Wis96] Michael J. Wise. YAP3: Improved detection of similarities in computer program and other texts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):130–134, March 1996. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/1996.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Object/Nierstrasz.bib; http://citeseer.
ist.psu.edu/wise96yap.html.

Wittenberg:1996:ULPb

[Wit96] L. Wittenberg. Using literate programming notation in introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1): 437, ???? 1996. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1999:SJPa

[WK99a] Ursula Wolz and Elliot Koffman. simpleIO: a Java package for novice interactive and graphics programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):139–142, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1999:SJPb

[WK99b] Ursula Wolz and Elliot Koffman. simpleIO: a Java package for novice interactive and graphics programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):212, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wu:1996:VRL

[WLC96] Cheng-Chih Wu, Janet Mei-Chuen Lin, and Guey-Fa Chiou. Visualizing recursion and linked lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (SI):232, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woit:1998:LLP

[WM98] Denise M. Woit and David V. Mason. Lessons from on-line programming examinations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):257–259, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1999:CAH

[WMZ^{+}99] Henry M. Walker, J. Paul Myers, Jr., Stu Zweben, Allen B. Tucker, Jr., and Grant Braught. The crisis in academic hiring in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):362–363, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wohlgehagen:1996:CBH

[WN96] Kathleen Wohlgehagen and Cathleen Norris. Computer-based high school algebra I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):230, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wojciechowski:1998:TIK

[WNKK98] Adam Wojciechowski, Jerzy R. Nawrocki, Karolina Kups, and Michal Kosiedowski. TETE: an intelligent knowledge testing engine (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):310, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolfe:1992:RSP

[Wol92] J. L. Wolfe. Reviving systems programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):255–258, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolfer:1998:GSE

[Wol98] James Wolfer. Graduate software engineering across international boundaries (poster): the first courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):311, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1997:CMC

[WP97] Ursula Wolz and Jacob Palme. Computer mediated communication in collaborative educational settings (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):145, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wade:1998:EDD

[WP98] Vincent P. Wade and Conor Power. Evaluating the design and delivery of WWW based educational environments and courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):243–248, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Whitehurst:1998:USM

[WPI98] R. Alan Whitehurst, Christopher L. Powell, and Jason S. Izatt. Utilising the student model in distance learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):254–256, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolf:1990:ESA

[WR90] Walter Wolf and Evelyn Rozanski. Expert systems: an applied course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(4):23–24, December 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wagner:1997:PAR

[WR97] Thomas D. Wagner and Eugene K. Ressler. A practical approach to reinforcing concepts in introductory operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):44–47, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wegner:1997:SDC

[WRRT97] Peter Wegner, Eric Roberts, Roy Rada, and Allen B. Tucker. Strategic directions in computer science education (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):371–372, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ward:1998:JFP

[WS98] Robert Ward and Martin Smith. JavaScript as a first programming language for multimedia students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):249–253, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wester:1997:VPJ

[WSK97] Frank Wester, Marleen Sint, and Peter Kluit. Visual programming with Java; an alternative approach to introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):57–58, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:1991:MHD

[WT91] David K. Walker and David S. Tucker. Managing hard-disk PCs in the introductory laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (4):37–40, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wu:1992:PNF

[Wu92] Margaret S. Wu. The practical need for fourth normal form. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):19–23, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wrycza:1999:CDM

[WUGV99] Stanisław Wrycza, Thaddeus W. Usowicz, Andras Gabor, and Borut Verber. The challenges and directions of MIS curriculum development in respect of transformation of business requirements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):177–178, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wagner:1991:AIT

[WV91] Christian Wagner and John Vinsonhaler. An artificial intelligence theory of computer competency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(2): 45–50, June 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wallace:1991:TNN

[WW91] Susan R. Wallace and F. Layne Wallace. Two neural network programming assignments using arrays. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):43–47, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1996:TIP

[WWDM96] Ursula Wolz, Scott Weisgarber, Daniel Domen, and Michael McAuliffe. Teaching introductory programming in the multimedia world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(SI):57–59, 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yahya:1992:LCC

[Yah92] Adnan H. Yahya. Local considerations in computer science curricula development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(1):123–128, March 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:1999:UJS

[YB99] Andrew Yang and Yael Bachar. Using Java and the socket interface in teaching client/server programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):206, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:1999:PCS

[YD99] Lan Yang and Zhiqiang Ding. Promoting computer software intellectual property right in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):115–118, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Young:1999:WWL

[YDM99] Stuart Young, Ross Dewstow, and Mae McSporran. Who wants to learn online? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):207, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yeager:1991:TCP

[Yea91] Dorian P. Yeager. Teaching concurrency in the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):155–161, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yerion:1994:DNF

[Yer94] Kathie A. Yerion. The Dutch National Flag Problem revisited as an introductory abstract data type. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(2): 36–40, June 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:1995:IPA

[YJ95] Lan Yang and Lan Jin. Integrating parallel algorithm design with parallel machine models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):131– 135, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Young:1999:IPG

[YJ99] Alison Young and Donald Joyce. Innovation in post graduate computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):205, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:1998:DIW

[YLQ98] Andrew Yang, James Linn, and David Quadrato. Developing integrated Web and database applications using JAVA applets and JDBC drivers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):302–306, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yerion:1995:GCL

[YR95] Kathie A. Yerion and Jane A. Rinehart. Guidelines for collaborative learning in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (4):29–34, December 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yaverbaum:1995:UMT

[YSKW95] Gayle Yaverbaum, Eric Stein, Russell C. Kick, and F. Stuart Wells. Using multimedia technology: different approaches and controversial issues (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):396– 397, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:1994:TBT

[Zac94] Joseph L. Zachary. Tutorial-based teaching of introductory programming classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):136–140, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:1997:CAT

[Zac97a] Joseph L. Zachary. A comprehensive approach to teaching programming to science and engineering majors (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):391, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:1997:GSP

[Zac97b] Joseph L. Zachary. The gestalt of scientific programming: problem, model, method, implementation, assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):238–242, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zagursky:1998:ITF

[Zag98] Valery Zagursky. Information technology for flexible and learning and training (poster). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):312, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zaidman:1996:CDA

[Zai96] Marsha Zaidman. C++ dynamic arrays vs. linked lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):26–28, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zoltowski:1999:CTD

[ZBD99] Bogdan Zoltowski, Roger Boyle, and John Davy. Curriculum and teaching delivery change in an international context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):33–35, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zweigoron:1990:OOP

[ZBK^{+}90] Julie Zweigoron, John Bierbauer, Scott Knaster, Tomasz Pietrzykowski, and John Pugh. Object oriented programming in the computer science curriculum (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):260, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ziegler:1999:IPD

[ZC99] Uta Ziegler and Thad Crews. An integrated program development tool for teaching and learning how to program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):276–280, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ziegler:1998:EIM

[Zie98] Uta Ziegler. Extending an iterator model for binary trees to backtracking problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):185–189, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ziegler:1999:DLI

[Zie99] Uta Ziegler. Discovery learning in introductory operating system courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):321–325, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:1995:ELC

[ZJEP95] Joseph L. Zachary, Christopher R. Johnson, Eric N. Eide, and Kenneth W. Parker. An entry-level course in computational engineering and science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):209–213, March 1995. CODEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic).  URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/fortran.bib;
ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Compiler/FORTRAN/fortran3.bib;   ftp:/
/ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/1995.bib.

Zhao:1998:TTC

[ZLS98] Yuan Zhao, John Lowther, and Ching-Kuang Shene. A tool for teaching curve design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):97–101, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zanconi:1995:EPC

[ZMS95] Marcelo Zanconi, Norma Moroni, and Perla Señas. An educational project in computer science for primary and high school. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(3):27–33, September 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhou:1999:TSD

[ZZLS99] Yan Zhou, Yuan Zhao, John L. Lowther, and Ching-Kuang Shene. Teaching surface design made easy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), REFERENCES 308

31(1):222–226, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).